### v1

# Fast, accurate training and sampling of Restricted Boltzmann Machines

Abstract
Thanks to their simple architecture, Restricted Boltzmann Machines (RBMs) are powerful tools for modeling complex systems and extracting interpretable insights from data. However, training RBMs, as other energy-based models, on highly structured data poses a major challenge, as effective training relies on mixing the Markov chain Monte Carlo simulations used to estimate the gradient. This process is often hindered by multiple second-order phase transitions and the associated critical slowdown. In this paper, we present an innovative method in which the principal directions of the dataset are integrated into a low-rank RBM through a convex optimization procedure. This approach enables efficient sampling of the equilibrium measure via a static Monte Carlo process. By starting the standard training process with a model that already accurately represents the main modes of the data, we bypass the initial phase transitions. Our results show that this strategy successfully trains RBMs to capture the full diversity of data in datasets where previous methods fail. Furthermore, we use the training trajectories to propose a new sampling method, parallel trajectory tempering, which allows us to sample the equilibrium measure of the trained model much faster than previous optimized MCMC approaches and obtain a better estimation of the log-likelihood. We illustrate the success of the training method on several highly structured datasets.
1 Introduction
Energy-based models (EBMs) are a classic approach to generative modeling that has been studied for decades. They were introduced using the Restricted Boltzmann Machine formulation by Smolensky [1] and later further developed by Sejnowski et al. [2]. They provide a straightforward method for modeling effective interactions within complex data distributions and for sufficiently simple energy functions, such as the Boltzmann machine (BM)  [3], it is also possible to interpret and infer the underlying constituent rules from the observed data. This inference strategy is often associated with the inverse Ising problem and pairwise interaction models [4], and it has found a great variety of applications in fields such as neuroscience [5] or computational biology [6]. A recent work has proposed replacing the use of pairwise models with the Restricted Boltzmann Machine (RBM) [7], as it allows the same direct interpretation of its energy function as an explicit many-body interaction model while greatly extending the expressive power of the model. RBMs are also very useful for grouping data into hierarchical families [8]. On the diametrically opposite side (on interpretability) are generative ConvNets [9, 10], where the energy function is formulated as a deep neural network, which are capable of synthesizing photorealistic images but are almost impossible to interpret as a physical model.
The applications of simple EBMs in science are very diverse. For example, they are often used today to encode the Hamiltonian of physical many-body systems, such as Quantum wave functions [11] or the accurate determination of ground state wave functions of strongly interacting and entangled quantum spins [12] or they have proven to be suitable for the representation of the AdS/CFT correspondence in theories of quantum gravity [13, 14]. Simple EBMs are also very common to encode the evolutionary constraints in protein families [6, 15], and to predict mutations [16], or to generate realistic synthetic sequences, such as fake human genomes [17, 18]. These examples show that, despite their somewhat old-fashioned architecture, shallow EBMs are increasingly seen as useful tools for better understanding modern physics/biology, as they allow for a certain level of analytical description.
Despite the appealing modeling properties of RBMs, they are notoriously difficult to train, a challenge common to EBMs in general. The main difficulty arises from the computation of the log-likelihood gradient, which requires an ergodic exploration of a dynamically evolving and potentially complex free energy landscape using Markov Chain Monte Carlo (MCMC) processes. Recent studies have shown that models trained with non-convergent MCMC processes suffer from out-of-equilibrium dynamic memory effects [19, 20, 21]. This dynamical behavior can be explained analytically using moment-matching arguments  [19, 22]. While exploiting these effects can yield fast and accurate generative models, even for highly structured data [23] or high-quality images with RBMs [24], this approach results in a sharp separation between the model’s Gibbs-Boltzmann distribution and the dataset distribution, thereby undermining the interpretability of the model parameters  [22, 7]. Thus, to extract meaningful information from datasets using RBMs, it is essential to ensure proper mixing of the chains during training, in short, one needs equilibrium models.
Both the ability to train an RBM in equilibrium and to generate convincing new samples from its equilibrium measure strongly depend on the dataset in question. For typical image datasets such as MNIST or CIFAR-10, good RBMs can be obtained by increasing the number of MCMC steps. However, this approach is no longer feasible for highly structured datasets [25]. Datasets from which one seeks scientific insights are often highly structured, such as genomics/proteomics data or low-temperature many-body physical systems. These datasets typically exhibit distinct clusters, identifiable via principal component analysis (PCA) which form distant groups of similar entries. We show an example of the PCA of 3 clustered dataset we will be studying in this work in Fig. 1; details about these datasets are given in the caption and in the Supplemental Information (SI). During training, the model must evolve from an initial normal distribution to an increasingly multimodal distribution. Sampling from multimodal distributions is particularly challenging because the mixing times are determined by the transition times between modes. But this is not the only difficulty. These distant modes are encoded by second-order phase transitions during training [26, 27, 28], leading to diverging mixing times in these regions — a phenomenon known as critical slowdown —, which means that mixing times are expected to grow with a power of their system size. This sampling challenge not only hinders the training process, but also limits the model’s ability to generate new samples. Obtaining new and independent configurations would require an impractically large number of sampling steps.
Figure 1: Clustered datasets. In A-C we show the 4 different clustered data sets that we will consider in this paper, projected onto their first two PCA components. In A we show the data of the MNIST 01 dataset (both projected and some instances), which contains only the 0-1 images of the complete MNIST dataset. In B, we show the Mickey dataset, an artificial dataset whose PCA forms a “Mickey"” face shape. In C, we show data from the Human Genome Dataset (HGD), which contains binary vectors each corresponding to a human individual and whose sites correspond to selected genes. A value of 1 at a particular position means that a mutation was observed there compared to an individual reference sequence. Details of these data sets can be found in the SI. In D-F we show the samples we generate with the low-rank RBMs that are used as initial point of a standard training.
2 Related work
Training EBMs by maximizing log-likelihood has long been a challenge in the community [29, 10]. EBMs gained popularity with the introduction of the contrastive divergence algorithm [30], in which a set of parallel chains is initialized on independent examples in the minibatch and the MCMC process iterates for a few steps. Despite its widespread use, this algorithm yields models with poor equilibrium properties that are ineffective as generative models [31, 32, 21]. An improvement is the persistent contrastive divergence (PCD) algorithm [33], which maintains a permanent chain in which the last configurations used to estimate the previous gradient update are reused. PCD acts like a slow annealing process improving gradient estimation quality. However, it often fails on clustered data as the statistical properties of the permanent chain quickly move away from the equilibrium measure and degrade the model [25]. This problem, which is primarily related to phase coexistence, can be addressed with constrained MCMC methods if appropriate order parameters are identified. For RBMs, these order parameters are related to the singular value decomposition of the model coupling matrix, which enables efficient reconstruction of multimodal distributions [25]. Although this method is effective for evaluating model quality, it is too computationally intensive to be used in training, even if it leads to models with good equilibrium properties. Other optimized MCMC methods, such as the Parallel Tempering (PT) [34] algorithm, simulate multiple models at different temperatures, facilitating mixing through temperature exchange [35, 32]. However, PT is costly and often ineffective, especially because EBMs undergo first-order phase transitions at the temperature where PT typically fails because one needs too many temperatures to make the moves accepted. We will see below that a more appropriate approach exchanges the models at different training times, which only implies crossing second-order phase transitions.
The population annealing algorithm, which reweights parallel chains during learning based on their relative weight changes during parameter updates, was proposed as an alternative [36]. Similarly, reweighting chains using non-equilibrium physics concepts such as the Jarzynski equality has been proposed [37]. Both approaches struggle with highly structured data sets. To prevent the different chains to get too correlated around the training phase transitions, one must either increase the number of sampling steps or decrease the learning rate, which in practice means very long training processes to ensure a proper equilibrium training. Another strategy is to use EBMs as corrections for straightforward-to-sample flow-based models [38]. This simplifies sampling and learning, but sacrifices the interpretability of the energy function, which was our goal. An evolving flow model can be used as a fast sampling moves proposer for the EBM [39] objective. This method requires the training of two different networks in parallel and may result in the drop of the move acceptancy as the EBM becomes specialized.
For RBMs, a recent method called “stacked tempering" [40] dramatically speeds up sampling by training smaller RBMs with latent variables from previous models, allowing fast updates to be proposed using a PT like algorithm. Authors also showed that this algorithm was much faster than the standard PT. While effective, it is too cumbersome for use in training. Also for RBMs, it has recently been shown that it is possible to train a low-rank RBM that accurately reproduces the statistics of the data projected along the $d$ first data principal directions through a convex and very fast optimization process (see [41] and the discussion below). This low-rank model can be seen as a good approximation to the correct RBM needed to describe the data, and has the nice property that it can be efficiently sampled via a static Monte Carlo process.
In this paper, we will show how to drastically reduce training times by starting the RBM training process at this low-rank RBM, as this means that the first and strongest dynamic effects associated with them are directly bypassed. We also show that one can exploit the training trajectory to develop an effective sampling method, the parallel trajectory tempering (PTT) that outperforms the “stacked tempering" [40] and only requires saving a reduced number of models during the training. This strategy also allows to obtain reliable estimations for the log-likelihood in well-trained models, much better than those obtained with the standard Annealing Important Sampling (AIS) techniques [42]. Using both strategies, we show that we are able to train and evaluate methods that accurately represent the different modes in the dataset, where standard methods lead to mode collapse effects.
3 The Restricted Boltzmann Machine
The RBM is composed by $N_{\mathrm{v}}$ visible nodes and $N_{\mathrm{h}}$ hidden nodes. In our study, we primarily use binary variables $\{0,1\}$ or ±$\pm 1$ for both layers. The two layers (visible and hidden) interact via a weight matrix $\bm{w}$, with no direct couplings within a given layer. Variables are also adjusted by visible and hidden local biases, $\bm{\theta}$ and $\bm{\eta}$, respectively. The Gibbs-Boltzmann distribution for this model is expressed as
$p(\bm{v},\bm{h})=\frac{1}{Z}\exp\left[-\mathcal{H}(\bm{v},\bm{h})\right]\text{where }\mathcal{H}(\bm{v},\bm{h})=-\sum_{ia}v_{i}w_{ia}h_{a}-\sum_{i}\theta_{i}v_{i}-\sum_{a}\eta_{a}h_{a},$ (1)
where $Z$ is the partition function of the system. As with other models containing hidden variables, the training objective is to minimize the distance between the empirical distribution of the data, $p_{\mathcal{D}}(\bm{v})$, and the model’s marginal distribution over the visible variables, $p(\bm{v})=\sum_{\bm{h}}\exp\left[-\mathcal{H}(\bm{v},\bm{h})\right]/Z=\exp\left[-H(\bm{v})\right]/Z$. Minimizing the Kullback-Leibler divergence is equivalent to maximizing the likelihood of observing the dataset in the model. Thus, the log-likelihood $\mathcal{L}=\langle-H(\bm{v})\rangle_{\mathcal{D}}-\log Z$ can be maximized using the classical stochastic gradient ascent. For a training dataset $\mathcal{D}=\{{\bm{v}}^{(m)}\}_{m=1,\dots,M}$, the log-likelihood gradient is given by
$\frac{\partial\mathcal{L}}{\partial w_{ia}}$ $=\langle v_{i}h_{a}\rangle_{\mathcal{D}}-\langle v_{i}h_{a}\rangle_{\rm RBM},\;\frac{\partial\mathcal{L}}{\partial\theta_{i}}=\langle v_{i}\rangle_{\mathcal{D}}-\langle v_{i}\rangle_{\rm RBM},\;\frac{\partial\mathcal{L}}{\partial\eta_{a}}=\langle h_{a}\rangle_{\mathcal{D}}-\langle h_{a}\rangle_{\rm RBM},$ (2)
where $\left\langle{\cdot}\right\rangle_{\mathcal{D}}$ denotes the average with respect to the entries in the dataset, and $\left\langle{\cdot}\right\rangle_{\mathrm{RBM}}$ with respect to $p(\bm{v},\bm{h})$. Since $Z$ is intractable, the model averages in the gradient are typically estimated using $N_{\mathrm{s}}$ independent MCMC processes, and observable averages $\left\langle{o(\bm{v},\bm{h})}\right\rangle_{\mathrm{RBM}}$ are replaced by $\sum_{r=1}^{R}o(\bm{v}^{(r)},\bm{h}^{(r)})/R$, with $(\bm{v}^{(r)},\bm{h}^{(r)})$ being the last configurations reached with each of the $r=1,\ldots,R$ parallel chains. To obtain reliable estimates, it should be ensured that each of the Markov chains mix well before each parameter update. However, ensuring equilibrium at each update is impractical, slow and tedious. The common use of non-convergent MCMC processes is the cause of most difficulties and weird dynamical behaviors encountered in training RBMs [21].
Typical MCMC mixing times in RBMs are very small at the beginning of the training and grow as it progresses [21], suffering with sharp increases every-time the training trajectory crosses each of the critical transitions that give birth to new modes [28]. In order to minimize out-of-equilibrium effects, it is often useful to keep $R$ permanent (or persistent) chains, which means that the last configurations reached with the MCMC process used to estimate the gradient at a given parameter update $t$, $\bm{P}_{t}\equiv\{(\bm{v}^{(r)}_{t},\bm{h}^{(r)}_{t})\}_{r=1}^{R}$, are used to initialize the chains of the subsequent update $t+1$. This algorithm is typically referred as PCD. In this scheme, the process of training can be mimicked to a slow cooling process, only that instead of varying a single parameter, the temperature, a whole set of parameters $\bm{\Theta}_{t}=(\bm{w}_{t},\bm{\theta}_{t},\bm{\eta}_{t})$ are updated at every step to $\bm{\Theta}_{t+1}=\bm{\Theta}_{t}+\gamma\bm{\nabla}\mathcal{L}_{t}$ with $\bm{\nabla}\mathcal{L}_{t}$ being the gradient in Eq. (2) estimated using the configurations in $\bm{P}_{t}$, and $\gamma$ being the learning rate.
4 The low-rank RBM pretrained
In Ref. [41], it was shown that it is possible to train exactly (i.e. by direct numerical integration instead of MCMC sampling) an RBM containing a reduced number of modes in the weight matrix $\bm{W}$ by exploiting a mapping between the RBM and a Restricted Coulomb Machine and solving a convex optimization problem, see the SI. In other words, it is possible to train a RBM with a coupling matrix of this simplified form
$\bm{W}=\sum_{\alpha=1}^{d}w_{\alpha}\bar{\bm{u}}_{\alpha}\bm{u}_{\alpha}^{\top},\qquad\text{with}\qquad(\bm{u}_{\alpha},\bar{\bm{u}}_{\alpha})\in{\mathbb{R}}^{N_{v}}\times{\mathbb{R}}^{N_{h}},$ (3)
and where the right singular vectors $\{\bm{u}_{\alpha}\}_{\alpha=1}^{d}$ correspond exactly to the first $d$ principal directions of the data set. Under this assumption, it is possible to write $p(\bm{v})$ only as a function of $d$ order parameters given by the magnetizations along each of the $u_{\alpha}$ components, $m_{\alpha}(\bm{v})=\bm{u}_{\alpha}\cdot\bm{v}/\sqrt{{N_{\mathrm{v}}}}$, and in particular,
$H(\bm{v})=-\sum_{a}\log\cosh\left(\sqrt{N_{\mathrm{v}}}\bar{u}_{a}\sum_{\alpha=1}^{d}w_{\alpha}m_{\alpha}+\eta_{a}\right)=\mathcal{H}(\bm{m}(\bm{v})),$ (4)
where $\bm{m}=\left(m_{1},\ldots,m_{\alpha}\right)$. As proposed in [41], the optimal parameters of such a model can basically be determined by solving a regression problem. We describe this method in details in the SI. This means that once the model is trained, we obtain a probability $p(\bm{m})$ defined on a much lower dimension than the original $p(\bm{v})$. Such a probability can be straightforwardly sampled using inverse transform sampling. Since this method requires a discretization of the $\bm{m}$-space both for training and generation, we cannot consider intrinsic space dimension $d>4$ dimensions in practice. These low-rank RBMs are then trained to reproduce the statistics of the dataset projected in its first $d$ principal components. Despite their simplicity, the low-rank models are already able to generate an approximate version of the dataset, as shown in Fig. 1–D-F for the 4 datasets previously presented.
In the initial stage of the standard learning process, the model encodes the strongest PCA components of the data through multiple critical transitions [26, 27, 28]. Pre-training with the low-rank construction allows us to bypass these transitions and avoid out-of-equilibrium effects caused by critical slowing down associated to these transitions. Once the main directions are incorporated, training can efficiently continue with standard algorithms like PCD, as the mixing times of pre-trained machines tend to be much shorter. In particular, in the PCD-100 training with MNIST01, relaxation times for the visible variables’ time correlation reach $5\cdot 10^{5}$ MCMC steps at the first three transitions, coinciding with the growth of singular values in the model weight matrix $W$. In contrast, the pre-trained machine has a much shorter relaxation time of $\sim 10^{3}$, allowing us to safely restart the PCD process from a set of equilibrium samples generated by static sampling of the low-rank RBM.
Figure 2: We compare the equilibrium samples generated by RBMs trained on the Mickey, MNIST01, and HGD datasets using three different training schemes: Jarzynski (JarRBM), PCD, and PCD initialized on low-rank RBMs (used to generate the samples in Fig. 1–D-F). To assess the fitting of the modes, we show a density plot of the projections of the data in the first two principal directions of each dataset. We compare these results with the density plot of the original datasets in the first column.
Overcoming these transitions has dramatic implications for the quality of the models we can train and how accurately they reproduce the statistics of the data. In Fig. 2, we show for 3 datasets the equilibrium samples drawn from 3 RBMs trained with identical number of samples, minibatch size, $k=100$ Gibbs steps, and learning rate $\gamma=0.01$, but different training strategies. In particular, we consider 2 RBMs trained from scratch with the standard PCD [33] and the recently proposed Jarzynski reweighing method [23] (see SI for our specific implementation in the RBM), and a final machine trained with PCD and pre-trained with a low-rank RBM. In all cases, the quality of the generated samples is significantly better when pre-training is used. For the Mickey dataset, neither JarRBM nor normal PCD are able to generate convincing data. For the MNIST01 dataset, all 3 methods are able to generate convincing data, but only Pretrain+PCD is able to correctly balance all modes, as can be seen in Fig. 3, where we compare the histograms of the generated data projected onto the first 3 PCA directions with those of the dataset and a random selection of the generated samples. We see that the pre-training+PCD training perfectly balances the different modes (here we show the first 3 directions, but it goes much further), unlike the other 2 methods, and also generates more diverse images. We can also compare the log-likelihood of all 3 models and find that the pre-trained RBM achieves higher values. At this point, it is important to emphasize that in order to properly quantify the increase in log-likelihood, we need to use the PTT algorithm (see section 5) to correctly thermalize in these well-trained machines. For comparison, we show our PTT measure in dark and solid lines, while the standard AIS [42] estimate is shown in light dashed lines.
Figure 3: We compare the samples generated by the 3 RBMs (JarRBM, PCD, pretrain+PCD) trained with MNIST01 data. In A, we show the histograms of the generated data projected on the first, second and third principal directions with those of the dataset. We see that only the pretrain+PCD correctly balances the different modes. In B we show 10 images generated by each machine. In C, we compare the log-likelihood of each model’s dataset as a function of training time. The dark and full curves were obtained using the PTT algorithm discussed in section  5, and the lighter and dashed curves using the AIS method [42].
Already from the scatter plots we see that the pre-training has a dramatic effect in obtaining models where all modes are properly balanced, but also has important effects in the maximum test-likelihood we can achieve. In all cases, these equilibrium samples are drawn using the trajectory PT algorithm that will be explained in the next section, and the log-likelihood obtained using the equilibrium configurations obtained at different epochs as a result of the trajectory PT flow.
5 Standard Gibbs sampling vs. Parallel Trajectory Tempering (PTT)
One major challenge with structured datasets is quantifying the model’s quality, since sampling the equilibrium measure of a well-trained model is often too time-consuming. This affects the reliability of generated samples and indirect measures as log-likelihood’s estimation through Annealing Importance Sampling (AIS) [42], making them inaccurate and meaningless.
Figure 4: Comparison between PTT and classical Gibbs sampling for the MNIST01 dataset (A and B, respectively) and the human genome dataset (C and D, respectively). In A and C, we show the trajectory of two independent chains (red and orange) projected onto the PCA along the sampling process of the pretraining+PCD model for $10^{4}$ MCMC steps. The black contour represents the density profile of the dataset and the position of the chains is plotted every 10 steps. In B and D we show the average number of jumps from one cluster to another as a function of the MCMC steps performed. The average is calculated over a population of 100 chains. In D, we show the average jump time between clusters along the first (solid line) and second (dashed line) principal components of the data.
To illustrate this problem, let us consider the MNIST01 and the HGD datasets. MNIST01 dataset is bimodal and the HGD highly multimodal as shown in their PCA in Figs. 1–A and C. Let us consider that we want to sample the equilibrium measure of the RBMs trained using low-rank RBM pretraining. In order to draw new samples from these models, one would typically run MCMC processes from random initialization and iterate them until convergence. The mixing time is controlled by the jumping time between clusters. To accurately estimate the relative weight between modes, the MCMC processes must be ergodic, requiring many back-and-forth jumps. However, as shown in Figs. 4–A and C for the MNIST01 and HGD datasets, Gibbs sampling dynamics are extremely slow, rarely producing jumps even after $10^{4}$ MCMC steps. The yellow curves in Figs. 4–B and D show the mean number of jumps over 100 independent chains as a function of MCMC steps, indicating that a proper equilibrium generation would require at least $10^{6}-10^{7}$ MCMC steps.
One effective way to accelerate the dynamics is to exploit the training trajectory, where the model progressively specializes through second-order phase transitions. To achieve this, we save RBMs trained at various epochs and propose swaps between configurations of similarly trained models. We call this the Parallel Trajectory Tempering (PTT) algorithm. Unlike the standard Parallel Tempering (PT) algorithm, which attempts swaps configurations between different temperatures, the PTT swaps between model parameters with different degrees of specialization. This approach is more natural for this problem because it involves crossing only second-order transitions, unlike the first-order transitions occurring in temperature annealing. And in fact, we show in Figs. 4–A and C, that this approach allows us to sharply accelerate the dynamics, as opposed to the standard PT algorithm (studied in detail for the MNIST dataset in [40]).
In the PTT algorithm, the configurations $\bm{x}=(\bm{v},\bm{h})$ of neighboring machines indexed by $t$ and $t-1$ are interchanged with the probability
$p_{\mathrm{acc}}(\bm{x}^{t}\leftrightarrow\bm{x}^{t-1})=\min\left(1,\exp\left(\Delta\mathcal{H}^{t}(\bm{x}^{t})-\Delta\mathcal{H}^{t}(\bm{x}^{t-1})\right)\right).$
This move satisfies detailed balance with our target equilibrium distribution $p(\bm{x})=\exp(-\mathcal{H}(\bm{x}))/Z$, ensuring that the moves lead to the same equilibrium measure. As “nonspecialized" models mix very quickly, either because the distribution is essentially Gaussian at the initialisation of a standard training, or because the low-rank RBM can be sampled with a static Monte Carlo process (yielding independent configurations each time), the trajectory flow significantly accelerates convergence to equilibrium. The time interval between successive machines is selected in such a way that the probability of accepting interchanges between neighboring machines remains around 0.3. Pre-trained machines require a significant fewer number of models to be effective, because most selected models are positioned at the most prominent phase transitions. We give the number of machines used for each sampling process in the SI. We also provide there a specific and detailed description of the algorithm used.
In the red curves in Fig. (4)–B and D, we show the number of jumps between clusters as a function of the number of elementary MCMC steps, which in the PTT scheme refer to 1 Gibbs sampling step + one swap proposal. For the DNA dataset, we have two measures corresponding to jumps along the two principal component directions. We observe at $10^{4}$ MCMC steps an increase of the number of jumps by a factor of 80 for MNIST01 and by a factor of 1350 for the HGD in this machine, although we achieve higher factors in other machines, as we show in the SI. The sampling of RBMs training on the MNIST01 dataset was the subject of the study of the “stacked tempering" algorithm in  [40]. If we compare the numbers with their work, we see that we achieve a 3-4 times higher speedup factor, where our model has the advantage that it does not need additional training, but simply uses the stored machines correctly.
Another desirable advantage of our PTT algorithm is that we can easily use it to compute an improved estimate of the AIS log-likelihood, except that in our case we consider the training trajectory instead of a cooling process and use the equilibrium samples obtained for each of the models to compute the model averages. In Figs. 3–C  5–A we compare the log-likelihood estimates obtained with our method (AIS-PTT) in full and dark lines and in light and dashed lines the AIS estimate (AIS). We see that both measures coincide for most parts of the training and that they split when the sampling becomes too long to thermalize along the temperature annealing curve in AIS. This effect is particularly evident for the JarRBM run in  5–A, where AIS takes a long time to recognize that the model suffers from a strong mode-collapse effect.
6 Overfitting and privacy loss as quality indicators
In this section, we examine the quality of the samples generated, regarding overfitting and privacy criteria which have been defined for genomic data in particular. We look at this on the models trained with PCD with and without pre-training. We do not include the Jarzysnki method here, as this method fails to obtain a reliable model as clearly shown in the evolution of the Log-likelihood in Fig. 5. We focus on the human genome dataset, as shown in Fig. 1–C, to evaluate the ability of various state-of-the-art generative models to generate realistic fake genomes while minimizing privacy concerns (i.e., reducing overfitting). Recent studies [17, 18] have thoroughly investigated this for a variety of generative models. Both studies concluded that the RBM was the most accurate method for generating high-quality and private synthetic genomes. The comparison between models relies primarily on the Nearest Neighbor Adversarial Accuracy ($AA_{\rm TS}$) and privacy loss indicators, introduced in Ref. [43], which quantify the similarity and the level of "privacy" of the data generated by a model w.r.t. the training set. We have $AA_{\rm TS}=\frac{1}{2}\bigl{(}AA_{\rm True}+AA_{\rm Synth}\bigr{)}$ where $AA_{\rm True}$ [resp. $AA_{\rm Synth}$] are two quantities in $[0,1]$ obtained by merging two sets of real and synthetic data of equal size $N_{s}$ and measuring respectively the frequency that a real [rep. synthetic] has a synthetic [resp. real] as nearest neighbor. If the generated samples are statistically indistinguishable from real samples, both frequencies $AA_{\rm True}$ and $AA_{\rm Synth}$ should converge to 0.5 at large ${N_{\mathrm{s}}}$. $AA_{\rm TS}$ can be evaluated both with train or test samples and the privacy loss indicator is defined as ${\rm Privacy\ loss}=AA_{\rm TS}^{\rm test}-AA_{\rm TS}^{\rm train}$ and is expected to be strictly positive. Fig. 5 shows the comparison of $AA_{\rm TS}$ and privacy loss values obtained with our two models, demonstrating that the pre-trained RBM clearly outperforms the other model, and even achieves better results ($AA_{\rm TS}$ values much closer to 0.5) than those discussed in [17, 18].
Figure 5: We compare the quality of the RBMs trained with the human genome data (HGD). In A, we show the log-likelihood as a function of the training epochs for the 3 training procedures. Solid lines correspond to AIS-PTT and dashed lines to AIS. The JarRBM falls down because the training breaks eventually. In B and C we compare privacy and overfitting based on the $AA_{\rm TS}$ indicator.
7 Conclusions
We have shown that the strategy of initiating the training on a pre-trained low-rank RBM is an extremely effective strategy to obtain high quality models for structured datasets that accurately represent all the modes in the datasets and with significantly higher log-likelihoods. We have also shown that the models obtained in that way are: (i) better generative models than those obtained with standard trainings, both, in the sense that they over-fit less at the same time they are more indistinguishable from the test samples, (ii) they display faster relaxational dynamics.
We have also proposed a new fast sampling method that exploits the progressive learning of features in the training of RBMs to design an efficient trajectory PT strategy that allows accelerating the parallel Gibbs sampling dynamics by many orders of magnitude and overcome the performance of recent efficient sampling methods without adding any extra cost than saving models during the training.
Both strategies for training and sampling are very general, and could be generalized to more complex EBMs. In this sense, the low-rank RBM model could be used as a more efficient pre-initialisation in deeper structures, and the trajectory PT algorithm is suitable to be directly used in any EBM no matter how complex it is.
8 Code availability
The code and datasets are available at https://github.com/nbereux/fast-RBM.
9 acknowledgments
The authors acknowledge financial support by the Comunidad de Madrid and the Complutense University of Madrid (UCM) through the Atracción de Talento program (Refs. 2019-T1/TIC-13298), and to project PID2021-125506NA-I00 financed by the Ministerio de Economía y Competitividad, Agencia Estatal de Investigación (MICIU/AEI /10.13039/501100011033), and the Fondo Europeo de Desarrollo Regional (FEDER, UE).
Appendix A Details of the pre-training of a low rank RBM
A.1 The low-rank RBM and its sampling procedure
Our goal is to pre-train an RBM to directly encode the first $d$ principal modes of the dataset in the model’s coupling matrix. This approach avoids the standard procedure of progressively encoding these modes through a series of second-order phase transitions, which negatively impact the quality of gradient estimates during standard training. It also helps prevent critical relaxation slowdown of MCMC dynamics in the presence of many separated clusters.
Given a dataset, we want to find a good set of model parameters ($\bm{w}$, $\bm{\theta}$ and $\bm{\eta}$) for which the statistics of the generated samples exactly match the statistics of the data projected onto the first $d$ directions of the PCA decomposition of the training set. Let us call each of these $\alpha=1,\ldots,d$ projections $m_{\alpha}=\bm{u}_{\alpha}\cdot\bm{v}/\sqrt{N_{\mathrm{v}}}$ the magnetizations along the mode $\alpha$, where $\bm{u}_{\alpha}$ is the $\alpha$-th mode of the PCA decomposition of the dataset. A simple way to encode these $d$-modes is to parameterize the $w$-matrix as:
$\bm{w}=\sum_{\alpha=1}^{d}w_{\alpha}\bar{\bm{u}}_{\alpha}\bm{u}_{\alpha}^{\top},\qquad\text{with}\qquad(\bm{u}_{\alpha},\bar{\bm{u}}_{\alpha})\in{\mathbb{R}}^{N_{v}}\times{\mathbb{R}}^{N_{h}},$ (5)
where $\bm{u}$ and $\hat{\bm{u}}$ are respectively the right-hand and left-hand singular vectors of $\bm{w}$, the former being directly given by the PCA, while $w_{\alpha}$ are the singular values of $\bm{w}$. Using this decomposition, the marginal energy on the visible variables, $\mathcal{H}(\bm{v})=\log\sum_{\bm{h}}\exp\mathcal{H}(\bm{v},\bm{h})$ can be rewritten in terms of these magnetizations $\bm{m}\equiv(m_{1},\ldots,m_{d})$
$\mathcal{H}(\bm{v})=-\sum_{a}\log\cosh\left(\sqrt{N_{\mathrm{v}}}\bar{u}_{a}\sum_{\alpha=1}^{d}w_{\alpha}m_{\alpha}+\eta_{a}\right)=\mathcal{H}(\bm{m}(\bm{v})).$ (6)
Now, the goal of our pre-training is not to match the entire statistics of the data set, but only the marginal probability of these magnetizations. In other words, we want to model the marginal distribution
$p_{\mathrm{emp}}(\bm{m})\equiv\sum_{\bm{v}}p_{\mathrm{emp}}(\bm{v})\prod_{\alpha=1}^{d}\delta\left(m_{\alpha}-\frac{1}{\sqrt{N_{\mathrm{v}}}}\bm{u}_{\alpha}^{T}\bm{v}\right),$ (7)
where $\delta$ is the Dirac $\delta$-distristribution. In this formulation, the distribution of the model over the magnetization $\bm{m}$ can be easily characterized
$p(\bm{m})$ $=\frac{1}{Z}\sum_{\bm{v}}e^{-\mathcal{H}(\bm{v})}\prod_{\alpha=1}^{d}\delta\left(m_{\alpha}-\frac{1}{\sqrt{N_{\mathrm{v}}}}\bm{u}_{\alpha}^{T}\bm{v}\right)$ (8)
$=\frac{1}{Z}\mathcal{N}(\bm{m})\exp{\sum_{a}\log\cosh\left(\bar{u}_{a}\sum_{\alpha=1}^{d}w_{\alpha}m_{\alpha}+\eta_{a}\right)}$ (9)
$=\frac{1}{Z}e^{-\mathcal{H}(\bm{m})+N_{\mathrm{v}}s(\bm{m})}=\frac{1}{Z}e^{-N_{\mathrm{v}}f(\bm{m})}$ (10)
where $\mathcal{N}(\bm{m})=\sum_{\bm{v}}\prod_{\alpha=1}^{d}\delta\left(m_{\alpha}-\frac{1}{\sqrt{N_{\mathrm{v}}}}\bm{u}_{\alpha}^{T}\bm{v}\right)$ is the number of configurations with magnetizations $\bm{m}$, and thus $S(\bm{m})=\log N(\bm{m})/N_{\mathrm{v}}$ is the associated entropy. Now, for large $N_{\mathrm{v}}$ the entropic term can be determined using large deviation theory, and in particular the Gärtner-Ellis theorem:
$p_{\mathrm{prior}}(\bm{m})=\frac{e^{N_{\mathrm{v}}s(\bm{m})}}{2^{N_{\mathrm{v}}}}\approx\exp\left(-N_{\mathrm{v}}\mathcal{I}(\bm{m})\right),$ (11)
with the rate function
$\mathcal{I}(\bm{m})=\sup_{\bm{\mu}}\left[\bm{m}^{T}\bm{\mu}-\phi(\bm{\mu})\right]=\bm{m}^{T}\bm{\mu}^{*}-\phi(\bm{\mu}^{*}),$ (12)
and
$\phi(\bm{\mu})$ $=\lim_{N_{\mathrm{v}}\to\infty}\frac{1}{N_{\mathrm{v}}}\log\left\langle e^{N_{\mathrm{v}}\bm{m}^{T}\bm{\mu}}\right\rangle=\lim_{N_{\mathrm{v}}\to\infty}\frac{1}{N_{\mathrm{v}}}\log\frac{1}{2^{N_{\mathrm{v}}}}\sum_{\bm{v}}e^{\sqrt{N_{\mathrm{v}}}\sum_{\alpha=1}^{d}\mu_{\alpha}\sum_{i}u_{\alpha,i}v_{i}}$ (13)
$=\lim_{N_{\mathrm{v}}\to\infty}\frac{1}{N_{\mathrm{v}}}\sum_{i=1}^{N_{\mathrm{v}}}\log\cosh\left(\sqrt{N_{\mathrm{v}}}\sum_{\alpha=1}^{d}\mu_{\alpha}u_{\alpha,i}\right).$ (14)
Then, given a magnetization $\bm{m}$, we can compute the minimizer $\bm{\mu}^{*}(\bm{m})$ of $\phi(\mu)-\bm{m}^{T}\mu$ which is convex, using e.g. Newton method which converge really fast since we are in small dimension. Note that in practice we will obviously use finite estimates of $\phi$, assuming $N_{v}$ is large enough. As a result we get $\bm{\mu}^{*}(\bm{m})$ satisfying implicit equations given by the constraints given at given $N_{v}$:
$m_{\alpha}=\frac{1}{\sqrt{{N_{\mathrm{v}}}}}\sum_{i=1}^{{N_{\mathrm{v}}}}u_{i}^{\alpha}\tanh\left({\sqrt{{N_{\mathrm{v}}}}\sum_{\beta=1}^{d}u_{i}^{\beta}\mu^{*}_{\beta}}\right).$ (15)
It is then straightforward to check that spins distributed as
$p_{\mathrm{prior}}(\bm{v}|\bm{m})\propto e^{{N_{\mathrm{v}}}{\bm{\mu}^{*}}^{T}\bm{m}(\bm{v})}$ (16)
fulfill well the requirement, as $\left\langle\bm{u}_{\alpha}^{T}\bm{v}/\sqrt{{N_{\mathrm{v}}}}\right\rangle_{p_{\mathrm{prior}}}=m_{\alpha}$. In other words, we can generate samples having mean magnetization $m_{\alpha}$ just by choosing $v_{i}$ as
$p_{\mathrm{prior}}(v_{i}=1|\bm{m})=\text{sigmoid}\left({2\sqrt{{N_{\mathrm{v}}}}\sum_{\alpha=1}^{d}u_{\alpha,i}\mu^{*}_{\alpha}(\bm{m})}\right)$ (17)
The training can therefore be done directly in the subspace of dimension $d$. In Ref. [41], it has been shown that such RBM can be trained by mean of the Restricted Coulomb Machine, where the gradient is actually convex in the parameter’s space. It is then possible to do a mapping from the RCM to the RBM to recover the RBM’s parameters. In brief, the training of the low-dimensional RBM is performed by the RCM, and then the parameters are obtrained via a direct relation between the RCM and the RBM’s parameters. The detail of the definition and of the training of the RCM is detailed in the appendix A.2.
A.2 The Restricted Coulomb Machine
As introduced in [41], it is possible to exactly train a surrogate model for the RBM, called the Restricted Coulomb Machine (RCM), on a low dimensional dataset without explicitly sampling the machine allowing to learn even heavily clustered datasets. We will briefly outline the main steps to train the RCM. A more detailed explanation can be found in Appendix A.2.
The RCM is an approximation of the marginal distribution of the RBM with $\{-1,1\}$ binary variables:
$\mathcal{H}(\bm{v})=-\sum_{i}v_{i}\theta_{i}-\sum_{a}\log\cosh\left(\sum_{i}w_{ia}v_{i}+\eta_{a}\right).$ (18)
We then project both the parameters and variables of the RBM on the first $d$ principal components of the dataset:
$m_{\alpha}\coloneqq\frac{1}{\sqrt{N_{v}}}\sum_{i=1}^{N_{v}}s_{i}u_{i\alpha},\quad w_{\alpha a}\coloneqq\sum_{i=1}^{N_{v}}w_{ia}u_{i\alpha},\quad\theta_{\alpha}\coloneqq\frac{1}{\sqrt{N_{v}}}\sum_{i=1}^{N_{v}}\theta_{i}u_{i\alpha}$ (19)
with $\alpha\in\{1,\ldots,d\}$ and $\bm{v}$ the projection matrix of the PCA. The projected distribution of the model is then given by
$p_{\text{RBM}}(\bm{m})=\frac{\exp\left(N_{v}\left[\mathcal{S}(\bm{m})+\sum_{\alpha=1}^{d}\theta_{\alpha}m_{\alpha}+\frac{1}{N_{v}}\sum_{a=1}^{N_{h}}\log\cosh\left(\sqrt{N_{v}}\sum_{\alpha=1}^{d}m_{\alpha}w_{\alpha a}+\eta_{a}\right)\right]\right)}{Z}$ (20)
where we ignore the fluctuations related to the transverse directions and $\mathcal{S}[\bm{m}]$ accounts for the non-uniform prior on $\bm{m}$ due to the projection of the uniform prior on $\bm{s}$ for the way to compute it.
The RCM is then built by approximating
$\log\cosh(x)\simeq|x|-\log 2,$ (21)
which is valid for $x$ large enough. The probability of the RCM is thus given by:
$p_{\text{RCM}}(\bm{m})=\frac{\exp\left(N_{v}\left[\mathcal{S}(\bm{m})+\sum_{\alpha=1}^{d}\theta_{\alpha}m_{\alpha}+\sum_{a=1}^{N_{h}}q_{a}\left|\sum_{\alpha=1}^{d}n_{\alpha}m_{\alpha}+z_{a}\right|\right]\right)}{Z}$ (22)
where
$q_{a}=\sqrt{N_{v}\sum_{\alpha=1}^{d}w_{\alpha a}^{2}},\quad n_{a}=\frac{w_{\alpha a}}{\sqrt{\sum_{\alpha=1}^{d}w_{\alpha a}^{2}}},\quad z_{a}=\frac{\eta_{a}}{\sqrt{N_{v}\sum_{\alpha=1}^{d}w_{\alpha a}^{2}}}.$ (23)
This can be easily inverted as
$w_{\alpha a}=\frac{1}{\sqrt{N_{v}}}q_{a}n_{a}\qquad\text{and}\qquad\eta_{a}=q_{a}z_{a},$
in order to obtain the RBM from the RCM. The model is then trained through log-likelihood maximization over its parameters. However, this objective is non-convex if all the parameters are trained through gradient ascent. To relax the problem, since we’re in low dimension, we can define a family of hyperplanes $(\bm{n},\bm{z})$ covering the space and let the model only learn the weights of each to the hyperplane. We can then discard the ones with a weight low enough for the approximation (21) to be bad.
The gradients are given by
$\frac{\partial J(\bm{\Theta})}{\partial q_{a}}=\mathbb{E}_{\bm{m}\sim p_{\mathcal{D}}(\bm{m})}\left[|\bm{n}_{a}^{T}\bm{m}+z_{a}|\right]-\mathbb{E}_{\bm{m}\sim p_{\text{RCM}}(\bm{m})}\left[|\bm{n}_{a}^{T}\bm{m}+z_{a}|\right],$ (24)
$\frac{\partial J(\bm{\Theta})}{\partial\theta_{\alpha}}=E_{\bm{m}\sim p_{\mathcal{D}}(\bm{m})}\left[m_{\alpha}\right]-\mathbb{E}_{\bm{m}\sim p_{\text{RCM}}(\bm{m})}\left[m_{\alpha}\right].$ (25)
The positive term is straightforward to compute. For the negative term, we rely on a discretization of the longitudinal space to estimate the probability density of the model and compute the averages.
Appendix B Sampling via Parallel Tempering using the learning trajectory
Figure 6: Scheme of PTT. We Initialize the chains of the models by starting from a configuration $\bm{x}_{0}^{(0)}$ and passing it through the machines along the training trajectory, each time performing $\tilde{k}$ mcmc steps. For pre-train+PCD, $\bm{x}_{0}^{(0)}$ is a sampling from the RCM, otherwise it is a uniform random initialization. The sampling consists of alternating one mcmc step for each model with a swap attempt between adjacent machines. For pre-train+PCD, at each step we sample a new independent configuration for $\mathrm{RBM}_{0}$ using the RCM.
Model Dataset # of machines Alg. # of steps acc. factor @ 104 steps
pre-train+PCD MNIST01 6 (+1) 10000 80
JarJar MNIST01 28 10000 50
PCD MNIST01 13 10000 30
pre-train+PCD Human Genome 6 (+1) 10000 1350
PCD Human Genome 13 10000 7100
Table 1: Performance comparison of different models on various datasets for the sampling using PTT versus Gibbs sampling for $10^{4}$ mcmc steps. The acceleration factor is defined as the ratio of the average number of jumps obtained until $10^{4}$ steps between PTT and Gibbs sampling. For pre-train+PCD, the RCM machine has not to be counted among the list of models (hence the +1) because it is very fast to sample from.
Assuming we have successfully trained a robust equilibrium model, there remains the challenge of efficiently generating equilibrium configurations from this model. Although models trained at equilibrium exhibit faster and more ergodic dynamics compared to poorly trained models, the sampling time can still be excessively long when navigating a highly rugged data landscape. Consequently, we devised a novel method for sampling equilibrium configurations that draws inspiration from the well-established parallel tempering approach. In this traditional method, multiple simulations are conducted in parallel at various temperatures, and configurations are exchanged among them using the Metropolis rule. Unlike this conventional technique, our method involves simultaneously simulating different models that are selected from various points along the training trajectory. This approach is motivated by the perspective that learning represents an annealing process for the model, encountering second-order type phase transitions during training. In contrast, annealing related to temperature changes involves first-order phase transitions, making traditional parallel tempering less effective for sampling from clustered multimodal distributions.
A sketch of the Parallel Trajectory Tempering (PTT) is represented in fig. 6. Specifically, we save $t_{\mathrm{f}}$ models at checkpoints $t=1,\dots,t_{\mathrm{f}}$ along the training trajectory. We denote the Hamiltonian of the model at checkpoint $t$ as $\mathcal{H}_{t}$, and refer to the Hamiltonian of the RCM model as $\mathcal{H}_{0}$. We define $\mathrm{GibbsSampling}(\mathcal{H},\bm{x},k)$ as the operation of performing $k$ Gibbs sampling updates using the model $\mathcal{H}$ starting from the state $\bm{x}$. In all our sampling simulations we used $k=1$.
The first step is to initialize the models’ configurations efficiently. This involves sampling $N$ chains from the RCM model, $\bm{x}_{0}^{(0)}\sim\mathrm{RCMSampling}(\mathcal{H}^{0})$, and then passing the chains through all the models from $t=1$ to $t=t_{\mathrm{f}}$, performing $k$ Gibbs steps at each stage: $\bm{x}_{t}^{(0)}\sim\mathrm{GibbsSampling}(\mathcal{H}_{t},\bm{x}_{t-1}^{(0)},k)$.
The sampling process proceeds in steps where we update the configuration of each model except $\mathcal{H}_{0}$ with $k$ Gibbs steps, and sample a completely new configuration for the RCM model $\mathcal{H}_{0}$. Following this update step, we propose swapping chains between adjacent models with an acceptance probability given by:
$p_{\mathrm{acc}}(\bm{x}_{t}\leftrightarrow\bm{x}_{t-1})=\min\left(1,\exp\left(\Delta\mathcal{H}_{t}(\bm{x}_{t})-\Delta\mathcal{H}_{t}(\bm{x}_{t-1})\right)\right),$ (26)
where $\Delta\mathcal{H}_{t}(\bm{x})=\mathcal{H}_{t}(\bm{x})-\mathcal{H}_{t-1}(\bm{x})$.
We continue alternating between the update step and the swap step until a total of $N_{\mathrm{mcmc}}$ steps is reached. The sampling procedure is illustrated in the following pseudo-code:
Input: Set of models $\{\mathcal{H}_{t}\}$, $t=0,\dots,t_{\mathrm{f}}$, Number of Gibbs steps $k$, Number of MCMC steps $N_{\mathrm{mcmc}}$
Output: Configurations $\bm{x}_{t}$ for $t=1,\ldots,t_{\mathrm{f}}$
Initialize: Sample $N$ chains from the RCM model $\bm{x}_{0}^{(0)}\sim\mathrm{RCMSampling}(\mathcal{H}_{0})$
for $t=1$ to $t_{\mathrm{f}}$ do
     $\bm{x}_{t}^{(0)}\sim\mathrm{GibbsSampling}(\mathcal{H}_{t},\bm{x}_{t-1}^{(0)},\tilde{k})$
end for
for $n=1$ to $\lfloor N_{\mathrm{mcmc}}/k\rfloor$ do
     for $t=1$ to $t_{\mathrm{f}}$ do
         $\bm{x}_{t}^{(n)}\sim\mathrm{GibbsSampling}(\mathcal{H}_{t},\bm{x}_{t}^{(n-1)},k)$
     end for
     Resample $\bm{x}_{0}^{(n)}\sim\mathrm{RCMSampling}(\mathcal{H}_{0})$
     for $t=1$ to $t_{\mathrm{f}}$ do
         Compute acceptance probability
$p_{\mathrm{acc}}(\bm{x}_{t}^{(n)}\leftrightarrow\bm{x}_{t-1}^{(n)})=\min\left(1,\exp\left(\Delta\mathcal{H}_{t}(\bm{x}_{t}^{(n)})-\Delta\mathcal{H}_{t}(\bm{x}_{t-1}^{(n)})\right)\right)$
         Swap $\bm{x}_{t}^{(n)}$ and $\bm{x}_{t-1}^{(n)}$ with probability $p_{\mathrm{acc}}(\bm{x}_{t}^{(n)}\leftrightarrow\bm{x}_{t-1}^{(n)})$
     end for
end for
A comparison of performances between PTT and standard Gibbs sampling is reported in Tab. 1.
Figure 7: Comparison between PTT and standard Gibbs Sampling for RBMs trained using PCD (A and B) and JarRBM (C and D) on the MNIST01 dataset. A and C show the sampling trajectory of two chains recorded every 10 steps for a total of $10^{4}$ mcmc steps. B and D show the average number of jumps of a population of 100 chains as a function of the sampling time.
Figure 8: Comparison between PTT and standard Gibbs Sampling for RBMs trained using PCD on the Human Genome dataset. The sampling has been performed under the same conditions of fig. 7.
Appendix C Training details
We describe in Tables 2 and 3 the datasets and hyperparameters used during training. The test set was used to evaluate the metrics.
Table 2: Details of the datasets used during training.
Name #Samples #Dimensions Train size Test size
Human Genome Dataset (HGD) 4500 805 60% 40%
MNIST-01 10 610 784 60% 40%
Mickey 16 000 1000 60% 40%
Table 3: Hyperparameters used for the training of RBMs.
Name Batch size #Chains #Epochs Learning rate #MCMC steps #Hidden nodes
HGD
PCD 2000 2000 10 000 0.01 100 185
Jar-RBM 2000 10 000 10 000 0.01 100 185
Pre-train+PCD 2000 2000 10 000 0.01 100 185
MNIST-01
PCD 2000 2000 10 000 0.01 100 200
Jar-RBM 2000 10 000 10 000 0.01 100 200
Pre-train+PCD 2000 2000 10 000 0.01 100 200
Mickey
PCD 2000 2000 10 000 0.01 100 100
Jar-RBM 2000 10 000 10 000 0.01 100 100
Pre-train+PCD 2000 2000 10 000 0.01 100 100
All experiments were run on a RTX 4090 with an AMD Ryzen 9 5950X.
Appendix D Training of the RBM using the Jarzynski equation
In this section, we describe a procedure similar to the one introduced in [37] for training the RBM by leveraging the Jarzynki equation. In one of its formulations, the Jarzynski equation states that we can relate the ensemble average of an observable $\mathcal{O}$ with the average obtained through many repetitions of an out-of-equilibrium dynamical process. If we consider the training trajectory of an RBM, $p_{0}\rightarrow p_{1}\rightarrow\dots\rightarrow p_{t-1}\rightarrow p_{t}$, we can write
$\langle\mathcal{O}\rangle_{t}=\frac{\left\langle\mathcal{O}e^{-W_{t}}\right\rangle_{\mathrm{traj}}}{\left\langle e^{-W_{t}}\right\rangle_{\mathrm{traj}}},$ (27)
where the average on the lhs is done over the last model $p_{t}$, the averages on the rhs are taken across many different trajectory realizations and $W_{t}$ is a trajectory-dependent importance factor. By all practical means, under the assumption of having quasi-adiabatic parameters updates, namely $p(\Theta_{t-1}\rightarrow\Theta_{t})=p(\Theta_{t}\rightarrow\Theta_{t-1})$, this means that we can assign to each Markov chain of the simulation $\bm{x}^{(r)}$, $r=1,\dots,R,$ an importance weight given by:
$W_{t}^{(r)}=\sum_{\tau=1}^{t}[\mathcal{H}_{\tau}(\bm{x}^{(r)}_{\tau-1})-\mathcal{H}_{\tau-1}(\bm{x}^{(r)}_{\tau-1})]$ (28)
and then compute the gradient of the log-likelihood by means of a weighted average over the chains:
$\langle\mathcal{O}\rangle_{t}\simeq\frac{\sum_{r=1}^{R}\mathcal{O}(\bm{x}^{(r)})~{}e^{-W^{(r)}_{t}}}{\sum_{r=1}^{R}e^{-W^{(r)}_{t}}}.$ (29)
Notice that, since Eq. (27) is an exact result, the importance weights should, in principle, eliminate the bias brought by the non-convergent chains used for approximating the log-likelihood gradient in the classical PCD scheme. However, after many updates of the importance weights, one finds that only a few chains carry almost all the importance mass. In other words, the vast majority of the chains we are simulating are statistically irrelevant, and we expect to get large fluctuations in the estimate of the gradient because of the small effective number of chains contributing to the statistical average. A good observable for monitoring this effect is the Effective Sample Size (ESS), defined as [37]
$\mathrm{ESS}=\frac{\left(R^{-1}\sum_{r=1}^{R}e^{-W^{(r)}}\right)^{2}}{R^{-1}\sum_{r=1}^{R}e^{-2W^{(r)}}}\in[0,1],$ (30)
which measures the relative dispersion of the weights distribution. A way of circumventing the weight concentration on a few chains, then, is to resample the chain population according to the importance weights every time the ESS drops below a certain threshold, for instance 0.5. After this resampling, all the chain weights have to be set to 1 ($W^{(r)}=0\,\,\forall r=1,\dots,R$).

### v2

@@
-Fast, accurate training and sampling of Restricted Boltzmann Machines
+Fast training and sampling of Restricted Boltzmann Machines
 Abstract
-Thanks to their simple architecture, Restricted Boltzmann Machines (RBMs) are powerful tools for modeling complex systems and extracting interpretable insights from data.
-However, training RBMs, as other energy-based models, on highly structured data poses a major challenge, as effective training relies on mixing the Markov chain Monte Carlo simulations used to estimate the gradient.
-This process is often hindered by multiple second-order phase transitions and the associated critical slowdown.
-In this paper, we present an innovative method in which the principal directions of the dataset are integrated into a low-rank RBM through a convex optimization procedure.
-This approach enables efficient sampling of the equilibrium measure via a static Monte Carlo process.
-By starting the standard training process with a model that already accurately represents the main modes of the data, we bypass the initial phase transitions.
-Our results show that this strategy successfully trains RBMs to capture the full diversity of data in datasets where previous methods fail.
-Furthermore, we use the training trajectories to propose a new sampling method, parallel trajectory tempering, which allows us to sample the equilibrium measure of the trained model much faster than previous optimized MCMC approaches and obtain a better estimation of the log-likelihood.
-We illustrate the success of the training method on several highly structured datasets.
+Restricted Boltzmann Machines (RBMs) are effective tools for modeling complex systems and deriving insights from data.
+However, training these models with highly structured data presents significant challenges due to the slow mixing characteristics of Markov Chain Monte Carlo (MCMC) processes.
+In this study, we build upon recent theoretical advancements in RBM training, focusing on the gradual encoding of data patterns into singular vectors of the coupling matrix, to significantly reduce the computational cost of training (in very clustered datasets) and evaluating and sampling in RBMs in general.
+The learning process is analogous to thermodynamic continuous phase transitions observed in ferromagnetic models, where new modes in the probability measure emerge in a continuous manner.
+Such continuous transitions are associated with the critical slowdown effect, which adversely affects the accuracy of gradient estimates, particularly during the initial stages of training with clustered data.
+To mitigate this issue, we propose a pre-training phase that encodes the principal components into a low-rank RBM through a convex optimization process.
+This approach facilitates efficient static Monte Carlo sampling and accurate computation of the partition function.
+Furthermore, we exploit the continuous and smooth nature of the parameter annealing trajectory to achieve reliable and computationally efficient log-likelihood estimations, enabling online assessment during the training process, and proposing a novel sampling strategy termed parallel trajectory tempering that outperforms previously optimized MCMC methods.
+Our results demonstrate that this innovative training strategy enables RBMs to effectively address highly structured datasets that conventional methods struggle with.
+Additionally, we provide evidence that our log-likelihood estimation is more accurate than traditional, more computationally intensive approaches in controlled scenarios.
+Moreover, the parallel trajectory tempering algorithm significantly accelerates MCMC processes compared to existing and conventional methods.
 1 Introduction
-Energy-based models (EBMs) are a classic approach to generative modeling that has been studied for decades.
-They were introduced using the Restricted Boltzmann Machine formulation by Smolensky [1] and later further developed by Sejnowski et al.
-[2].
-They provide a straightforward method for modeling effective interactions within complex data distributions and for sufficiently simple energy functions, such as the Boltzmann machine (BM)  [3], it is also possible to interpret and infer the underlying constituent rules from the observed data.
-This inference strategy is often associated with the inverse Ising problem and pairwise interaction models [4], and it has found a great variety of applications in fields such as neuroscience [5] or computational biology [6].
-A recent work has proposed replacing the use of pairwise models with the Restricted Boltzmann Machine (RBM) [7], as it allows the same direct interpretation of its energy function as an explicit many-body interaction model while greatly extending the expressive power of the model.
-RBMs are also very useful for grouping data into hierarchical families [8].
-On the diametrically opposite side (on interpretability) are generative ConvNets [9, 10], where the energy function is formulated as a deep neural network, which are capable of synthesizing photorealistic images but are almost impossible to interpret as a physical model.
-The applications of simple EBMs in science are very diverse.
-For example, they are often used today to encode the Hamiltonian of physical many-body systems, such as Quantum wave functions [11] or the accurate determination of ground state wave functions of strongly interacting and entangled quantum spins [12] or they have proven to be suitable for the representation of the AdS/CFT correspondence in theories of quantum gravity [13, 14].
-Simple EBMs are also very common to encode the evolutionary constraints in protein families [6, 15], and to predict mutations [16], or to generate realistic synthetic sequences, such as fake human genomes [17, 18].
-These examples show that, despite their somewhat old-fashioned architecture, shallow EBMs are increasingly seen as useful tools for better understanding modern physics/biology, as they allow for a certain level of analytical description.
+Energy-based models (EBMs) are a long-established approach to generative modeling, with the Boltzmann Machine (BM)  (Hinton & Sejnowski, 1983) and Restricted Boltzmann Machine (RBM) (Smolensky, 1986; Ackley et al., 1985) being among the earliest examples.
+They offer a clear framework for capturing complex data interactions and, with simple energy functions, can reveal underlying patterns, making them especially valuable for scientific applications.
+Pairwise-interacting models (Nguyen et al., 2017) have been widely used for decades for inference applications in fields like neuroscience (Hertz et al., 2011) and computational biology (Cocco et al., 2018).
+More recent research has shown that RBMs can also be reinterpreted as many-body physical models (Decelle et al., 2024; Bulso & Roudi, 2021), analyzed to identify hierarchical relationships in the data (Decelle et al., 2023), or used to uncover biologically relevant constituent features (Tubiana et al., 2019; Fernandez-de Cossio-Diaz, 2025).
+In essence, RBMs present a great compromise between expressibility and interpretability: they are universal approximators capable of modeling complex datasets and simple enough to allow a direct extraction of insights from data.
+This interpretative task is particularly challenging when dealing with more powerful models such as generative convolutional networks EBMs (Xie et al., 2016; Song & Kingma, 2021) and other state-of-the-art models in general.
+At this point, it is also important to stress that RBM are not only interpretable, but they perform particularly well when dealing with tabular datasets with discrete variables, and where data is not abundant, which is often the case in genomics/proteomics or neural recording datasets.
+For instance, RBMs have shown to be among the most reliable approaches to generate synthetic human genomes (Yelmen et al., 2021; 2023) or to model inmunological response (Bravi et al., 2021a; b; 2023) or neural activity (van der Plas et al., 2023; Quiroz Monnens et al., 2024).
 Despite the appealing modeling properties of RBMs, they are notoriously difficult to train, a challenge common to EBMs in general.
 The main difficulty arises from the computation of the log-likelihood gradient, which requires an ergodic exploration of a dynamically evolving and potentially complex free energy landscape using Markov Chain Monte Carlo (MCMC) processes.
-Recent studies have shown that models trained with non-convergent MCMC processes suffer from out-of-equilibrium dynamic memory effects [19, 20, 21].
-This dynamical behavior can be explained analytically using moment-matching arguments  [19, 22].
-While exploiting these effects can yield fast and accurate generative models, even for highly structured data [23] or high-quality images with RBMs [24], this approach results in a sharp separation between the model’s Gibbs-Boltzmann distribution and the dataset distribution, thereby undermining the interpretability of the model parameters  [22, 7].
+Recent studies have shown that models trained with non-convergent MCMC processes suffer from out-of-equilibrium dynamic memory effects (Nijkamp et al., 2019; 2020; Decelle et al., 2021).
+This dynamical behavior can be explained analytically using moment-matching arguments  (Nijkamp et al., 2019; Agoritsas et al., 2023).
+While exploiting these effects can yield fast and accurate generative models, even for highly structured data (Carbone et al., 2023) or high-quality images with RBMs (Liao et al., 2022), this approach results in a sharp separation between the model’s Gibbs-Boltzmann distribution and the dataset distribution, thereby undermining the interpretability of the model parameters as investigated in Agoritsas et al.
+(2023); Decelle et al.
+(2024).
 Thus, to extract meaningful information from datasets using RBMs, it is essential to ensure proper mixing of the chains during training, in short, one needs equilibrium models.
@@
 For typical image datasets such as MNIST or CIFAR-10, good RBMs can be obtained by increasing the number of MCMC steps.
-However, this approach is no longer feasible for highly structured datasets [25].
-Datasets from which one seeks scientific insights are often highly structured, such as genomics/proteomics data or low-temperature many-body physical systems.
-These datasets typically exhibit distinct clusters, identifiable via principal component analysis (PCA) which form distant groups of similar entries.
-We show an example of the PCA of 3 clustered dataset we will be studying in this work in Fig.
-1; details about these datasets are given in the caption and in the Supplemental Information (SI).
+However, this approach is no longer feasible for highly structured datasets (Béreux et al., 2023).
+Natural datasets from which one seeks scientific insights are often highly structured, such as genomics/proteomics data, neural-recordings or low-temperature configurations in physical systems.
+These datasets typically exhibit distinct clusters, identifiable via principal component analysis (PCA) which form distant groups of similar entries often associated to different labels.
+We show an example of the PCA of the 5 datasets we will be studying in this work in Fig.
+1.
+The first 4 datasets are very clustered, while CelebA is not, as it presents a rather compact PCA.
+A very similar (non-clustered) situation is obtained if the entire MNIST dataset is considered.
+Details about these datasets are given in the caption of figure 1.
 During training, the model must evolve from an initial normal distribution to an increasingly multimodal distribution.
-Sampling from multimodal distributions is particularly challenging because the mixing times are determined by the transition times between modes.
+Sampling from multimodal distributions is particularly challenging because the mixing times are controlled by the transition times between modes.
 But this is not the only difficulty.
-These distant modes are encoded by second-order phase transitions during training [26, 27, 28], leading to diverging mixing times in these regions — a phenomenon known as critical slowdown —, which means that mixing times are expected to grow with a power of their system size.
+In RBMs, these distant modes can be shown analytically(Decelle et al., 2017; 2018; Bachtis et al., 2024) to gradually appear in the model’s equilibrium measure during training by a process formally analogous to the ferromagnetic phase transition in simple spin models, the process by which the magnetic spins of a material spontaneously align with a fixed pattern when the temperature goes below a certain transition temperature.
+The RBM training is characterized as a sequence of such phase transitions in which the specialization of the model parameters plays the role of temperature reduction.
+These pattern-encoding phase transitions are continuous phase transitions (also known as second-order transitions or critical transitions), and these types of transitions are associated with a sudden arrest in the characteristic times of the visible variables, a phenomenon known as critical slowdown, which causes the MCMC mixing time to diverge with a power of the system size when the transition point is approached.
+The critical slowdown effects that occur in the training trajectory were recently studied in Bachtis et al.
+(2024).
+A brief introduction to phase transitions and the associated dynamical effects can be found in the SI G.
 This sampling challenge not only hinders the training process, but also limits the model’s ability to generate new samples.
 Obtaining new and independent configurations would require an impractically large number of sampling steps.
-Figure 1: Clustered datasets.
-In A-C we show the 4 different clustered data sets that we will consider in this paper, projected onto their first two PCA components.
-In A we show the data of the MNIST 01 dataset (both projected and some instances), which contains only the 0-1 images of the complete MNIST dataset.
-In B, we show the Mickey dataset, an artificial dataset whose PCA forms a “Mickey"” face shape.
-In C, we show data from the Human Genome Dataset (HGD), which contains binary vectors each corresponding to a human individual and whose sites correspond to selected genes.
-A value of 1 at a particular position means that a mutation was observed there compared to an individual reference sequence.
-Details of these data sets can be found in the SI.
-In D-F we show the samples we generate with the low-rank RBMs that are used as initial point of a standard training.
+Figure 1: Datasets.
+Panels A-E display 5 distinct datasets projected onto their first two PCA components.
+In some instances, the dots are color-coded to indicate different labels.
+In A, the MNIST 01 dataset, featuring images of the digits 0 and 1 from the complete MNIST collection, along with a few sample images.
+In B, the “Mickey” dataset, an artificial dataset whose PCA representation forms the shape of Mickey Mouse’s face.
+In C, the Human Genome Dataset (HGD), which consists of binary vectors representing mutations or non-mutations for individuals compared to a reference genome across selected genes.
+In D, the Ising dataset, showcasing equilibrium configurations of the 2D ferromagnetic Ising model at an inverse temperature of $\beta=0.44$.
+In E, the CelebA dataset in black and white, resized to 32x32 pixels.
+For more details on these datasets, please refer to the SI.
 2 Related work
-Training EBMs by maximizing log-likelihood has long been a challenge in the community [29, 10].
-EBMs gained popularity with the introduction of the contrastive divergence algorithm [30], in which a set of parallel chains is initialized on independent examples in the minibatch and the MCMC process iterates for a few steps.
-Despite its widespread use, this algorithm yields models with poor equilibrium properties that are ineffective as generative models [31, 32, 21].
-An improvement is the persistent contrastive divergence (PCD) algorithm [33], which maintains a permanent chain in which the last configurations used to estimate the previous gradient update are reused.
+Training EBMs by maximizing log-likelihood has long been a challenge in the community (LeCun et al., 2006; Song & Kingma, 2021).
+EBMs gained popularity with the introduction of the contrastive divergence algorithm (Hinton, 2002), in which a set of parallel chains is initialized on independent examples in the minibatch and the MCMC process iterates for a few steps.
+Despite its widespread use, this algorithm yields models with poor equilibrium properties that are ineffective as generative models (Salakhutdinov & Murray, 2008; Desjardins et al., 2010; Decelle et al., 2021).
+An improvement is the persistent contrastive divergence (PCD) algorithm (Tieleman, 2008), which maintains a permanent chain in which the last configurations used to estimate the previous gradient update are reused.
 PCD acts like a slow annealing process improving gradient estimation quality.
-However, it often fails on clustered data as the statistical properties of the permanent chain quickly move away from the equilibrium measure and degrade the model [25].
+However, it often fails on clustered data as the statistical properties of the permanent chain quickly move away from the equilibrium measure and degrade the model (Béreux et al., 2023).
 This problem, which is primarily related to phase coexistence, can be addressed with constrained MCMC methods if appropriate order parameters are identified.
-For RBMs, these order parameters are related to the singular value decomposition of the model coupling matrix, which enables efficient reconstruction of multimodal distributions [25].
+For RBMs, these order parameters are related to the singular value decomposition of the model coupling matrix, which enables efficient reconstruction of multimodal distributions (Béreux et al., 2023).
 Although this method is effective for evaluating model quality, it is too computationally intensive to be used in training, even if it leads to models with good equilibrium properties.
-Other optimized MCMC methods, such as the Parallel Tempering (PT) [34] algorithm, simulate multiple models at different temperatures, facilitating mixing through temperature exchange [35, 32].
-However, PT is costly and often ineffective, especially because EBMs undergo first-order phase transitions at the temperature where PT typically fails because one needs too many temperatures to make the moves accepted.
-We will see below that a more appropriate approach exchanges the models at different training times, which only implies crossing second-order phase transitions.
-The population annealing algorithm, which reweights parallel chains during learning based on their relative weight changes during parameter updates, was proposed as an alternative [36].
-Similarly, reweighting chains using non-equilibrium physics concepts such as the Jarzynski equality has been proposed [37].
-Both approaches struggle with highly structured data sets.
-To prevent the different chains to get too correlated around the training phase transitions, one must either increase the number of sampling steps or decrease the learning rate, which in practice means very long training processes to ensure a proper equilibrium training.
-Another strategy is to use EBMs as corrections for straightforward-to-sample flow-based models [38].
-This simplifies sampling and learning, but sacrifices the interpretability of the energy function, which was our goal.
-An evolving flow model can be used as a fast sampling moves proposer for the EBM [39] objective.
-This method requires the training of two different networks in parallel and may result in the drop of the move acceptancy as the EBM becomes specialized.
-For RBMs, a recent method called “stacked tempering" [40] dramatically speeds up sampling by training smaller RBMs with latent variables from previous models, allowing fast updates to be proposed using a PT like algorithm.
-Authors also showed that this algorithm was much faster than the standard PT.
-While effective, it is too cumbersome for use in training.
-Also for RBMs, it has recently been shown that it is possible to train a low-rank RBM that accurately reproduces the statistics of the data projected along the $d$ first data principal directions through a convex and very fast optimization process (see [41] and the discussion below).
+Other optimized MCMC methods, such as the Parallel Tempering (PT) algorithm  (Hukushima & Nemoto, 1996) (in which the model is simulated in parallel at different temperatures which are exchanged once in a while using the Metropolis rule), have been proposed to improve the training process (Salakhutdinov, 2009; Desjardins et al., 2010) and show important improvements in certain datasets (Krause et al., 2020).
+However, PT is expensive because it requires sampling at multiple temperatures while using samples from only one model.
+It’s also often ineffective with highly clustered data due to first-order phase transitions in EBMs, where modes disappear abruptly at certain temperatures, as discussed by Decelle & Furtlehner (2021a).
+This leads to sharp drops in acceptance rates for exchanges between temperatures near the transition.
+For this reason, first-order transitions typically hinder PT, as many intermediate temperatures are needed to allow sufficient acceptance rates in the exchanges.
+In contrast, continuous transitions, as seen during the training trajectory, cause modes to gradually separate as the model parameters are changed.
+The population annealing algorithm, which adjusts the statistical weights of parallel chains during learning each time the parameters are updated, has been proposed as an alternative approach (Krause et al., 2018).
+Reweighting chains using non-equilibrium physics concepts like the Jarzynski equation has also been explored (Carbone et al., 2024).
+However, both methods struggle with highly structured data, as frequent reweighting (which involves duplicating only a few samples) biases the equilibrium measure due to an insufficient number of effective independent chains.
+To prevent chains from becoming too correlated around transition regions during training, either more sampling steps or a lower learning rate is required, which leads to very long training times to ensure equilibrium.
+In our case, we tried to used both methods with our datasets but the training becomes too unstable or two slow if learning rate and number of Gibbs steps are auto adjusted.
+Another approach uses EBMs as corrections for flow-based models that are easier to sample (Nijkamp et al., 2022), simplifying learning but sacrificing the interpretability of the energy function, a key goal of this work.
+Alternatively, evolving flow models can serve as fast sampling proposers for the EBM (Grenioux et al., 2023), but this requires training two separate networks and the acceptance of the proposed moves is significantly reduced as the EBM becomes more specialized.
+A recent method for RBMs, called “stacked tempering” (Roussel et al., 2023), significantly accelerates sampling by training smaller RBMs using latent variables from previous models.
+This allows for fast updates through a PT-like algorithm, and the authors demonstrated that it is much faster than standard PT.
+However, while this method is effective for sampling a trained model, it is not suitable for learning, as it requires sequentially training multiple models stacked on top of each other.
+Also for RBMs, it has recently been shown in a theoretical work that it should be possible to train a low-rank RBM that accurately reproduces the statistics of the data projected along the $d$ first data principal directions through a convex and very fast optimization process (see Decelle & Furtlehner (2021a) and the discussion below) using a mapping to another model, the so-called Restricted Coulomb Machine.
 This low-rank model can be seen as a good approximation to the correct RBM needed to describe the data, and has the nice property that it can be efficiently sampled via a static Monte Carlo process.
-In this paper, we will show how to drastically reduce training times by starting the RBM training process at this low-rank RBM, as this means that the first and strongest dynamic effects associated with them are directly bypassed.
-We also show that one can exploit the training trajectory to develop an effective sampling method, the parallel trajectory tempering (PTT) that outperforms the “stacked tempering" [40] and only requires saving a reduced number of models during the training.
-This strategy also allows to obtain reliable estimations for the log-likelihood in well-trained models, much better than those obtained with the standard Annealing Important Sampling (AIS) techniques [42].
-Using both strategies, we show that we are able to train and evaluate methods that accurately represent the different modes in the dataset, where standard methods lead to mode collapse effects.
+While this work showed the efficiency of this training process to describe simple synthetic datasets embedded in only one or two dimensions, the use of this algorithm to obtain reliable low-rank RBMs with real data requires further developments that will be discussed in this paper, in particular: (i) the training of a bias, which is crucial to gain almost for free an additional dimension, and to deal with image datasets, and (ii) a tuning of the RBM fast sampling process, which was not properly investigated in the original paper.
+In this paper, we show how starting the RBM training with a low-rank RBM in very clustered datasets can drastically improve the quality of the models by circumventing the initial dynamical arrest effects that occur during the training process, at a very similar computational cost.
+We also show that the training trajectory can be leveraged to estimate the log-likelihood at a very little computational effort, so that both the training and test set log-likelihood can be easily and reliably computed online during training.
+Furthermore, we propose an efficient sampling method, which we call Parallel Trajectory Tempering (PTT).
+This method outperforms optimized MCMC methods such as the traditional Parallel Tempering (PT) or the “Stacked Tempering” (Roussel et al., 2023) only requiring storing the parameters of the model at a few stages during training.
+By combining these strategies, we are able to train equilibrium models that accurately capture different modes in datasets prone to mode collapse using standard methods, easily evaluate RBMs during the training process, and easily sample from multimodal distribution.
 3 The Restricted Boltzmann Machine
@@
 As with other models containing hidden variables, the training objective is to minimize the distance between the empirical distribution of the data, $p_{\mathcal{D}}(\bm{v})$, and the model’s marginal distribution over the visible variables, $p(\bm{v})=\sum_{\bm{h}}\exp\left[-\mathcal{H}(\bm{v},\bm{h})\right]/Z=\exp\left[-H(\bm{v})\right]/Z$.
-Minimizing the Kullback-Leibler divergence is equivalent to maximizing the likelihood of observing the dataset in the model.
+Minimizing the Kullback-Leibler divergence is equivalent to maximizing the likelihood of observing the training dataset in the model.
 Thus, the log-likelihood $\mathcal{L}=\langle-H(\bm{v})\rangle_{\mathcal{D}}-\log Z$ can be maximized using the classical stochastic gradient ascent.
@@
 However, ensuring equilibrium at each update is impractical, slow and tedious.
-The common use of non-convergent MCMC processes is the cause of most difficulties and weird dynamical behaviors encountered in training RBMs [21].
-Typical MCMC mixing times in RBMs are very small at the beginning of the training and grow as it progresses [21], suffering with sharp increases every-time the training trajectory crosses each of the critical transitions that give birth to new modes [28].
+The common use of non-convergent MCMC processes is the cause of most difficulties and weird dynamical behaviors encountered in training RBMs as discussed in Decelle et al.
+(2021).
 In order to minimize out-of-equilibrium effects, it is often useful to keep $R$ permanent (or persistent) chains, which means that the last configurations reached with the MCMC process used to estimate the gradient at a given parameter update $t$, $\bm{P}_{t}\equiv\{(\bm{v}^{(r)}_{t},\bm{h}^{(r)}_{t})\}_{r=1}^{R}$, are used to initialize the chains of the subsequent update $t+1$.
 This algorithm is typically referred as PCD.
-In this scheme, the process of training can be mimicked to a slow cooling process, only that instead of varying a single parameter, the temperature, a whole set of parameters $\bm{\Theta}_{t}=(\bm{w}_{t},\bm{\theta}_{t},\bm{\eta}_{t})$ are updated at every step to $\bm{\Theta}_{t+1}=\bm{\Theta}_{t}+\gamma\bm{\nabla}\mathcal{L}_{t}$ with $\bm{\nabla}\mathcal{L}_{t}$ being the gradient in Eq.
-(2) estimated using the configurations in $\bm{P}_{t}$, and $\gamma$ being the learning rate.
+In this scheme, the process of training can be mimicked to a slow cooling process, only that instead of varying a single parameter, e.g.
+the temperature, a whole set of parameters $\bm{\Theta}_{t}=(\bm{w}_{t},\bm{\theta}_{t},\bm{\eta}_{t})$ are updated at every step to $\bm{\Theta}_{t+1}=\bm{\Theta}_{t}+\gamma\bm{\nabla}\mathcal{L}_{t}$ with $\bm{\nabla}\mathcal{L}_{t}$ being the gradient in equation 2 estimated using the configurations in $\bm{P}_{t}$, and $\gamma$ being the learning rate.
+Typical MCMC mixing times in RBMs are short at the beginning of training but increase as it progresses (Decelle et al., 2021).
+Recent analytical and numerical studies show that the mixing times exhibit sharp, localized peaks when the model encodes a new mode during training (Bachtis et al., 2024).
+This phenomenon, known in physics as critical slowing down, is related to the system going through critical phase transitions.
+For RBMs, the first phase transitions during maximum likelihood training are associated with the encoding of the principal components of the dataset into the singular vectors of the weight matrix $\bm{W}$ (Decelle et al., 2017; 2018; Decelle & Furtlehner, 2021b; Bachtis et al., 2024).
+Since these directions can be determined before training by performing PCA on the data, one can pre-train an RBM to encode them accurately.
+This avoids crossing the initial phase transitions encountered during the standard training and the associated degradation of the gradient calculation.
 4 The low-rank RBM pretrained
-In Ref.
-[41], it was shown that it is possible to train exactly (i.e.
-by direct numerical integration instead of MCMC sampling) an RBM containing a reduced number of modes in the weight matrix $\bm{W}$ by exploiting a mapping between the RBM and a Restricted Coulomb Machine and solving a convex optimization problem, see the SI.
-In other words, it is possible to train a RBM with a coupling matrix of this simplified form
+In Decelle & Furtlehner (2021a), it was shown that it is possible to train exactly an RBM containing a reduced number of modes in the weight matrix $\bm{W}$ by exploiting a mapping between the RBM and a Restricted Coulomb Machine and solving a convex optimization problem, see the SI A.2.
+In other words, it is possible to train an RBM whose coupling matrix has this simplified form
 $\bm{W}=\sum_{\alpha=1}^{d}w_{\alpha}\bar{\bm{u}}_{\alpha}\bm{u}_{\alpha}^{\top},\qquad\text{with}\qquad(\bm{u}_{\alpha},\bar{\bm{u}}_{\alpha})\in{\mathbb{R}}^{N_{v}}\times{\mathbb{R}}^{N_{h}},$ (3)
-and where the right singular vectors $\{\bm{u}_{\alpha}\}_{\alpha=1}^{d}$ correspond exactly to the first $d$ principal directions of the data set.
-Under this assumption, it is possible to write $p(\bm{v})$ only as a function of $d$ order parameters given by the magnetizations along each of the $u_{\alpha}$ components, $m_{\alpha}(\bm{v})=\bm{u}_{\alpha}\cdot\bm{v}/\sqrt{{N_{\mathrm{v}}}}$, and in particular,
-$H(\bm{v})=-\sum_{a}\log\cosh\left(\sqrt{N_{\mathrm{v}}}\bar{u}_{a}\sum_{\alpha=1}^{d}w_{\alpha}m_{\alpha}+\eta_{a}\right)=\mathcal{H}(\bm{m}(\bm{v})),$ (4)
-where $\bm{m}=\left(m_{1},\ldots,m_{\alpha}\right)$.
-As proposed in [41], the optimal parameters of such a model can basically be determined by solving a regression problem.
-We describe this method in details in the SI.
-This means that once the model is trained, we obtain a probability $p(\bm{m})$ defined on a much lower dimension than the original $p(\bm{v})$.
-Such a probability can be straightforwardly sampled using inverse transform sampling.
-Since this method requires a discretization of the $\bm{m}$-space both for training and generation, we cannot consider intrinsic space dimension $d>4$ dimensions in practice.
-These low-rank RBMs are then trained to reproduce the statistics of the dataset projected in its first $d$ principal components.
-Despite their simplicity, the low-rank models are already able to generate an approximate version of the dataset, as shown in Fig.
-1–D-F for the 4 datasets previously presented.
-In the initial stage of the standard learning process, the model encodes the strongest PCA components of the data through multiple critical transitions [26, 27, 28].
-Pre-training with the low-rank construction allows us to bypass these transitions and avoid out-of-equilibrium effects caused by critical slowing down associated to these transitions.
-Once the main directions are incorporated, training can efficiently continue with standard algorithms like PCD, as the mixing times of pre-trained machines tend to be much shorter.
-In particular, in the PCD-100 training with MNIST01, relaxation times for the visible variables’ time correlation reach $5\cdot 10^{5}$ MCMC steps at the first three transitions, coinciding with the growth of singular values in the model weight matrix $W$.
-In contrast, the pre-trained machine has a much shorter relaxation time of $\sim 10^{3}$, allowing us to safely restart the PCD process from a set of equilibrium samples generated by static sampling of the low-rank RBM.
-Figure 2: We compare the equilibrium samples generated by RBMs trained on the Mickey, MNIST01, and HGD datasets using three different training schemes: Jarzynski (JarRBM), PCD, and PCD initialized on low-rank RBMs (used to generate the samples in Fig.
-1–D-F).
-To assess the fitting of the modes, we show a density plot of the projections of the data in the first two principal directions of each dataset.
-We compare these results with the density plot of the original datasets in the first column.
-Overcoming these transitions has dramatic implications for the quality of the models we can train and how accurately they reproduce the statistics of the data.
+where $\{w_{\alpha}\}$ are the singular values of the coupling matrix and the right singular vectors $\{\bm{u}_{\alpha}\}_{\alpha=1}^{d}$ correspond exactly to the first $d$ principal directions of the data set.
+Under this assumption, $p(\bm{v})$ a function of only $d$ order parameters, given by the magnetizations along each of the $\boldsymbol{u}_{\alpha}$ components, i.e.
+$m_{\alpha}(\bm{v})=\bm{u}_{\alpha}\cdot\bm{v}/\sqrt{{N_{\mathrm{v}}}}$.
+In particular,
+$H(\bm{v})=-\sum_{\alpha=0}^{d}\theta_{\alpha}m_{\alpha}-\sum_{a}\log\cosh\left(\sqrt{N_{\mathrm{v}}}\bar{u}_{a}\sum_{\alpha=1}^{d}w_{\alpha}m_{\alpha}+\eta_{a}\right)=\mathcal{H}(\bm{m}(\bm{v})),$ (4)
+where $\bm{m}=\left(m_{1},\ldots,m_{\alpha}\right)$ and $\theta_{\alpha}$ is the projection of the visible bias on the extended $\bm{u}$ matrix.
+In order to initialize the visible bias as suggested in Montavon & Müller (2012), a direction $\bm{u}_{0}$ is used for the magnetization $m_{0}$ that is only present in the bias term.
+This is an improvement of this work with respect to Decelle & Furtlehner (2021a), as it allows us to gain almost for free one additional direction.
+The optimal parameters can be determined by solving a convex optimization problem, detailed in SI A.2.
+Once the model is trained, we obtain a probability distribution $p(\bm{m})$ in a much lower-dimensional space than the original $p(\bm{v})$, which can be easily sampled globally using the inverse transform sampling method.
+However, since this approach requires discretizing the $\bm{m}$-space for both training and generation, it becomes impractical for intrinsic space dimensions $d>4$ due to memory constraints.
+These low-rank RBMs are trained to replicate the statistics of the dataset projected onto its first $d$ principal components.
+Despite their simplicity, the low-rank models can still generate an approximate version of the dataset, as shown in Supplemental Figure 15 for the five datasets previously presented.
+In the initial stage of the standard RBM learning process, the model encodes the strongest PCA components of the data through multiple critical transitions (Decelle et al., 2017; 2018; Bachtis et al., 2024).
+Pre-training with the low-rank construction allows us to bypass these transitions and avoid out-of-equilibrium effects caused by critical slowing down associated to these transitions, which becomes prohibitive in the case of very clustered datasets.
+Once the main directions are incorporated, training can efficiently continue with standard algorithms like PCD, as the mixing times of pre-trained machines tend to be much shorter that at the transitions.
+Figure 2: We compare the equilibrium samples generated by RBMs trained on the MNIST01, Mickey, HGD, Ising2D and CelebA datasets using two different training schemes: PCD, and PCD initialized on low-rank RBMs.
+To assess the fitting of the modes, we show a scatter plot of the projections of the samples in the first two principal directions of each dataset, together with the histograms along each direction.
+Black dots and histograms refer to the original data distribution, while the red color is used to represent the generated data.
+In D we show the evolution of the log-likelihood in time for the two training methods computed with the online AIS trajectory.
+In continuous lines we show the train and in dashed, the test.
+Results are obtained by averaging with respect to 10 trainings initialized at the same low-rank RBM, and the standard deviation between all is shown in a shade around the line, which is hardly seen due to the small variation from run to run.
+Overcoming these transitions has dramatic implications for the quality of the models we can train and how accurately they reproduce the statistics of the data, especially for short training times.
 In Fig.
-2, we show for 3 datasets the equilibrium samples drawn from 3 RBMs trained with identical number of samples, minibatch size, $k=100$ Gibbs steps, and learning rate $\gamma=0.01$, but different training strategies.
-In particular, we consider 2 RBMs trained from scratch with the standard PCD [33] and the recently proposed Jarzynski reweighing method [23] (see SI for our specific implementation in the RBM), and a final machine trained with PCD and pre-trained with a low-rank RBM.
-In all cases, the quality of the generated samples is significantly better when pre-training is used.
-For the Mickey dataset, neither JarRBM nor normal PCD are able to generate convincing data.
-For the MNIST01 dataset, all 3 methods are able to generate convincing data, but only Pretrain+PCD is able to correctly balance all modes, as can be seen in Fig.
-3, where we compare the histograms of the generated data projected onto the first 3 PCA directions with those of the dataset and a random selection of the generated samples.
-We see that the pre-training+PCD training perfectly balances the different modes (here we show the first 3 directions, but it goes much further), unlike the other 2 methods, and also generates more diverse images.
-We can also compare the log-likelihood of all 3 models and find that the pre-trained RBM achieves higher values.
-At this point, it is important to emphasize that in order to properly quantify the increase in log-likelihood, we need to use the PTT algorithm (see section 5) to correctly thermalize in these well-trained machines.
-For comparison, we show our PTT measure in dark and solid lines, while the standard AIS [42] estimate is shown in light dashed lines.
-Figure 3: We compare the samples generated by the 3 RBMs (JarRBM, PCD, pretrain+PCD) trained with MNIST01 data.
-In A, we show the histograms of the generated data projected on the first, second and third principal directions with those of the dataset.
-We see that only the pretrain+PCD correctly balances the different modes.
-In B we show 10 images generated by each machine.
-In C, we compare the log-likelihood of each model’s dataset as a function of training time.
-The dark and full curves were obtained using the PTT algorithm discussed in section  5, and the lighter and dashed curves using the AIS method [42].
-Already from the scatter plots we see that the pre-training has a dramatic effect in obtaining models where all modes are properly balanced, but also has important effects in the maximum test-likelihood we can achieve.
-In all cases, these equilibrium samples are drawn using the trajectory PT algorithm that will be explained in the next section, and the log-likelihood obtained using the equilibrium configurations obtained at different epochs as a result of the trajectory PT flow.
-5 Standard Gibbs sampling vs. Parallel Trajectory Tempering (PTT)
+2, we show for 5 datasets the equilibrium samples drawn from 2 RBMs trained with identical number of samples, minibatch size, $k=100$ Gibbs steps, and learning rate $\gamma=0.01$, but different training strategies.
+In particular, we consider an RBMs trained from scratch with the standard PCD (Tieleman, 2008), and a final machine trained with PCD and pre-trained with a low-rank RBM.
+We also considered machines trained with the recently proposed Jarzynski reweighing method (Carbone et al., 2023) (see SI E for our specific implementation in the RBM), but we do not include the analysis here because in most of the datasets the training method eventually becomes very unstable reaching very poor results.
+In all datasets, the quality of the generated samples and the test log-likelihood are significantly better when pre-training is used.
+We see that the pre-training+PCD training perfectly balances the different modes (here we show the projections on the first 2 PCA directions, but the agreement extends further to several other principal directions).
+We can also compare the log-likelihood of all 5 datasets and find that the pre-trained RBM achieves higher values.
+In all cases, the log-likelihood is measured using the annealed importance sampling trajectory (Tr-AIS) method, and the samples are drawn using the trajectory PT (PTT) algorithm.
+Both algorithms will be described in the next section 5.
+Already from the scatter plots, we see that the pre-training is key in obtaining models where all modes are properly balanced in clustered datasets, but it’s also beneficial in the maximum test log-likelihood we can achieve.
+We observe that the pretraining is not particularly useful when the dataset is not much clustered, as it is the case of CelebA and the entire MNIST dataset as shown in Fig.
+16 of the SI I.
+This result can be easily understood as the encoding of the first modes of the data do not involve splitting the distribution into two isolated modes, something that happens in the case of clustered data.
+5 Trajectory annealing
+The evolution of parameters during training can be seen as a smoother annealing process compared to the standard temperature annealing used in algorithms like annealing importance sampling (AIS) for log-likelihood estimation or the PT algorithm for sampling.
+This is because it avoids the first-order transitions that typically occur with temperature in clustered datasets.
+However, these methods can still be applied to trajectory annealing by using the parameters from different stages of training, rather than scaling the model parameters by a common factor $\beta_{i}$, as is done in temperature-based protocols.
+5.1 Trajectory Annealing importance sampling
+The standard method for computing the log-likelihood (LL) in the literature is Annealed Importance Sampling (AIS) (Neal, 2001).
+For a detailed review of its application in RBMs, see Krause et al.
+(2020).
+In this approach, the partition function of a model with parameters $\bm{\Theta}=\{\bm{w},\bm{\theta},\bm{\eta}\}$ is estimated by gradually increasing the inverse temperature $\beta$ from 0 to 1 over $N_{\beta}$ steps.
+At $\beta=0$ (where $p_{0}=p_{\mathrm{ref}}$), the partition function $Z_{0}$ is tractable, enabling the estimation of $Z_{i}$ at neighboring $\beta_{i}$ values through path sampling methods.
+The reference configuration $p_{\mathrm{ref}}$ is commonly taken uniform, but improvements have been reported in RBM when fixing it to the independent site distribution that describes the empirical center of the dataset.
+For a detailed description of the algorithm, see SI D.
+AIS is quite costly because it requires a large number of annealing steps $N_{\beta}$ for reliable estimations, making it impractical to compute the log-likelihood online during training.
+As a result, it is often calculated offline using a set of models at different temperatures.
+However, since the training process itself follows an annealing trajectory, it makes more sense to use the evolution of parameters during training as annealing steps rather than relying on temperature.
+Additionally, at the start of training, parameters are set to zero, making $Z_{0}$ tractable, which also holds when initializing with a low-rank RBM.
+Since updating the partition function only involves calculating the energy difference between the new and old parameters, this can be efficiently computed online during each parameter update, allowing for very small integration steps.
+We refer to this method as online Tr-AIS.
+Alternatively, similar to standard AIS, multiple models can be saved during training for an offline Tr-AIS estimation.
+To evaluate the reliability of our new approach, we trained several RBMs using datasets with a maximum of 20 hidden nodes, allowing for the exact computation of the partition function through direct enumeration.
+We then calculated the difference between the true partition function and the estimations obtained from various methods.
+The results for the HGD dataset are presented in Fig.
+3, while similar qualitative results for the other datasets can be found in the SI, see Figs.
+11, 12, 13 and 14.
+In Fig.
+3, we display the estimation errors in the LL using three different methods: standard temperature Annealed Importance Sampling (AIS) (panel A), AIS with a reference distribution (panel B), and Tr-AIS (panel C).
+For the temperature AIS estimates, we show results for different values of $N_{\beta}$ in distinct colors.
+In panel C, we compare the online estimation in black with the offline Tr-AIS estimation, using the same models analyzed in panels A and B.
+Additionally, we present the error in LL estimation using a few machines and the trajectory parallel tempering (PTT), which will be detailed in the next section.
+Overall, the online Tr-AIS estimation consistently outperforms the other methods, and only the PTT estimation yields similar results in offline analyses.
+Figure 3: Comparison of the error in (LL) estimation (with respect to the exact value) using different methods: AIS (left), AIS with a reference distribution (middle), and Tr-AIS (ours, right).
+For Tr-AIS, we present three approaches to computing it: online during training, offline using the models saved during training, and with PTT.
+The RBM was pretrained and trained with a total of 20 hidden nodes (allowing for an exact computation of the LL) for 10,000 gradient steps.
+PTT uses a subset of saved models such that the acceptance rate between two consecutive models is 0.25.
+5.2 Standard Gibbs sampling vs. Parallel Trajectory Tempering
 One major challenge with structured datasets is quantifying the model’s quality, since sampling the equilibrium measure of a well-trained model is often too time-consuming.
-This affects the reliability of generated samples and indirect measures as log-likelihood’s estimation through Annealing Importance Sampling (AIS) [42], making them inaccurate and meaningless.
-Figure 4: Comparison between PTT and classical Gibbs sampling for the MNIST01 dataset (A and B, respectively) and the human genome dataset (C and D, respectively).
-In A and C, we show the trajectory of two independent chains (red and orange) projected onto the PCA along the sampling process of the pretraining+PCD model for $10^{4}$ MCMC steps.
-The black contour represents the density profile of the dataset and the position of the chains is plotted every 10 steps.
-In B and D we show the average number of jumps from one cluster to another as a function of the MCMC steps performed.
-The average is calculated over a population of 100 chains.
-In D, we show the average jump time between clusters along the first (solid line) and second (dashed line) principal components of the data.
-To illustrate this problem, let us consider the MNIST01 and the HGD datasets.
-MNIST01 dataset is bimodal and the HGD highly multimodal as shown in their PCA in Figs.
-1–A and C.
+This affects the reliability of generated samples and indirect measures as log-likelihood’s estimation through Annealing Importance Sampling (AIS) (Krause et al., 2020), making them inaccurate and meaningless.
+Figure 4: Comparison of the performance of different MCMC sampling methods on RBMs trained with the pretraining+PCD procedure on the MNIST01 (row-1), the HGD (row-2) and Ising 2D datasets (row-3).
+In A and B columns, we show the trajectory of two independent Markov chains (red and orange) iterated for $10^{4}$ MCMC steps using either PTT or alternate Gibbs sampling (AGS), projected onto the first two principal components of the dataset.
+The position of the chains is plotted every 10 MCMC steps.
+The black contour represents the density profile of the dataset.
+In column C, we show the average number of jumps between the two regions separated by the dashed grew line in the PCA plots using different MCMC methods: Alternative Gibbs Sampling (AGS), Parallel Tempering (PT) (Hukushima & Nemoto, 1996), without or with reference configuration (PTref) (Krause et al., 2020), the Stacked Tempering (Roussel et al., 2023), and the Trajectory Parallel Tempering (PTT) proposed in this work.
+The average is calculated over a population of 1000 chains and the shadow around the lines indicates the error of the mean.
+To illustrate this problem, let us consider the MNIST01, the HGD and the Ising 2D datasets.
+MNIST01 and Ising 2D are strongly bimodal in the PCA and the HGD trimodal, as shown in Fig.
+1.
 Let us consider that we want to sample the equilibrium measure of the RBMs trained using low-rank RBM pretraining.
 In order to draw new samples from these models, one would typically run MCMC processes from random initialization and iterate them until convergence.
-The mixing time is controlled by the jumping time between clusters.
+The mixing time in such case is controlled by the jumping time between clusters.
 To accurately estimate the relative weight between modes, the MCMC processes must be ergodic, requiring many back-and-forth jumps.
 However, as shown in Figs.
-4–A and C for the MNIST01 and HGD datasets, Gibbs sampling dynamics are extremely slow, rarely producing jumps even after $10^{4}$ MCMC steps.
-The yellow curves in Figs.
-4–B and D show the mean number of jumps over 100 independent chains as a function of MCMC steps, indicating that a proper equilibrium generation would require at least $10^{6}-10^{7}$ MCMC steps.
-One effective way to accelerate the dynamics is to exploit the training trajectory, where the model progressively specializes through second-order phase transitions.
-To achieve this, we save RBMs trained at various epochs and propose swaps between configurations of similarly trained models.
-We call this the Parallel Trajectory Tempering (PTT) algorithm.
-Unlike the standard Parallel Tempering (PT) algorithm, which attempts swaps configurations between different temperatures, the PTT swaps between model parameters with different degrees of specialization.
-This approach is more natural for this problem because it involves crossing only second-order transitions, unlike the first-order transitions occurring in temperature annealing.
-And in fact, we show in Figs.
-4–A and C, that this approach allows us to sharply accelerate the dynamics, as opposed to the standard PT algorithm (studied in detail for the MNIST dataset in [40]).
-In the PTT algorithm, the configurations $\bm{x}=(\bm{v},\bm{h})$ of neighboring machines indexed by $t$ and $t-1$ are interchanged with the probability
-$p_{\mathrm{acc}}(\bm{x}^{t}\leftrightarrow\bm{x}^{t-1})=\min\left(1,\exp\left(\Delta\mathcal{H}^{t}(\bm{x}^{t})-\Delta\mathcal{H}^{t}(\bm{x}^{t-1})\right)\right).$
+4–A1-3, Alternate Gibbs sampling (AGS) dynamics are extremely slow, rarely producing jumps even after $10^{4}$ MCMC steps.
+The red curves in Figs.
+4–C1-3 show the mean number of jumps over 1000 independent chains as a function of MCMC steps, indicating that a proper equilibrium generation would require at least $10^{6}-10^{7}$ MCMC steps.
+In order to speed up sampling, we consider other traditional optimized MCMC methods such as the PT (with or without reference), the Stack Tempering (Roussel et al., 2023) and our Parallel Trajectory Tempering (PTT) that will be explained below.
+The evolution of the number of jumps between modes observed with each of these algorithms, as a function of the total number of AGS steps performed (i.e.
+the number of parallel models multiplied by the number of AGS steps performed at each model) is shown in Fig.
+4–C1-3.
+Before discussing our PTT algorithm and the results, let us briefly introduce the PT algorithm.
+5.2.1 Parallel Tempering approaches
+Parallel tempering (PT) is commonly used to simulate multiple systems in parallel at different inverse temperatures $\beta$ (Hukushima & Nemoto, 1996; Marinari & Parisi, 1992), with occasional temperature swaps between models based on the Metropolis rule.
+Since this move satisfies detailed balance, the dynamics will converge to the Boltzmann distribution $p_{\beta}(\bm{x})\propto\exp(-\beta\mathcal{H}(\bm{x}))$ at each $\beta$.
+However, a more general approach involves simulating multiple systems, each with its own set of parameters.
+We can define a system as $p_{\bm{\Theta}^{t}}(\bm{x})$, where $\bm{\Theta}^{t}$ denotes the parameters of the model indexed by $t$.
+This includes the classic PT case, where temperature scaling is applied, $\bm{\Theta}^{t}=\beta^{(t)}\bm{\Theta}$, with $\bm{\Theta}$ being the parameters of the model to be sampled.
+In this work, we introduce a new algorithm that exchanges parameters of models saved at different stages of learning, using the learning trajectory to define the parameter set $\bm{\Theta}^{t}$, where $t$ refers to a specific point in the training process.
+We call this method Parallel Trajectory Tempering (PTT).
+The rule for swapping configurations $\bm{x}=(\bm{v},\bm{h})$ between neighboring machines indexed by $t$ and $t-1$ is given by
+$p_{\mathrm{acc}}(\bm{x}^{t}\leftrightarrow\bm{x}^{t-1})=\min\left(1,\exp\left(\Delta\mathcal{H}^{t}(\bm{x}^{t})-\Delta\mathcal{H}^{t}(\bm{x}^{t-1})\right)\right),$ (5)
+where again the index $t$ indicates that the Hamiltonian is evaluated using the parameters $\bm{\Theta}^{t}$.
 This move satisfies detailed balance with our target equilibrium distribution $p(\bm{x})=\exp(-\mathcal{H}(\bm{x}))/Z$, ensuring that the moves lead to the same equilibrium measure.
-As “nonspecialized" models mix very quickly, either because the distribution is essentially Gaussian at the initialisation of a standard training, or because the low-rank RBM can be sampled with a static Monte Carlo process (yielding independent configurations each time), the trajectory flow significantly accelerates convergence to equilibrium.
-The time interval between successive machines is selected in such a way that the probability of accepting interchanges between neighboring machines remains around 0.3.
-Pre-trained machines require a significant fewer number of models to be effective, because most selected models are positioned at the most prominent phase transitions.
-We give the number of machines used for each sampling process in the SI.
-We also provide there a specific and detailed description of the algorithm used.
-In the red curves in Fig.
-(4)–B and D, we show the number of jumps between clusters as a function of the number of elementary MCMC steps, which in the PTT scheme refer to 1 Gibbs sampling step + one swap proposal.
-For the DNA dataset, we have two measures corresponding to jumps along the two principal component directions.
-We observe at $10^{4}$ MCMC steps an increase of the number of jumps by a factor of 80 for MNIST01 and by a factor of 1350 for the HGD in this machine, although we achieve higher factors in other machines, as we show in the SI.
-The sampling of RBMs training on the MNIST01 dataset was the subject of the study of the “stacked tempering" algorithm in  [40].
-If we compare the numbers with their work, we see that we achieve a 3-4 times higher speedup factor, where our model has the advantage that it does not need additional training, but simply uses the stored machines correctly.
-Another desirable advantage of our PTT algorithm is that we can easily use it to compute an improved estimate of the AIS log-likelihood, except that in our case we consider the training trajectory instead of a cooling process and use the equilibrium samples obtained for each of the models to compute the model averages.
-In Figs.
-3–C  5–A we compare the log-likelihood estimates obtained with our method (AIS-PTT) in full and dark lines and in light and dashed lines the AIS estimate (AIS).
-We see that both measures coincide for most parts of the training and that they split when the sampling becomes too long to thermalize along the temperature annealing curve in AIS.
-This effect is particularly evident for the JarRBM run in  5–A, where AIS takes a long time to recognize that the model suffers from a strong mode-collapse effect.
-6 Overfitting and privacy loss as quality indicators
-In this section, we examine the quality of the samples generated, regarding overfitting and privacy criteria which have been defined for genomic data in particular.
-We look at this on the models trained with PCD with and without pre-training.
-We do not include the Jarzysnki method here, as this method fails to obtain a reliable model as clearly shown in the evolution of the Log-likelihood in Fig.
-5.
-We focus on the human genome dataset, as shown in Fig.
-1–C, to evaluate the ability of various state-of-the-art generative models to generate realistic fake genomes while minimizing privacy concerns (i.e., reducing overfitting).
-Recent studies [17, 18] have thoroughly investigated this for a variety of generative models.
-Both studies concluded that the RBM was the most accurate method for generating high-quality and private synthetic genomes.
-The comparison between models relies primarily on the Nearest Neighbor Adversarial Accuracy ($AA_{\rm TS}$) and privacy loss indicators, introduced in Ref.
-[43], which quantify the similarity and the level of "privacy" of the data generated by a model w.r.t.
-the training set.
-We have $AA_{\rm TS}=\frac{1}{2}\bigl{(}AA_{\rm True}+AA_{\rm Synth}\bigr{)}$ where $AA_{\rm True}$ [resp.
-$AA_{\rm Synth}$] are two quantities in $[0,1]$ obtained by merging two sets of real and synthetic data of equal size $N_{s}$ and measuring respectively the frequency that a real [rep.
-synthetic] has a synthetic [resp.
-real] as nearest neighbor.
-If the generated samples are statistically indistinguishable from real samples, both frequencies $AA_{\rm True}$ and $AA_{\rm Synth}$ should converge to 0.5 at large ${N_{\mathrm{s}}}$.
-$AA_{\rm TS}$ can be evaluated both with train or test samples and the privacy loss indicator is defined as ${\rm Privacy\ loss}=AA_{\rm TS}^{\rm test}-AA_{\rm TS}^{\rm train}$ and is expected to be strictly positive.
-Fig.
-5 shows the comparison of $AA_{\rm TS}$ and privacy loss values obtained with our two models, demonstrating that the pre-trained RBM clearly outperforms the other model, and even achieves better results ($AA_{\rm TS}$ values much closer to 0.5) than those discussed in [17, 18].
-Figure 5: We compare the quality of the RBMs trained with the human genome data (HGD).
-In A, we show the log-likelihood as a function of the training epochs for the 3 training procedures.
-Solid lines correspond to AIS-PTT and dashed lines to AIS.
-The JarRBM falls down because the training breaks eventually.
-In B and C we compare privacy and overfitting based on the $AA_{\rm TS}$ indicator.
+The reason for using the trajectory is that PT algorithms work very well when a system undergoes a continuous phase transition.
+As discussed in the appendix G, adjusting the inverse temperature can lead to discontinuous transitions for which PT is not known to perform very well, while following the trajectory, the system undergoes several continuous transitions as discussed in Decelle et al.
+(2018); Bachtis et al.
+(2024).
+Our PTT approach relies on the ease of sampling configurations from an initial RBM.
+In classical PCD training, the early “non-specialized” models thermalize quickly because they behave like Gaussian distributions.
+For RCM+PCD, independent configurations can be efficiently sampled from the low-rank RBM using a static Monte Carlo algorithm, as described in Section A.1 of the SI.
+The trajectory flow then significantly speeds up convergence to equilibrium, especially in the presence of a multimodal structure.
+The time interval between successive models is chosen to keep the acceptance probability for exchanges between neighboring models (Eq.
+5) around 0.25 (we observed that the total time (number of AGS multiplied by number of machines) needed to thermalize a PTT chain did not depend on the acceptance fixed in the interval [0.2,0.5].
+In practice, effective models can be saved during training by monitoring the acceptance rate between the last saved model and the current one.
+Pre-trained models require far fewer saved models, as many are naturally positioned near major phase transitions.
+The number of models used for each sampling process, along with a detailed description of the algorithm, is provided in the SI.
+In Fig.
+4–C1-3, we present the number of jumps between clusters as a function of the total number of AGS steps, where in PT-like schemes this refers to 1 (AGS step + one swap proposal) multiplied by the total number of models run in parallel.
+In all cases, PTT sampling outperforms the other methods, showing particularly remarkable results for HGD.
+Additionally, our method has the advantage of not requiring extra training like Stacked Tempering or simulating unnecessary temperatures.
+Moreover, the equilibrium samples obtained from different models during sampling can be used to compute the log-likelihood, similar to the standard PT approach Krause et al.
+(2020).
+In Fig.
+3–right we show that the quality of this estimation is comparable with the online trajectory AIS.
+Another desirable advantage of the PT algorithms is that one can easily check the thermalisation of the MCMC process by investigating the ergodicity of the random walk in models Banos et al.
+(2010).
+In addition, we evaluated other indicators of sample quality and potential overfitting.
+As detailed analysis is in Sect.
+H of the SI.
+Our method consistently provides high-quality samples across all datasets and, thanks to precise likelihood computation, can effectively detect overfitting.
 6 Conclusions
-We have shown that the strategy of initiating the training on a pre-trained low-rank RBM is an extremely effective strategy to obtain high quality models for structured datasets that accurately represent all the modes in the datasets and with significantly higher log-likelihoods.
-We have also shown that the models obtained in that way are: (i) better generative models than those obtained with standard trainings, both, in the sense that they over-fit less at the same time they are more indistinguishable from the test samples, (ii) they display faster relaxational dynamics.
-We have also proposed a new fast sampling method that exploits the progressive learning of features in the training of RBMs to design an efficient trajectory PT strategy that allows accelerating the parallel Gibbs sampling dynamics by many orders of magnitude and overcome the performance of recent efficient sampling methods without adding any extra cost than saving models during the training.
-Both strategies for training and sampling are very general, and could be generalized to more complex EBMs.
-In this sense, the low-rank RBM model could be used as a more efficient pre-initialisation in deeper structures, and the trajectory PT algorithm is suitable to be directly used in any EBM no matter how complex it is.
+In this work, we tackle the challenges of training high-quality equilibrium RBMs on highly structured datasets.
+We focused on improving the sampling process and accurately evaluating test log-likelihood throughout training.
+Our findings reveal that for datasets that are tightly clustered in their principal component analysis representation, introducing a pre-training phase can enhance training quality.
+This phase involves encoding the principal directions of the data into the model using a convex optimization approach.
+However, this pre-training method, that is specific to RBMs, appears to have limited benefits for less structured datasets or for long trainings, as chains get out of equilibrium, and should be combined in the future with optimized MCMC methods to achieve its maximum capacity.
+Furthermore, we developed a new approach that exploits the progressive learning of features during RBM training to enhance evaluation and sampling.
+This approach enables online and precise computation of log-likelihood and leads to an efficient sampling strategy akin to the standard parallel tempering algorithm.
+Our results demonstrate that both our proposed methods outperform traditional algorithms without the need for simulating models at unnecessary temperatures or training additional models to facilitate sampling.
+Furthermore, both the trajectory AIS method for log-likelihood estimation and the trajectory parallel tempering algorithm are highly versatile, making them applicable to more complex energy-based models.
+Additionally, the low-rank RBM model can serve as an effective pre-initialization step for deeper architectures, and the trajectory PT algorithm is suitable for direct use in any EBM, regardless of complexity.
 7 Code availability
-The code and datasets are available at https://github.com/nbereux/fast-RBM.
-9 acknowledgments
-The authors acknowledge financial support by the Comunidad de Madrid and the Complutense University of Madrid (UCM) through the Atracción de Talento program (Refs.
-2019-T1/TIC-13298), and to project PID2021-125506NA-I00 financed by the Ministerio de Economía y Competitividad, Agencia Estatal de Investigación (MICIU/AEI /10.13039/501100011033), and the Fondo Europeo de Desarrollo Regional (FEDER, UE).
+The code and datasets are available at https://github.com/anonymousrbm/iclr-2025
+Supplemental Information
 Appendix A Details of the pre-training of a low rank RBM
@@
 In Ref.
-[41], it has been shown that such RBM can be trained by mean of the Restricted Coulomb Machine, where the gradient is actually convex in the parameter’s space.
+Decelle & Furtlehner (2021a), it has been shown that such RBM can be trained by mean of the Restricted Coulomb Machine, where the gradient is actually convex in the parameter’s space.
 It is then possible to do a mapping from the RCM to the RBM to recover the RBM’s parameters.
@@
 A.2 The Restricted Coulomb Machine
-As introduced in [41], it is possible to exactly train a surrogate model for the RBM, called the Restricted Coulomb Machine (RCM), on a low dimensional dataset without explicitly sampling the machine allowing to learn even heavily clustered datasets.
-We will briefly outline the main steps to train the RCM.
-A more detailed explanation can be found in Appendix A.2.
+As introduced in Decelle & Furtlehner (2021a), it is possible to exactly train a surrogate model for the RBM, called the Restricted Coulomb Machine (RCM), on a low dimensional dataset without explicitly sampling the machine allowing to learn even heavily clustered datasets.
 The RCM is an approximation of the marginal distribution of the RBM with $\{-1,1\}$ binary variables:
@@
 We then project both the parameters and variables of the RBM on the first $d$ principal components of the dataset:
-$m_{\alpha}\coloneqq\frac{1}{\sqrt{N_{v}}}\sum_{i=1}^{N_{v}}s_{i}u_{i\alpha},\quad w_{\alpha a}\coloneqq\sum_{i=1}^{N_{v}}w_{ia}u_{i\alpha},\quad\theta_{\alpha}\coloneqq\frac{1}{\sqrt{N_{v}}}\sum_{i=1}^{N_{v}}\theta_{i}u_{i\alpha}$ (19)
-with $\alpha\in\{1,\ldots,d\}$ and $\bm{v}$ the projection matrix of the PCA. The projected distribution of the model is then given by
+$m_{\alpha}\coloneqq\frac{1}{\sqrt{N_{v}}}\sum_{i=1}^{N_{v}}v_{i}u_{i\alpha},\quad w_{\alpha a}\coloneqq\sum_{i=1}^{N_{v}}w_{ia}u_{i\alpha},\quad\theta_{\alpha}\coloneqq\frac{1}{\sqrt{N_{v}}}\sum_{i=1}^{N_{v}}\theta_{i}u_{i\alpha}$ (20)
+with $\alpha\in\{0,\ldots,d\}$ and $\bm{u}$ the projection matrix of the PCA.
+The projected distribution of the model is then given by
 $p_{\text{RBM}}(\bm{m})=\frac{\exp\left(N_{v}\left[\mathcal{S}(\bm{m})+\sum_{\alpha=1}^{d}\theta_{\alpha}m_{\alpha}+\frac{1}{N_{v}}\sum_{a=1}^{N_{h}}\log\cosh\left(\sqrt{N_{v}}\sum_{\alpha=1}^{d}m_{\alpha}w_{\alpha a}+\eta_{a}\right)\right]\right)}{Z}$ (21)
@@
 To relax the problem, since we’re in low dimension, we can define a family of hyperplanes $(\bm{n},\bm{z})$ covering the space and let the model only learn the weights of each to the hyperplane.
-We can then discard the ones with a weight low enough for the approximation (21) to be bad.
+We can then discard the ones with a weight low enough for the approximation equation 22 to be bad.
 The gradients are given by
@@
 The positive term is straightforward to compute. For the negative term, we rely on a discretization of the longitudinal space to estimate the probability density of the model and compute the averages.
+A.3 Learning the bias
+To initialize the model with its bias set to the dataset’s mean, we define the 00-th direction of the longitudinal space as the direction obtained by projecting the configuration space onto the normalized mean vector of the dataset.
+The RBM will only contribute to this direction through the visible bias:
+$H(\bm{m}(v))=-\sum_{\alpha=0}^{N_{v}}\theta_{\alpha}m_{\alpha}-\frac{1}{N_{v}}\sum_{a}\log\cosh\left(\sqrt{N_{v}}\sum_{\alpha=1}^{d}m_{\alpha}w_{\alpha a}+\eta_{a}\right).$ (27)
+This implies that this direction is independent of the others within the model.
+Consequently, we can create an independent mesh for this direction to learn it.
+The rest of the training is performed using the same procedure as before.
 Appendix B Sampling via Parallel Tempering using the learning trajectory
@@
 For pre-train+PCD, at each step we sample a new independent configuration for $\mathrm{RBM}_{0}$ using the RCM.
-Model Dataset # of machines Alg. # of steps acc. factor @ 104 steps
-pre-train+PCD MNIST01 6 (+1) 10000 80
-JarJar MNIST01 28 10000 50
-PCD MNIST01 13 10000 30
-pre-train+PCD Human Genome 6 (+1) 10000 1350
-PCD Human Genome 13 10000 7100
-Table 1: Performance comparison of different models on various datasets for the sampling using PTT versus Gibbs sampling for $10^{4}$ mcmc steps.
-The acceleration factor is defined as the ratio of the average number of jumps obtained until $10^{4}$ steps between PTT and Gibbs sampling.
-For pre-train+PCD, the RCM machine has not to be counted among the list of models (hence the +1) because it is very fast to sample from.
 Assuming we have successfully trained a robust equilibrium model, there remains the challenge of efficiently generating equilibrium configurations from this model.
@@
 A sketch of the Parallel Trajectory Tempering (PTT) is represented in fig.
-6.
+5.
 Specifically, we save $t_{\mathrm{f}}$ models at checkpoints $t=1,\dots,t_{\mathrm{f}}$ along the training trajectory.
@@
 end for
-A comparison of performances between PTT and standard Gibbs sampling is reported in Tab. 1.
-Figure 7: Comparison between PTT and standard Gibbs Sampling for RBMs trained using PCD (A and B) and JarRBM (C and D) on the MNIST01 dataset.
-A and C show the sampling trajectory of two chains recorded every 10 steps for a total of $10^{4}$ mcmc steps.
-B and D show the average number of jumps of a population of 100 chains as a function of the sampling time.
-Figure 8: Comparison between PTT and standard Gibbs Sampling for RBMs trained using PCD on the Human Genome dataset. The sampling has been performed under the same conditions of fig. 7.
+A comparison of performances between PTT and standard Gibbs sampling is reported in Fig.4.
+B.1 Controlling equilibration times with PTT
+One of the advantages of PT-like approaches is that they allow easy control over thermalization times by studying the random walk in temperatures (or models for the PTT) visited by each run over time, as proposed in Ref.
+Banos et al.
+(2010).
+From now on, we will refer only to the number of models used in the PTT sampling, denoted as $N_{\mathrm{m}}$.
+In an ergodic sampling process, each run should spend roughly the same amount of time at each of the $N_{\mathrm{m}}$ different model indices, as all models have the same probability.
+The mean temperature/model index is known to be $\left\langle n\right\rangle=(N_{\mathrm{m}}-1)/2$ (assuming the model indices run from 00 to $N_{\mathrm{m}}-1$).
+This simplifies the computation of the time-autocorrelation function of the model indices $n_{i}(t)$ visited by each sample $i$ at time $t$:
+$C(t)=\frac{\overline{(n_{i}(t+t_{0})-\left\langle n\right\rangle)(n_{i}(t_{0})-\left\langle n\right\rangle)}}{\overline{(n_{i}(t_{0})-\left\langle n\right\rangle)^{2}}}$
+Here, the symbol $\overline{(\cdots)}$ refers to the average over all samples $i$ and initial times $t_{0}$ obtained from our PTT runs.
+This allows us to compute the exponential autocorrelation time, $\tau_{\mathrm{exp}}$, from a fit to
+$C(t)\sim Ae^{-t/\tau_{\mathrm{exp}}}\quad\text{for large }t,$
+as well as the integrated autocorrelation time, $\tau_{\mathrm{int}}$, using the self-consistent equation Sokal (1997); Amit & Martin-Mayor (2005):
+$\tau_{\mathrm{int}}=\frac{1}{2}+\sum_{t=0}^{6\tau_{\mathrm{int}}}C(t).$
+These two times provide different insights: $\tau_{\mathrm{exp}}$ gives insight into the time needed to thermalize the simulations (setting the length of the runs to at least $20\tau_{\mathrm{exp}}$ is generally a safe (and very conservative) choice to ensure thermalization Sokal (1997)), while $\tau_{\mathrm{int}}$ gives the time needed to generate independent samples.
+To keep track of the fluctuations, we compute a different $C^{m}(t)$ for each $m=1,\ldots,N_{\mathrm{m}}$ system index simulated in parallel in the PTT, where each $C^{m}(t)$ is averaged over 1000 independent realizations (with all runs initialized to the same model index).
+We then compute the average over all parallel runs $C(t)=\overline{C^{m}(t)}$ and use this curve to extract $\tau_{\mathrm{exp}}$ and $\tau_{\mathrm{int}}$.
+The errors in $C$ and the times are then extracted from the error of the mean of the values extracted with each of the $C^{m}(t)$ curves.
+For example, in Fig.
+6 (left), we show the averaged curves $C(t)$ obtained with the PTT sampling of the machines discussed in the main text in the MNIST01 (top) and HGD (bottom) datasets.
+The extracted times are $\tau_{\mathrm{exp}}=258(29)$ and $\tau_{\mathrm{int}}=9.0(1.5)$ for MNIST01, and $\tau_{\mathrm{exp}}=48(10)$ and $\tau_{\mathrm{int}}=3.8(7)$ for HGD.
+These numbers guarantee us that the $10^{4}$ steps used to generate the scatterplots in Fig.
+2 were long enough to ensure equilibration in this dataset, but also in the others, as shown in Table 1.
+Figure 6: (Left) We show the autocorrelation function $C(t)$ obtained as in Eq.
+equation B.1 when we analyze the random walk in the model indices during PTT sampling with the same RBMs analyzed in the main text (trained with pretraining+PCD) for the MNIST01 (top) and HGD (bottom) datasets.
+Both curves were obtained with a PTT run of $5\cdot 10^{4}$ MCMC steps, where the models used were automatically selected during training using the PTT moves acceptance criterion of 0.25.
+(Right) In blue we show the $\tau_{\mathrm{exp}}$ as a function of the target acceptance rate for MNIST01 (top) and HGD (bottom).
+The error bar results from the error of the mean of the values extracted from the different $C_{m}(t)$ and the shading marks the range between the largest and the smallest value of the $N_{\mathrm{m}}$ curves.
+In orange, we show the total simulation costs ($\tau_{\mathrm{exp}}$ multiplied by the number of models to be simulated in parallel) as a function of the target acceptance rate.
+This number of models saved, as a function of the acceptance rate, is shown in green in the inset.
+dataset $\tau_{\mathrm{exp}}$ $\tau_{\mathrm{int}}$
+MNIST01 258(29) 9.0(15)
+HGD 48(10) 3.8(7)
+Mickey 2(0) 1(0)
+Ising 18(3) 2.4(5)
+Table 1: Autocorrelation times associated to the random walk dynamics between models in the PTT sampling obtained with the models trained with pre-training+PCD on different datasets.
+Models used in the PTT are all chosen with the $a=0.25$ criterion.
+With these measures, we can try to optimize the number of models that we use for the PTT and that are automatically selected during training.
+The criterion here is that the acceptance of PTT exchanges between neighboring models must not go below a target acceptance $a$.
+In the figures shown so far, $a=0.25$.
+Let us justify this choice.
+The total physical time we need to simulate $N_{\mathrm{m}}$ models grows proportionally with $N_{\mathrm{m}}$ (unless one manages to properly parallelize the sampling, which is feasible but we did not do it), while the time needed to flow between very well trained and poorly trained models controlling the mixing time scales as $\sqrt{N_{\mathrm{m}}}/a$, where $a$ is the average acceptance of the exchanges between the models.
+Since $N_{\mathrm{m}}$ grows with $a$ in practice (as shown in the inset of Fig.
+6–right), it is expected that the total time can be optimized for a given value of $a$.
+In Fig.
+6–right we show $\tau_{\mathrm{exp}}$ in blue (and $\tau_{\mathrm{exp}}N_{\mathrm{m}}$ in orange) as a function of $a$ for the data sets MNIST01 (top) and HGD (bottom).
+These figures show that the choice of $a$ is not particularly crucial as long as it is between 0.2 and 0.5.
 Appendix C Training details
@@
 Name #Samples #Dimensions Train size Test size
+CelebA 30 000 1024 60% 40%
 Human Genome Dataset (HGD) 4500 805 60% 40%
+Ising 20 000 64 60% 40%
+Mickey 16 000 1000 60% 40%
 MNIST-01 10 610 784 60% 40%
-Mickey 16 000 1000 60% 40%
+MNIST 50 000 784 60% 40%
 Table 3: Hyperparameters used for the training of RBMs.
 Name Batch size #Chains #Epochs Learning rate #MCMC steps #Hidden nodes
+CelebA
+PCD 2000 2000 10 000 0.01 100 500
+Pre-train+PCD 2000 2000 10 000 0.01 100 500
 HGD
-PCD 2000 2000 10 000 0.01 100 185
-Jar-RBM 2000 10 000 10 000 0.01 100 185
-Pre-train+PCD 2000 2000 10 000 0.01 100 185
+PCD 2000 2000 10 000 0.01 100 200
+Pre-train+PCD 2000 2000 10 000 0.01 100 200
+Ising
+PCD 2000 2000 10 000 0.01 100 200
+Pre-train+PCD 2000 2000 10 000 0.01 100 200
+Mickey
+PCD 2000 2000 10 000 0.01 100 200
+Pre-train+PCD 2000 2000 10 000 0.01 100 200
 MNIST-01
 PCD 2000 2000 10 000 0.01 100 200
-Jar-RBM 2000 10 000 10 000 0.01 100 200
 Pre-train+PCD 2000 2000 10 000 0.01 100 200
+MNIST
+PCD 2000 2000 10 000 0.01 100 200
+Pre-train+PCD 2000 2000 10 000 0.01 100 200
+Table 4: Training time of the RBMs
+Name Training time (s)
+CelebA
+PCD $119.2\pm 0.8$
+Pre-train+PCD $3134.0\pm 0.6$
+HGD
+PCD $106.2\pm 0.3$
+Pre-train+PCD $266\pm 3$
+Ising
+PCD $62\pm 1$
+Pre-train+PCD $104\pm 2$
 Mickey
-PCD 2000 2000 10 000 0.01 100 100
-Jar-RBM 2000 10 000 10 000 0.01 100 100
-Pre-train+PCD 2000 2000 10 000 0.01 100 100
+PCD $118.3\pm 3$
+Pre-train+PCD $246.0\pm 0.7$
+MNIST-01
+PCD $100.6\pm 0.3$
+Pre-train+PCD $258.0\pm 0.9$
+MNIST
+PCD $105.7\pm 0.2$
+Pre-train+PCD $233.9\pm 0.2$
+Table 5: RCM Hyperparameters
+Dataset name # Dimensions # Hidden nodes Training time (s)
+CelebA 4 100 3006
+HGD 3 100 150
+Ising 3 100 32
+Mickey 2 100 118
+MNIST-01 3 100 148
+MNIST 3 100 119
 All experiments were run on a RTX 4090 with an AMD Ryzen 9 5950X.
-Appendix D Training of the RBM using the Jarzynski equation
-In this section, we describe a procedure similar to the one introduced in [37] for training the RBM by leveraging the Jarzynki equation.
+Appendix D Computation of the likelihood using AIS
+To compute the likelihood of the RBMs we used the Annealed Importance Sampling method introduced by Neal (2001) and recently studied in the case of the RBM Krause et al. (2020).
+D.1 Offline AIS
+We follow the method described in Krause et al.
+(2020).
+The AIS estimate is based on creating a sequence of machines, at one end a machine from which we can compute the partition function and at the other end the machine from which we want to compute the partition function.
+Let’s consider that we have $i=0,\dots,N_{\rm rbm}$ machines described by
+$p_{i}(\bm{x})=\frac{e^{-\mathcal{H}_{i}(\bm{x})}}{Z}=\frac{p_{i}^{*}(\bm{x})}{Z},$
+where $p_{0}=p_{\rm ref}$ denote the RBM from which we can compute the partition function (typically a RBM where the weight matrix is zero) and $p_{N_{\rm rbm}}=p_{\rm target}$ the machine from which we want to estimate the partition function.
+For each machine, we denote its set of parameters $\bm{\Theta}_{i}=\{\bm{w}_{i},\bm{\theta}_{i},\bm{\eta}_{i}\}$.
+Following Krause et al.
+(2020), defining a transition rule $T_{i}(\bm{x}_{i-1}\to\bm{x}_{i})$ in order to pass from a configuration $\bm{x}_{i-1}$ of the system $i-1$ to $\bm{x}_{i}$ of the system $i$, that we have the following identity
+$\frac{Z_{N_{\rm rbm}}}{Z_{0}}=\left\langle\exp\left(-\sum_{i=1}^{N_{\rm rbm}}\mathcal{H}_{i+1}(\bm{x}_{i})-\mathcal{H}_{i}(\bm{x}_{i})\right)\right\rangle_{\rm ForwardPath}.$ (29)
+There, if one can compute $Z_{0}$, the estimate of $Z_{N_{\rm rbm}}$ is obtained by averaging the r.h.s.
+of Eq.
+29 over the forward trajectory starting from a configuration $\bm{x}_{0}$ using the operators $T_{i}$.
+While the method is exact, the variance of the estimation relies on on close are two successive distribution $p_{i}$, $p_{i+1}$.
+Temperature AIS:
+a commonly used implementation of AIS, is to multiply the Hamiltonian of the system by a temperature, and to the different systems indexed by $i$, by a set of $\beta_{i}$ where $\beta_{0}=0$, $\beta_{i}<\beta_{i+1}$ and $\beta_{N_{\rm rbm}}=1$.
+Therefore, it is equivalent to choosing $\Theta_{i}=\beta_{i}\{\bm{w},\bm{\theta},\bm{\eta}\}$.
+Another version of Temperature AIS consist in choosing a slightly different reference distribution $p_{0}$ that take into account the bias of the model.
+In this setting, the intermediate distribution are taken as
+$p_{i}(\bm{x})=\frac{p_{\rm ref}^{(1-\beta_{i})}e^{-\beta_{i}\mathcal{H}(\bm{x})}}{Z_{i}}.$
+The main difference is that a prior distribution is used as reference $p_{0}=p_{\rm ref}$ and the temperatures interpolate between this prior and the target distribution. It is quite common to take
+$p_{\rm ref}(\bm{x})=\frac{\exp\left(\sum_{i}\theta_{i}v_{i}+\sum_{a}\eta_{a}h_{a}\right)}{\prod_{i}2\cosh(\theta_{i})\prod_{a}2\cosh(\eta_{a})}.$
+Trajectory AIS:
+the general derivation of AIS allow in principle to choose the set forward probability and/or intermediate system as one sees fit.
+Since again, a change of temperature in the system can induce discontinuous transition for which, successive distribution will be far away.
+While, when considering system that are neighbors in the learning trajectory, the system should pass through continuous transition which do not suffer this problem.
+Hence, we consider a set of RBMs along the learning trajectory.
+The RBMs are chosen following the prescription that the exchange rate between configuration (see Eq.
+5) is about $\sim 0.3$).
+Then the formula Eq.
+29 is used to estimate the partition function.
+D.2 Online AIS
+The online AIS (Annealed Importance Sampling) is the trajectory AIS evaluated at every gradient update during training.
+A set of permanent chains is initialized at the beginning of the training process.
+Every time the parameters are updated, a step is performed on these chains before using them to compute the current importance weights.
+This approach is computationally efficient relative to the cost of sampling (1 step vs.
+100 steps) and allows for a very precise estimation.
+Appendix E Training of the RBM using the Jarzynski equation
+In this section, we describe the procedure introduced in Carbone et al.
+(2024) for training energy-based models by leveraging the Jarzynki equation, and we adapt it to the specific case of the RBM.
+At variance with Carbone et al.
+(2024), we introduce an additional assumption on the transition probability of the parameters that will lead us to the the reweighting implemented in population annealing methods Weigel et al.
+(2021).
 In one of its formulations, the Jarzynski equation states that we can relate the ensemble average of an observable $\mathcal{O}$ with the average obtained through many repetitions of an out-of-equilibrium dynamical process.
-If we consider the training trajectory of an RBM, $p_{0}\rightarrow p_{1}\rightarrow\dots\rightarrow p_{t-1}\rightarrow p_{t}$, we can write
-$\langle\mathcal{O}\rangle_{t}=\frac{\left\langle\mathcal{O}e^{-W_{t}}\right\rangle_{\mathrm{traj}}}{\left\langle e^{-W_{t}}\right\rangle_{\mathrm{traj}}},$ (27)
-where the average on the lhs is done over the last model $p_{t}$, the averages on the rhs are taken across many different trajectory realizations and $W_{t}$ is a trajectory-dependent importance factor.
-By all practical means, under the assumption of having quasi-adiabatic parameters updates, namely $p(\Theta_{t-1}\rightarrow\Theta_{t})=p(\Theta_{t}\rightarrow\Theta_{t-1})$, this means that we can assign to each Markov chain of the simulation $\bm{x}^{(r)}$, $r=1,\dots,R,$ an importance weight given by:
-$W_{t}^{(r)}=\sum_{\tau=1}^{t}[\mathcal{H}_{\tau}(\bm{x}^{(r)}_{\tau-1})-\mathcal{H}_{\tau-1}(\bm{x}^{(r)}_{\tau-1})]$ (28)
-and then compute the gradient of the log-likelihood by means of a weighted average over the chains:
-$\langle\mathcal{O}\rangle_{t}\simeq\frac{\sum_{r=1}^{R}\mathcal{O}(\bm{x}^{(r)})~{}e^{-W^{(r)}_{t}}}{\sum_{r=1}^{R}e^{-W^{(r)}_{t}}}.$ (29)
+Let us consider a path $\boldsymbol{y}=\{(\boldsymbol{x}^{0},\boldsymbol{\theta}^{0}),(\boldsymbol{x}^{1},\boldsymbol{\theta}^{1}),\dots,(\boldsymbol{x}^{t},\boldsymbol{\theta}^{t})\}$, where $\bm{x}^{i}=(\bm{v}^{i},\bm{h}^{t})$ refers to the model variables and $\bm{\theta}^{i}=(w^{i},\eta^{i},\theta^{i})$, the model parameters, both at time $i$.
+If we consider the training trajectory of an RBM, $p_{0}\rightarrow p_{1}\rightarrow\dots\rightarrow p_{t-1}\rightarrow p_{t}$, we can write the average of an observable $\mathcal{O}$ over the last model of the trajectory $p_{t}$ as
+$\langle\mathcal{O}\rangle_{t}=\frac{\left\langle\mathcal{O}~{}e^{-W^{t}}\right\rangle_{\mathrm{traj}}}{\left\langle e^{-W^{t}}\right\rangle_{\mathrm{traj}}}=\frac{\sum_{r=1}^{R}\mathcal{O}(\boldsymbol{y}_{r})~{}e^{-W_{r}^{t}}}{\sum_{r=1}^{R}e^{-W_{r}^{t}}},$ (30)
+where $W^{t}$ is a trajectory-dependent importance factor and the averages on the rhs are taken across many different trajectory realizations. In particular, $W_{t}$ is given by the recursive formula:
+$W^{t}(\bm{y})$ $=$ $\mathcal{H}^{t}(\bm{x}^{t})-\mathcal{H}^{0}(\bm{x}^{0})-\sum_{i=1}^{t}\log\tilde{T}^{i}(\bm{y}^{i}\rightarrow\bm{y}^{i-1})+\sum_{i=1}^{t}\log T^{i}(\bm{y}^{i-1}\rightarrow\bm{y}^{i})$ (31)
+$=$ $W^{t-1}(\bm{y})+\mathcal{H}^{t}(\bm{x}^{t})-\mathcal{H}^{t-1}(\bm{x}^{t-1})-\log\frac{\tilde{T}^{t}(\bm{y}^{t}\rightarrow\bm{y}^{t-1})}{T^{t}(\bm{y}^{t-1}\rightarrow\bm{y}^{t})}$ (32)
+Where $T$ and $\tilde{T}$ are, respectively, the forward and backward transition probabilities.
+The transition matrix between two states $\bm{y}^{t-1}\rightarrow\bm{y}^{t}$ is factorized into two different subsequent movements:
+  1. 1.
+    Update the variables $\bm{x}^{t-1}\to\bm{x}^{t}$ on the fixed model $\bm{\theta}^{t-1}$,
+  2. 2.
+    Update the parameters $\bm{\theta}^{t-1}\to\bm{\theta}^{t}$ on fixed variables $\bm{x}^{t}$.
+In other words,
+$T(\bm{y}^{t-1}\to\bm{y}^{t})=w_{\bm{\theta}^{t-1}}(\bm{x}^{t-1}\to\bm{x}^{t})t_{\bm{x}^{t}}(\bm{\theta}^{t-1}\to\bm{\theta}^{t}),$ (33)
+and
+$\tilde{T}(\bm{y}^{t}\to\bm{y}^{t-1})=t_{\bm{x}^{t}}(\bm{\theta}^{t}\to\bm{\theta}^{t-1})w_{\bm{\theta}^{t-1}}(\bm{x}^{t}\to\bm{x}^{t-1}),$ (34)
+so that
+$\log\frac{\tilde{T}^{t}(\bm{y}^{t}\rightarrow\bm{y}^{t-1})}{T^{t}(\bm{y}^{t-1}\rightarrow\bm{y}^{t})}=\log\frac{w_{\bm{\theta}^{t-1}}(\bm{x}^{t}\to\bm{x}^{t-1})}{w_{\bm{\theta}^{t-1}}(\bm{x}^{t-1}\to\bm{x}^{t})}+\log\frac{t_{\bm{x}^{t}}(\bm{\theta}^{t}\to\bm{\theta}^{t-1})}{t_{\bm{x}^{t}}(\bm{\theta}^{t-1}\to\bm{\theta}^{t})}.$ (35)
+We know the ratio between the transition probabilities between $\bm{x}^{t-1}\to\bm{x}^{t}$, because evolution of the variables during the simulation is performed using MCMC simulations that satisfy detailed balance.
+This means that:
+$\log\frac{w_{\bm{\theta}^{t-1}}(\bm{x}^{t}\to\bm{x}^{t-1})}{w_{\bm{\theta}^{t-1}}(\bm{x}^{t-1}\to\bm{x}^{t})}=\mathcal{H}^{t-1}(\bm{x}^{t})-\mathcal{H}^{t-1}(\bm{x}^{t-1}),$ (36)
+and we are left with
+$W^{t}(\bm{y})$ $=$ $W^{t-1}(\bm{x})+\mathcal{H}^{t}(\bm{x}^{t})-\mathcal{H}^{t-1}(\bm{x}^{t})-\log\frac{t_{\bm{x}^{t}}(\bm{\theta}^{t}\to\bm{\theta}^{t-1})}{t_{\bm{x}^{t}}(\bm{\theta}^{t-1}\to\bm{\theta}^{t})}.$ (37)
+The standard formula that we use in the paper assumes that the changes in $\bm{\theta}$ are so small that we can consider the transition probabilities to be symmetrical and the last term in the rhs cancels out.
+In that case, we recover the known expressions of the population annealing.
 Notice that, since Eq.
-(27) is an exact result, the importance weights should, in principle, eliminate the bias brought by the non-convergent chains used for approximating the log-likelihood gradient in the classical PCD scheme.
+equation 30 is an exact result, the importance weights should, in principle, eliminate the bias brought by the non-convergent chains used for approximating the log-likelihood gradient in the classical PCD scheme.
 However, after many updates of the importance weights, one finds that only a few chains carry almost all the importance mass.
 In other words, the vast majority of the chains we are simulating are statistically irrelevant, and we expect to get large fluctuations in the estimate of the gradient because of the small effective number of chains contributing to the statistical average.
-A good observable for monitoring this effect is the Effective Sample Size (ESS), defined as [37]
+A good observable for monitoring this effect is the Effective Sample Size (ESS), defined as Carbone et al.
+(2024)
 $\mathrm{ESS}=\frac{\left(R^{-1}\sum_{r=1}^{R}e^{-W^{(r)}}\right)^{2}}{R^{-1}\sum_{r=1}^{R}e^{-2W^{(r)}}}\in[0,1],$ (38)
 which measures the relative dispersion of the weights distribution.
-A way of circumventing the weight concentration on a few chains, then, is to resample the chain population according to the importance weights every time the ESS drops below a certain threshold, for instance 0.5.
+Then, a way of circumventing the weight concentration on a few chains is to resample the chain population according to the importance weights every time the ESS drops below a certain threshold, for instance 0.5.
 After this resampling, all the chain weights have to be set to 1 ($W^{(r)}=0\,\,\forall r=1,\dots,R$).
+Appendix F Tuning of the hyperparameters
+In order to justify our choice of hyperparameters in the main text, we illustrate on Fig.
+7 the value of the log-likelihood as a function of the number of hidden nodes and of the learning of the machine.
+In order to be fair, the comparison is done at a constant number of gradient updates of $10000$.
+In this regime, the machines that perform the best have a learning rate of $\gamma\sim 0.1,0.005$ and about $N_{h}=500,1000$ hidden nodes.
+We can expect that smaller learning rates reach the same value of the log-likelihood but for a much longer training.
+We clearly see the benefit of the RCM, where the log-likelihood achieves much higher values in general.
+Figure 7: Comparison of the end log-likelihood reached by the RBM using different pairs of values for the learning rate and the hidden layer dimension on MNIST-01.
+The left panel shows these result when performing PCD training and the right one the results when pre-training the RBM before PCD.
+Appendix G Continuous phase transition
+In statistical physics, and more precisely in mean-field model, a model described by the Gibbs-Boltzmann distribution is studied in the infinite size limit, $N\to\infty$, and characterize the state of the system by investigating the free energy of the system, defined by minus the log partition function
+$F[\bm{\theta}]=-\log(Z[\bm{\theta}])$
+One of the goals is thus to determine the typical configurations that should be expected given a set of model parameters.
+In the case of the RBM, the question is what are the typical visible (and/or hidden) variables that one expects to observe.
+For example, one may ask what typical energies one expects to observe for the parameters $\bm{\theta}$.
+For this purpose, let us rewrite the partition function as a sum over all possible values of the energy
+$Z=\sum_{E}e^{-E}g(E)=\sum_{E}e^{-E+S(E)}=\sum_{E}e^{-F(E)}=\sum_{E}e^{-Nf(E)}$
+where $g(E)$ the number of states with a given energy $E$, and thus $S(E)=\log g(E)$ is the entropy.
+Finally, the free energy is defined as $F(E)=E-S(E)$ and $f(E)=F(E)/N$ is the free energy per variable.
+All $Z$, $E$ and $F$ depend on the model parameters.
+$N$ refers to the number of variables.
+For $N\to\infty$, the sum in Z is then dominated by the states with minimum free energy.
+A “first order”, or discontinuous transition, occurs in the parameters if the first derivative of $f(E)$ is discontinuous with respect to the parameters (in physics, the parameter is typically the temperature).
+A “second order”, or continuous transition, occurs when the first derivative is continuous, but the second derivative is discontinuous.
+In the case of the RBM, the first learning transition can be mapped to a very simple model, the so-called Curie-Weiss model, as recently described in Ref.
+Bachtis et al.
+(2024).
+The Curie-Weiss model is fully solvable, and for this reason the meaning of a continuous transition is perhaps easier to understand there.
+Let’s give it a try.
+The Curie-Weiss model consists in a set of $N$ discrete variables $s_{i}=\pm 1$.
+We define the (exponential) distribution over these variables as
+$p(s)=\frac{1}{Z}\exp(\frac{\beta}{N}\sum_{i<j}s_{i}s_{j}+\beta H\sum_{i}s_{i})$
+where $\beta$ and $H$ are parameters of the model and $Z$ the normalization constant also called partition function.
+The question is to understand the “structure” of the distribution as a function of $\beta$.
+Case $H=0$:
+we expect that for $\beta$ small each spin behaves as an isolated Rademacher random variables (each spin being ±$\pm 1$ with probability one-half), while for large value of $\beta$, the distribution of the system is dominated by configurations where the variables have (almost) all the same signs.
+The key mathematical aspect is to study the system in the limit where $N\to\infty$.
+A way to study this distribution is to investigate the moment generating function.
+It is possible to show (rigorously in this precise model), that the probability of $m=N^{-1}\sum_{i}s_{i}$, also called the magnetization of the system, is given by $p(m)\propto\exp(-N\Omega(m))$ (in the large $N$ limit) where $\Omega(m)$ is a large deviation function.
+In the CW model, it is given by
+$\Omega(m)=\frac{\beta m^{2}}{2}-\log\left[2\cosh(\beta m)\right].$
+The structure of $\Omega(m)$ is such that for $\beta<\beta_{c}=1$, it is a convex function with only one minimum in $m=0$, while for $\beta>\beta_{c}$ it has two symmetric minima $m=\pm m(\beta)\neq 0$ with $\Omega(m(\beta))=\Omega(-m(\beta))$ (the two minimum are equally probable).
+The values (in function of $\beta$) of the $m$ are continuous: $m=0$ for $\beta\in[0,1]$, and then positive until it saturates to $1$ as $\beta\to\infty$.
+In practice, it means that the distribution $p(m)$ pass from an unimodal distribution to a bimodal one as $\beta$ increases.
+The point where it happens is at $\beta_{c}=1$ and is called a second order (continuous) phase transition because de value of $m$ changes continuously from zero to non-zero values.
+A particular interesting properties is that at $\beta_{c}$, the function $\Omega(m)$ develops a non-analyticity in its second-order derivative, that can be linked with long-range fluctuations.
+For a recent review Kochmański et al.
+(2013), and for more rigorous results Sinai (2014).
+Case $H\neq 0$:
+in this case we expect that the system will always be biased toward configurations where the spins have the same direction(sign) as the bias (or field) $H$.
+Yet, for small value of $\beta$, the bias will be quite small, hence the magnetization will be small but not zero, $m\sim\beta H$.
+When $\beta$ is large, the system will now have a stronger and strong magnetization toward the magnetic field $m\sim{\rm sign}(H)$.
+Now we can again compute the large deviation function of the system
+$\Omega(m)=\frac{\beta m^{2}}{2}-\log\left[2\cosh(\beta m)\right]+Hm.$
+Again, it is possible to analyze the function $\Omega$ as a function of $\beta$ and $H$.
+For $\beta<1$ the function is convex and there is only one minimum located at $m>0$.
+When $\beta>1$ two situations can be distinguished.
+First, let’s denote
+$H_{\rm sp}(\beta)=\sqrt{\frac{\beta-1}{\beta}}-{\rm atanh}\left(\frac{\beta-1}{\beta}\right).$
+If $H>H_{\rm sp}(\beta)$, then the large deviation function has only one minimum for positive $m$.
+Now if $H<H_{\rm sp}(\beta)$, then $\Omega$ will present two minima.
+A first one for positive $m$ and a second one for negative $m$.
+The reason is that the pairwise interactions “manage” to create a metastable state for negative magnetization even in the presence of a positive field or bias.
+It is clear that this state will be unstable w.r.t.
+the one with positive magnetization which traduces in a minimum of higher value in the large deviation function.
+By changing the value of $H$ from large positive values toward large negative ones, we then pass from a system with only one state, then two states with one being subdominant, and finally back to one state but where the magnetization has changed its sign.
+In the infinite size limit, the system will always dominated by the state where $\Omega$ is at its lowest value.
+However, if one prepare a system close to a metastable state, the system will remain trap there until it disappears.
+Figure 8: Left: The case $H=0$ for various temperature.
+We illustrate the symmetric shape of the large deviation function and how the two minima emerges from the centers.
+The minima of $m$ go continuously from zero to non-zero values.
+In the inset we also illustrate the shape of the probability distribution for moderately large size $N=500$.
+For $N\to\infty$ they converge toward a $\delta$ distribution.
+Right: The case $H\neq 0$.
+In this example, we illustrate what happen where the bias parameter $H$ is changed at a fix temperature $\beta=1.15>1$.
+We see that for small values of $H$, $\Omega$ has two minima, yet the dominant one suppresses exponentially the other one for large system size.
+For larger values of the bias the metastable minimum disappears.
+We show how the place of the minimum jump when the bias goes from negative to positive values.
+How is this related to the training of the RBM ?
+First, it has been shown in previous works Decelle et al.
+(2018); Bachtis et al.
+(2024) that the RBM’s training undergoes a set of continuous transitions, at least at the beginning of the learning.
+Hence, changing the parameters of the RBM following the learning trajectories, the transitions are continuous.
+Then in Decelle & Furtlehner (2021a) it is shown that the relaxed version of the RBM, namely the RCM undergoes a first-order transition when changing the temperature.
+Schematically to understand that, it is enough to see how the parameters $\beta$ acts on the system.
+In the case of the RCM, the energy term behaves in the standard way, i.e.
+linearly with the inverse temperature $\beta$ and the free energy takes the ordinary form
+$\log(Z)=-\beta F=-\beta E+S.$
+Therefore $\beta$ act as a trade-off parameters between the energy $E$ and the entropy $S$.
+By varying $\beta$, we change the contribution of the energy while keeping the entropy fixed.
+Hence, when the corresponding large deviation function of the RCM has reached a point having many minima with the same value of the free energy up to ${}\mathcal{O}(1/N)$ corrections, all these minima realize in principle different trade-off between energy and entropy.
+As a result a slight change in $\beta$ will offset this sensible balance and the respective free energies will acquire differences of order ${\mathcal{O}}(N)$ leading to favor dramatically one state against the others, i.e.
+the one having originally the lowest [resp.
+the largest] energy when $\beta$ is increased.
+It is very likely that by changing $\beta$ a discontinuous transition is encountered, where the minima of $\Omega$ vanishes far away from each other rather than a continuous one.
+Concerning the RBM the argument for the occurence of such a phase transition, is only slightly more involved than for Coulomb machine because for RBM the energy (obtained after summing over latent variables) has a non-linear behavior w.r.t.
+$\beta$.
+A formal argument would go schematically as follows: just writing the local minima of the free energy (function of magnetization in the reduced intrinsic space) and looking at how the equilibrium is displaced with temperature shows that their energy vary indeed in a non-linear way with temperature and non-uniformly regarding their position on the magnetization manifold, while the entropy contribution again do not change.
+This again necessarily implies that changing even slightly the temperature will break the highly sensitive equilibrium obtained between these state corresponding to the multimodal distribution and will again typically favor one state among all.
+This scenario is expected to occur as soon as the data are located on low dimensional space ($d=O(1)$ ) compared to a large embedding space ($d=O(N),\ N\gg 1$) which is quite common in our point of view, at least for the type of data we are interested in.
+To conclude, moving the parameters along the learning trajectory, leads the system to undergo continuous transition where $p_{\bm{\Theta(t)}}(\bm{s})$ is very close to $p_{\bm{\Theta}(t+\delta t)}(\bm{s})$.
+Changing the temperature tends to produces discontinuous transition where $p_{\beta(t)\bm{\Theta}}(\bm{s})$ can happen to be very far from $p_{\beta(t+\delta t)\bm{\Theta}}(\bm{s})$.
+Appendix H Overfitting and privacy loss as quality indicators
+In this section, we examine the quality of the samples generated, regarding overfitting and privacy criteria which have been defined for genomic data in particular.
+We look at this on the models trained with PCD with and without pre-training.
+We focus on the human genome dataset, as shown in Fig.
+1–C, to evaluate the ability of various state-of-the-art generative models to generate realistic fake genomes while minimizing privacy concerns (i.e., reducing overfitting).
+Recent studies (Yelmen et al., 2021; 2023) have thoroughly investigated this for a variety of generative models.
+Both studies concluded that the RBM was the most accurate method for generating high-quality and private synthetic genomes.
+The comparison between models relies primarily on the Nearest Neighbor Adversarial Accuracy ($AA_{\rm TS}$) and privacy loss indicators, introduced in Ref.
+Yale et al.
+(2020), which quantify the similarity and the level of ”privacy” of the data generated by a model w.r.t.
+the training set.
+We have $AA_{\rm TS}=\frac{1}{2}\bigl{(}AA_{\rm True}+AA_{\rm Synth}\bigr{)}$ where $AA_{\rm True}$ [resp.
+$AA_{\rm Synth}$] are two quantities in $[0,1]$ obtained by merging two sets of real and synthetic data of equal size $N_{s}$ and measuring respectively the frequency that a real [rep.
+synthetic] has a synthetic [resp.
+real] as nearest neighbor.
+If the generated samples are statistically indistinguishable from real samples, both frequencies $AA_{\rm True}$ and $AA_{\rm Synth}$ should converge to 0.5 at large ${N_{\mathrm{s}}}$.
+$AA_{\rm TS}$ can be evaluated both with train or test samples and the privacy loss indicator is defined as ${\rm Privacy\ loss}=AA_{\rm TS}^{\rm test}-AA_{\rm TS}^{\rm train}$ and is expected to be strictly positive.
+Fig.
+10 shows the comparison of $AA_{\rm TS}$ and privacy loss values obtained with our two models, demonstrating that the pre-trained RBM clearly outperforms the other model, and even achieves better results ($AA_{\rm TS}$ values much closer to 0.5) than those discussed in Yelmen et al.
+(2021; 2023).
+Figure 9: Evolution of the AATS (first row) and privacy loss (second row) indicator during the training of pretrained RBMs.
+The online estimation of the train and test LL is given at the third row.
+The overfitting is clearly detected in the HGD and MICKEY dataset and the divergence between the train and test LL coincides with a sharp rise in the privacy loss.
+Figure 10: Comparison of the evolution of overfitting metrics along the log-likelihood between a pretrained RBM (A) and a RBM trained from scratch using PCD (B).
+(C) shows the comparison between the AASyn and AATruth between the two machines.
+The machines were selected in order to have a comparable privacy loss (D), right before starting to overfit.
+Figure 11: Comparison of log-likelihood estimation w.r.t. the exact ll on the MNIST-01 dataset.
+Figure 12: Comparison of log-likelihood estimation w.r.t. the exact ll on the MNIST dataset.
+Figure 13: Comparison of log-likelihood estimation w.r.t. the exact ll on the Ising dataset.
+Figure 14: Comparison of log-likelihood estimation w.r.t. the exact ll on the MICKEY dataset.
+Appendix I Extra figures
+Figure 15: Samples generated with the low-rank RBM trained with 3 directions plus a bias.
+We show the samples generated by the pre-trained machined for the 5 datasets discussed in the paper and projected on the PCA of the dataset.
+For the image datasets, we also show the images generated.
+Figure 16: We show the evolution of the tr-AIS log-likelihood in the training of the entire MNIST dataset (containing digits from 0 to 9) with and without the pre-training.
+The pre-training does not suppose a particular advantage but neither is a disadvantage.

### v3

@@
 Abstract
-Restricted Boltzmann Machines (RBMs) are effective tools for modeling complex systems and deriving insights from data.
-However, training these models with highly structured data presents significant challenges due to the slow mixing characteristics of Markov Chain Monte Carlo (MCMC) processes.
-In this study, we build upon recent theoretical advancements in RBM training, focusing on the gradual encoding of data patterns into singular vectors of the coupling matrix, to significantly reduce the computational cost of training (in very clustered datasets) and evaluating and sampling in RBMs in general.
-The learning process is analogous to thermodynamic continuous phase transitions observed in ferromagnetic models, where new modes in the probability measure emerge in a continuous manner.
-Such continuous transitions are associated with the critical slowdown effect, which adversely affects the accuracy of gradient estimates, particularly during the initial stages of training with clustered data.
-To mitigate this issue, we propose a pre-training phase that encodes the principal components into a low-rank RBM through a convex optimization process.
-This approach facilitates efficient static Monte Carlo sampling and accurate computation of the partition function.
-Furthermore, we exploit the continuous and smooth nature of the parameter annealing trajectory to achieve reliable and computationally efficient log-likelihood estimations, enabling online assessment during the training process, and proposing a novel sampling strategy termed parallel trajectory tempering that outperforms previously optimized MCMC methods.
-Our results demonstrate that this innovative training strategy enables RBMs to effectively address highly structured datasets that conventional methods struggle with.
-Additionally, we provide evidence that our log-likelihood estimation is more accurate than traditional, more computationally intensive approaches in controlled scenarios.
-Moreover, the parallel trajectory tempering algorithm significantly accelerates MCMC processes compared to existing and conventional methods.
+Restricted Boltzmann Machines (RBMs) are powerful tools for modeling complex systems and extracting insights from data, but their training is hindered by the slow mixing of Markov Chain Monte Carlo (MCMC) processes, especially with highly structured datasets.
+In this study, we build on recent theoretical advances in RBM training and focus on the stepwise encoding of data patterns into singular vectors of the coupling matrix, significantly reducing the cost of generating new samples and evaluating the quality of the model, as well as the training cost in highly clustered datasets.
+The learning process is analogous to the thermodynamic continuous phase transitions observed in ferromagnetic models, where new modes in the probability measure emerge in a continuous manner.
+We leverage the continuous transitions in the training process to define a smooth annealing trajectory that enables reliable and computationally efficient log-likelihood estimates.
+This approach enables online assessment during training and introduces a novel sampling strategy called Parallel Trajectory Tempering (PTT) that outperforms previously optimized MCMC methods.
+To mitigate the critical slowdown effect in the early stages of training, we propose a pre-training phase.
+In this phase, the principal components are encoded into a low-rank RBM through a convex optimization process, facilitating efficient static Monte Carlo sampling and accurate computation of the partition function.
+Our results demonstrate that this pre-training strategy allows RBMs to efficiently handle highly structured datasets where conventional methods fail.
+Additionally, our log-likelihood estimation outperforms computationally intensive approaches in controlled scenarios, while the PTT algorithm significantly accelerates MCMC processes compared to conventional methods.
 1 Introduction
-Energy-based models (EBMs) are a long-established approach to generative modeling, with the Boltzmann Machine (BM)  (Hinton & Sejnowski, 1983) and Restricted Boltzmann Machine (RBM) (Smolensky, 1986; Ackley et al., 1985) being among the earliest examples.
+Energy-based models (EBMs) are a long-established approach to generative modeling, with the Boltzmann Machine (BM)  (hinton1983optimal) and Restricted Boltzmann Machine (RBM) (Smolensky; ackley1985learning) being among the earliest examples.
 They offer a clear framework for capturing complex data interactions and, with simple energy functions, can reveal underlying patterns, making them especially valuable for scientific applications.
-Pairwise-interacting models (Nguyen et al., 2017) have been widely used for decades for inference applications in fields like neuroscience (Hertz et al., 2011) and computational biology (Cocco et al., 2018).
-More recent research has shown that RBMs can also be reinterpreted as many-body physical models (Decelle et al., 2024; Bulso & Roudi, 2021), analyzed to identify hierarchical relationships in the data (Decelle et al., 2023), or used to uncover biologically relevant constituent features (Tubiana et al., 2019; Fernandez-de Cossio-Diaz, 2025).
+Pairwise-interacting models (nguyen2017inverse) have been widely used for decades for inference applications in fields like neuroscience (hertz2011ising) and computational biology (cocco2018inverse).
+More recent research has shown that RBMs can also be reinterpreted as many-body physical models (decelle2024inferring; bulso2021restricted), analyzed to identify hierarchical relationships in the data (decelle2023unsupervised), or used to uncover biologically relevant constituent features (tubiana2019learning; Fernandez-de-Cossio-Diaz2025).
 In essence, RBMs present a great compromise between expressibility and interpretability: they are universal approximators capable of modeling complex datasets and simple enough to allow a direct extraction of insights from data.
-This interpretative task is particularly challenging when dealing with more powerful models such as generative convolutional networks EBMs (Xie et al., 2016; Song & Kingma, 2021) and other state-of-the-art models in general.
-At this point, it is also important to stress that RBM are not only interpretable, but they perform particularly well when dealing with tabular datasets with discrete variables, and where data is not abundant, which is often the case in genomics/proteomics or neural recording datasets.
-For instance, RBMs have shown to be among the most reliable approaches to generate synthetic human genomes (Yelmen et al., 2021; 2023) or to model inmunological response (Bravi et al., 2021a; b; 2023) or neural activity (van der Plas et al., 2023; Quiroz Monnens et al., 2024).
-Despite the appealing modeling properties of RBMs, they are notoriously difficult to train, a challenge common to EBMs in general.
-The main difficulty arises from the computation of the log-likelihood gradient, which requires an ergodic exploration of a dynamically evolving and potentially complex free energy landscape using Markov Chain Monte Carlo (MCMC) processes.
-Recent studies have shown that models trained with non-convergent MCMC processes suffer from out-of-equilibrium dynamic memory effects (Nijkamp et al., 2019; 2020; Decelle et al., 2021).
-This dynamical behavior can be explained analytically using moment-matching arguments  (Nijkamp et al., 2019; Agoritsas et al., 2023).
-While exploiting these effects can yield fast and accurate generative models, even for highly structured data (Carbone et al., 2023) or high-quality images with RBMs (Liao et al., 2022), this approach results in a sharp separation between the model’s Gibbs-Boltzmann distribution and the dataset distribution, thereby undermining the interpretability of the model parameters as investigated in Agoritsas et al.
-(2023); Decelle et al.
-(2024).
-Thus, to extract meaningful information from datasets using RBMs, it is essential to ensure proper mixing of the chains during training, in short, one needs equilibrium models.
-Both the ability to train an RBM in equilibrium and to generate convincing new samples from its equilibrium measure strongly depend on the dataset in question.
-For typical image datasets such as MNIST or CIFAR-10, good RBMs can be obtained by increasing the number of MCMC steps.
-However, this approach is no longer feasible for highly structured datasets (Béreux et al., 2023).
-Natural datasets from which one seeks scientific insights are often highly structured, such as genomics/proteomics data, neural-recordings or low-temperature configurations in physical systems.
-These datasets typically exhibit distinct clusters, identifiable via principal component analysis (PCA) which form distant groups of similar entries often associated to different labels.
-We show an example of the PCA of the 5 datasets we will be studying in this work in Fig.
+This interpretative task is particularly challenging when dealing with more powerful models, such as generative convolutional network EBMs (pmlr-v48-xiec16; song2021train) and other state-of-the-art models in general.
+At this point, it is also important to stress that RBM are not only interpretable, but they perform particularly well when dealing with tabular datasets with discrete variables and where data are not abundant, which is often the case in genomics/proteomics or neural recording datasets.
+For example, RBMs have been shown to be among the most reliable approaches to generate synthetic human genomes (yelmen2021creating; yelmen2023deep) or to model the immune response (bravi2021probing; bravi2021rbm; bravi2023transfer) or neural activity (van2023neural; Quiroz_Monnens_2024).
+RBMs, like other EBMs, are challenging to train due to the difficulty of computing the log-likelihood gradient, which requires ergodic exploration of a complex free-energy landscape via Markov Chain Monte Carlo (MCMC).
+Training with non-convergent MCMC introduces out-of-equilibrium memory effects, as shown in recent studies  (nijkamp2019learning; nijkamp2020anatomy; decelle2021equilibrium), which can be analytically explained using moment-matching arguments  (nijkamp2019learning; agoritsas2023explaining).
+While these effects enable fast and accurate generative models for structured data (carbone2024fast) and high-quality images with RBMs (liao2022gaussian), they create a sharp disconnect between the model’s Gibbs-Boltzmann and dataset distributions, undermining parameter interpretability (agoritsas2023explaining; decelle2024inferring).
+For meaningful insights, RBMs require proper chain mixing during training, emphasizing the need for equilibrium models.
+Working with equilibrium RBMs poses three primary challenges: training, evaluating model quality, and sampling from the trained model.
+Both training and sampling rely on MCMC methods to draw equilibrium samples, but their efficiency depends heavily on the dataset’s structure.
+While RBMs can be effectively trained on image datasets like MNIST or CIFAR-10 with sufficient MCMC steps, this approach falters for highly structured datasets (bereux2023learning), such as genomics, neural recordings, or low-temperature physical systems.
+These datasets often exhibit multimodal distributions with distinct clusters, as revealed by PCA (Fig.
+1), which hinder mixing.
+Clustering is prominent in the first four datasets studied here, in contrast to the more compact PCA pattern seen in CelebA or full MNIST.
+Dataset details are provided in Fig.
 1.
-The first 4 datasets are very clustered, while CelebA is not, as it presents a rather compact PCA.
-A very similar (non-clustered) situation is obtained if the entire MNIST dataset is considered.
-Details about these datasets are given in the caption of figure 1.
-During training, the model must evolve from an initial normal distribution to an increasingly multimodal distribution.
-Sampling from multimodal distributions is particularly challenging because the mixing times are controlled by the transition times between modes.
-But this is not the only difficulty.
-In RBMs, these distant modes can be shown analytically(Decelle et al., 2017; 2018; Bachtis et al., 2024) to gradually appear in the model’s equilibrium measure during training by a process formally analogous to the ferromagnetic phase transition in simple spin models, the process by which the magnetic spins of a material spontaneously align with a fixed pattern when the temperature goes below a certain transition temperature.
-The RBM training is characterized as a sequence of such phase transitions in which the specialization of the model parameters plays the role of temperature reduction.
-These pattern-encoding phase transitions are continuous phase transitions (also known as second-order transitions or critical transitions), and these types of transitions are associated with a sudden arrest in the characteristic times of the visible variables, a phenomenon known as critical slowdown, which causes the MCMC mixing time to diverge with a power of the system size when the transition point is approached.
-The critical slowdown effects that occur in the training trajectory were recently studied in Bachtis et al.
-(2024).
-A brief introduction to phase transitions and the associated dynamical effects can be found in the SI G.
-This sampling challenge not only hinders the training process, but also limits the model’s ability to generate new samples.
-Obtaining new and independent configurations would require an impractically large number of sampling steps.
+Multimodal distributions exacerbate sampling challenges since mixing times depend on the MCMC chains’ ability to transition between clusters.
+In RBMs, these distant modes emerge during training through continuous phase transitions (e.g., second-order or critical transitions) that encode patterns (Decelle_2017; decelle2018thermodynamics; bachtis2024cascade).
+These transitions lead to critical slowdown (hohenberg1977theory), where mixing times diverge with system size as the transition point nears.
+Critical slowdown along the training trajectory was recently analyzed in (bachtis2024cascade).
+For a brief introduction to phase transitions and associated dynamical effects, see Supplemental Information (SI) G.
+All this hampers both the training process and the model’s ability to generate independent samples, as achieving adequate mixing requires prohibitively long simulations.
+Model quality evaluation adds to the challenge, particularly in unsupervised learning.
+A common qualitative approach compares equilibrium samples to data via PCA projections.
+Quantitatively, moment comparisons between data and generated distributions offer insights but fail to detect mode collapse.
+Log-likelihood computation provides a single-score metric and tracks overfitting during training, but exact computation is infeasible due to the intractable partition function.
+Instead, methods such as Annealed Importance Sampling (AIS) (neal2001annealed) are used for estimation but become unreliable for highly multimodal distributions and are computationally impractical for online evaluations (see Section 4.1).
+In this paper, we build on recent advances in understanding the evolution of the landscape during the training of RBM models to propose 3 innovative approaches for tackling the challenges of evaluating, generating samples and training RBMs on highly structured datasets:
+  1. 1.
+    We propose a novel method for estimating log-likelihood (LL) by leveraging the softness of the training trajectory.
+This approach enables efficient online LL computation during training at minimal cost or offline reconstruction using trajectory annealing or parallel tempering.
+Moreover, we demonstrate that this method is not only less computationally expensive than existing techniques but also more accurate in controlled experiments.
+The method is introduced in Sect.
+4.1.
+  2. 2.
+    We propose a variation of the standard parallel tempering algorithm, where parameters from different stages of the learning process are exchanged instead of temperatures.
+This new sampling approach significantly outperforms previous optimized methods and is detailed in Section 4.2.
+  3. 3.
+    We introduce a pretraining strategy based on the low-rank RBM framework of (decelle2021exact) to mitigate gradient degradation in early training, caused by mixing time divergence when encoding distant modes in structured datasets.
+Our approach extends this framework to real datasets by increasing constrained directions and incorporating a trainable bias, essential for image data.
+We also refine static Monte Carlo sampling for efficient equilibrium sampling.
+See Section 5 and SI A for details.
+We also provide an in-depth discussion in Sect. 2 on the physical reasons behind the failure of previous algorithms when handling clustered datasets.
 Figure 1: Datasets.
@@
 2 Related work
-Training EBMs by maximizing log-likelihood has long been a challenge in the community (LeCun et al., 2006; Song & Kingma, 2021).
-EBMs gained popularity with the introduction of the contrastive divergence algorithm (Hinton, 2002), in which a set of parallel chains is initialized on independent examples in the minibatch and the MCMC process iterates for a few steps.
-Despite its widespread use, this algorithm yields models with poor equilibrium properties that are ineffective as generative models (Salakhutdinov & Murray, 2008; Desjardins et al., 2010; Decelle et al., 2021).
-An improvement is the persistent contrastive divergence (PCD) algorithm (Tieleman, 2008), which maintains a permanent chain in which the last configurations used to estimate the previous gradient update are reused.
-PCD acts like a slow annealing process improving gradient estimation quality.
-However, it often fails on clustered data as the statistical properties of the permanent chain quickly move away from the equilibrium measure and degrade the model (Béreux et al., 2023).
-This problem, which is primarily related to phase coexistence, can be addressed with constrained MCMC methods if appropriate order parameters are identified.
-For RBMs, these order parameters are related to the singular value decomposition of the model coupling matrix, which enables efficient reconstruction of multimodal distributions (Béreux et al., 2023).
-Although this method is effective for evaluating model quality, it is too computationally intensive to be used in training, even if it leads to models with good equilibrium properties.
-Other optimized MCMC methods, such as the Parallel Tempering (PT) algorithm  (Hukushima & Nemoto, 1996) (in which the model is simulated in parallel at different temperatures which are exchanged once in a while using the Metropolis rule), have been proposed to improve the training process (Salakhutdinov, 2009; Desjardins et al., 2010) and show important improvements in certain datasets (Krause et al., 2020).
+Training EBMs by maximizing log-likelihood has long been a challenge (lecun2006tutorial; song2021train).
+The introduction of contrastive divergence (hinton2002training) popularized EBMs by using short MCMC runs initialized on minibatch examples, but this approach often produces models with poor equilibrium properties, limiting their effectiveness as generative models (salakhutdinov2008quantitative; desjardins2010tempered; decelle2021equilibrium).
+Persistent contrastive divergence (PCD) (tieleman2008training) improves gradient estimation by maintaining a continuous MCMC chain between updates, acting like a slow annealing process.
+However, PCD struggles with clustered data, as the chain’s properties can drift from the equilibrium measure, degrading the model (bereux2023learning).
+This issue, tied to phase coexistence, can be mitigated with constrained MCMC methods, leveraging order parameters from the singular value decomposition of the RBM coupling matrix to reconstruct multimodal distributions (bereux2023learning).
+While effective for model evaluation, this approach is too computationally intensive for practical training, despite producing models with strong equilibrium properties.
+The population annealing algorithm, which adjusts the statistical weights of parallel chains after each parameter update, has been proposed as an alternative (krause2018population).
+Similarly, reweighting chains using non-equilibrium physics concepts like the Jarzynski equation has been explored (carbone2024efficient).
+However, both methods face challenges with highly structured data, where frequent reweighting (duplicating only a few samples) skews the equilibrium measure due to an insufficient number of independent chains.
+To mitigate this, more sampling steps or lower learning rates are required, leading to excessively long training times to maintain equilibrium.
+In our case, applying these methods to our datasets resulted in training that was either too unstable or prohibitively slow when auto-adjusting learning rates and Gibbs steps.
+Another approach uses EBMs as corrections for flow-based models (nijkamp2022mcmc), simplifying sampling but sacrificing the interpretability of the energy function, which is central to our work.
+Alternatively, evolving flow models can act as fast sampling proposers for EBMs (grenioux2023balanced), but this requires training two separate networks and leads to lower acceptance rates as the EBM becomes more specialized.
+The Parallel Tempering (PT) algorithm  (hukushima1996exchange), in which the model is simulated in parallel at different temperatures which are exchanged once in a while using the Metropolis rule, has been proposed to both speed up the sampling and to improve the training process (salakhutdinov2009learning; desjardins2010tempered) showing important improvements in certain datasets (krause_algorithms_2020).
 However, PT is expensive because it requires sampling at multiple temperatures while using samples from only one model.
-It’s also often ineffective with highly clustered data due to first-order phase transitions in EBMs, where modes disappear abruptly at certain temperatures, as discussed by Decelle & Furtlehner (2021a).
+It’s also often ineffective with highly clustered data due to first-order phase transitions in EBMs, where modes disappear abruptly at certain temperatures, as discussed by (decelle2021exact).
 This leads to sharp drops in acceptance rates for exchanges between temperatures near the transition.
@@
 In contrast, continuous transitions, as seen during the training trajectory, cause modes to gradually separate as the model parameters are changed.
-The population annealing algorithm, which adjusts the statistical weights of parallel chains during learning each time the parameters are updated, has been proposed as an alternative approach (Krause et al., 2018).
-Reweighting chains using non-equilibrium physics concepts like the Jarzynski equation has also been explored (Carbone et al., 2024).
-However, both methods struggle with highly structured data, as frequent reweighting (which involves duplicating only a few samples) biases the equilibrium measure due to an insufficient number of effective independent chains.
-To prevent chains from becoming too correlated around transition regions during training, either more sampling steps or a lower learning rate is required, which leads to very long training times to ensure equilibrium.
-In our case, we tried to used both methods with our datasets but the training becomes too unstable or two slow if learning rate and number of Gibbs steps are auto adjusted.
-Another approach uses EBMs as corrections for flow-based models that are easier to sample (Nijkamp et al., 2022), simplifying learning but sacrificing the interpretability of the energy function, a key goal of this work.
-Alternatively, evolving flow models can serve as fast sampling proposers for the EBM (Grenioux et al., 2023), but this requires training two separate networks and the acceptance of the proposed moves is significantly reduced as the EBM becomes more specialized.
-A recent method for RBMs, called “stacked tempering” (Roussel et al., 2023), significantly accelerates sampling by training smaller RBMs using latent variables from previous models.
-This allows for fast updates through a PT-like algorithm, and the authors demonstrated that it is much faster than standard PT.
-However, while this method is effective for sampling a trained model, it is not suitable for learning, as it requires sequentially training multiple models stacked on top of each other.
-Also for RBMs, it has recently been shown in a theoretical work that it should be possible to train a low-rank RBM that accurately reproduces the statistics of the data projected along the $d$ first data principal directions through a convex and very fast optimization process (see Decelle & Furtlehner (2021a) and the discussion below) using a mapping to another model, the so-called Restricted Coulomb Machine.
+The “Stacked Tempering” algorithm (roussel2023accelerated), recently proposed for RBMs, accelerates sampling by training smaller RBMs with latent variables from previous models, enabling fast updates via a PT-like algorithm.
+While significantly faster than PT, it is unsuitable for learning, as it requires sequential training of multiple stacked models.
+Separately, a theoretical study (decelle2021exact) demonstrated that a low-rank RBM can be trained via a convex, fast optimization process to reproduce the statistics of the data projected along the $d$ first data principal directions through a convex and very fast optimization process (see (decelle2021exact) and the discussion below) using a mapping to another model, the so-called Restricted Coulomb Machine.
 This low-rank model can be seen as a good approximation to the correct RBM needed to describe the data, and has the nice property that it can be efficiently sampled via a static Monte Carlo process.
-While this work showed the efficiency of this training process to describe simple synthetic datasets embedded in only one or two dimensions, the use of this algorithm to obtain reliable low-rank RBMs with real data requires further developments that will be discussed in this paper, in particular: (i) the training of a bias, which is crucial to gain almost for free an additional dimension, and to deal with image datasets, and (ii) a tuning of the RBM fast sampling process, which was not properly investigated in the original paper.
-In this paper, we show how starting the RBM training with a low-rank RBM in very clustered datasets can drastically improve the quality of the models by circumventing the initial dynamical arrest effects that occur during the training process, at a very similar computational cost.
-We also show that the training trajectory can be leveraged to estimate the log-likelihood at a very little computational effort, so that both the training and test set log-likelihood can be easily and reliably computed online during training.
-Furthermore, we propose an efficient sampling method, which we call Parallel Trajectory Tempering (PTT).
-This method outperforms optimized MCMC methods such as the traditional Parallel Tempering (PT) or the “Stacked Tempering” (Roussel et al., 2023) only requiring storing the parameters of the model at a few stages during training.
-By combining these strategies, we are able to train equilibrium models that accurately capture different modes in datasets prone to mode collapse using standard methods, easily evaluate RBMs during the training process, and easily sample from multimodal distribution.
+While this work showed the efficiency of this training process to describe simple synthetic datasets embedded in only one or two dimensions, the use of this algorithm to obtain reliable low-rank RBMs with real data requires further developments that will be discussed in this paper, in particular: (i) the training of a bias, which is crucial to gain almost for free an additional dimension and to deal with image datasets, and (ii) a tuning of the RBM fast sampling process, which was not properly investigated in the original paper.
+In this paper, we show that the training trajectory can be exploited to estimate log-likelihood with minimal computational cost, enabling reliable online computation.
+We also introduce Parallel Trajectory Tempering (PTT), an efficient sampling method that outperforms optimized MCMC approaches like Parallel Tempering (PT) and Stacked Tempering (roussel2023accelerated), requiring only a few model parameters stored at different stages of the training.
+Additionally, we demonstrate that initializing RBM training with a low-rank RBM in highly clustered datasets mitigates the effect of initial dynamical arrests, significantly improving model quality at a comparable cost.
+Combining these strategies allows us to train equilibrium models that capture diverse modes in datasets prone to mode collapse, efficiently evaluate RBMs during training, and sample from multimodal distributions.
 3 The Restricted Boltzmann Machine
-The RBM is composed by $N_{\mathrm{v}}$ visible nodes and $N_{\mathrm{h}}$ hidden nodes.
-In our study, we primarily use binary variables $\{0,1\}$ or ±$\pm 1$ for both layers.
-The two layers (visible and hidden) interact via a weight matrix $\bm{W}$, with no direct couplings within a given layer.
-Variables are also adjusted by visible and hidden local biases, $\bm{\theta}$ and $\bm{\eta}$, respectively.
-The Gibbs-Boltzmann distribution for this model is expressed as
+The RBM consists of $N_{\mathrm{v}}$ visible nodes and $N_{\mathrm{h}}$ hidden nodes.
+In our study, we primarily use binary variables $\{0,1\}$ or $\pm 1$ for both layers.
+The visible and hidden layers interact through a weight matrix $\bm{W}$, with no intra-layer couplings.
+Each variable is further influenced by local biases, $\bm{\theta}$ for visible units and $\bm{\eta}$ for hidden units.
+The Gibbs-Boltzmann distribution for this model is given by
 $p(\bm{v},\bm{h})=\frac{1}{Z}\exp\left[-\mathcal{H}(\bm{v},\bm{h})\right]\text{ where }\mathcal{H}(\bm{v},\bm{h})=-\sum_{ia}v_{i}W_{ia}h_{a}-\sum_{i}\theta_{i}v_{i}-\sum_{a}\eta_{a}h_{a},$ (1)
@@
 For a training dataset $\mathcal{D}=\{{\bm{v}}^{(m)}\}_{m=1,\dots,M}$, the log-likelihood gradient is given by
-$\frac{\partial\mathcal{L}}{\partial w_{ia}}$ $=\langle v_{i}h_{a}\rangle_{\mathcal{D}}-\langle v_{i}h_{a}\rangle_{\rm RBM},\;\frac{\partial\mathcal{L}}{\partial\theta_{i}}=\langle v_{i}\rangle_{\mathcal{D}}-\langle v_{i}\rangle_{\rm RBM},\;\frac{\partial\mathcal{L}}{\partial\eta_{a}}=\langle h_{a}\rangle_{\mathcal{D}}-\langle h_{a}\rangle_{\rm RBM},$ (2)
-where $\left\langle{\cdot}\right\rangle_{\mathcal{D}}$ denotes the average with respect to the entries in the dataset, and $\left\langle{\cdot}\right\rangle_{\mathrm{RBM}}$ with respect to $p(\bm{v},\bm{h})$.
-Since $Z$ is intractable, the model averages in the gradient are typically estimated using $N_{\mathrm{s}}$ independent MCMC processes, and observable averages $\left\langle{o(\bm{v},\bm{h})}\right\rangle_{\mathrm{RBM}}$ are replaced by $\sum_{r=1}^{R}o(\bm{v}^{(r)},\bm{h}^{(r)})/R$, with $(\bm{v}^{(r)},\bm{h}^{(r)})$ being the last configurations reached with each of the $r=1,\ldots,R$ parallel chains.
-To obtain reliable estimates, it should be ensured that each of the Markov chains mix well before each parameter update.
-However, ensuring equilibrium at each update is impractical, slow and tedious.
-The common use of non-convergent MCMC processes is the cause of most difficulties and weird dynamical behaviors encountered in training RBMs as discussed in Decelle et al.
-(2021).
+$\textstyle\frac{\partial\mathcal{L}}{\partial W_{ia}}$ $=\langle v_{i}h_{a}\rangle_{\mathcal{D}}-\langle v_{i}h_{a}\rangle_{\rm RBM},\;\frac{\partial\mathcal{L}}{\partial\theta_{i}}=\langle v_{i}\rangle_{\mathcal{D}}-\langle v_{i}\rangle_{\rm RBM},\;\frac{\partial\mathcal{L}}{\partial\eta_{a}}=\langle h_{a}\rangle_{\mathcal{D}}-\langle h_{a}\rangle_{\rm RBM},$ (2)
+where $\langle f(\bm{v},\bm{h})\rangle_{\mathcal{D}}=M^{-1}\sum_{m}\sum_{\{\bm{h}\}}f(\bm{v}^{(m)},\bm{h})p(\bm{h}|\bm{v}^{(m)})$ denotes the average with respect to the entries in the dataset, and $\left\langle{f(\bm{v},\bm{h})}\right\rangle_{\mathrm{RBM}}$ with respect to $p(\bm{v},\bm{h})$.
+Since $Z$ is intractable, gradient estimates typically rely on $N_{\mathrm{s}}$ independent MCMC processes, replacing observable averages $\left\langle{o(\bm{v},\bm{h})}\right\rangle_{\mathrm{RBM}}$ with $\sum_{r=1}^{R}o(\bm{v}^{(r)},\bm{h}^{(r)})/R$, where $(\bm{v}^{(r)},\bm{h}^{(r)})$ are the final states of $R$ parallel chains.
+Reliable estimates require proper mixing before each parameter update, but enforcing equilibrium at every step is impractical and slow.
+The widespread use of non-convergent MCMC processes underlies most training difficulties and anomalous dynamics in RBMs, as discussed in (decelle2021equilibrium).
 In order to minimize out-of-equilibrium effects, it is often useful to keep $R$ permanent (or persistent) chains, which means that the last configurations reached with the MCMC process used to estimate the gradient at a given parameter update $t$, $\bm{P}_{t}\equiv\{(\bm{v}^{(r)}_{t},\bm{h}^{(r)}_{t})\}_{r=1}^{R}$, are used to initialize the chains of the subsequent update $t+1$.
-This algorithm is typically referred as PCD.
-In this scheme, the process of training can be mimicked to a slow cooling process, only that instead of varying a single parameter, e.g.
-the temperature, a whole set of parameters $\bm{\Theta}_{t}=(\bm{w}_{t},\bm{\theta}_{t},\bm{\eta}_{t})$ are updated at every step to $\bm{\Theta}_{t+1}=\bm{\Theta}_{t}+\gamma\bm{\nabla}\mathcal{L}_{t}$ with $\bm{\nabla}\mathcal{L}_{t}$ being the gradient in equation 2 estimated using the configurations in $\bm{P}_{t}$, and $\gamma$ being the learning rate.
-Typical MCMC mixing times in RBMs are short at the beginning of training but increase as it progresses (Decelle et al., 2021).
-Recent analytical and numerical studies show that the mixing times exhibit sharp, localized peaks when the model encodes a new mode during training (Bachtis et al., 2024).
-This phenomenon, known in physics as critical slowing down, is related to the system going through critical phase transitions.
-For RBMs, the first phase transitions during maximum likelihood training are associated with the encoding of the principal components of the dataset into the singular vectors of the weight matrix $\bm{W}$ (Decelle et al., 2017; 2018; Decelle & Furtlehner, 2021b; Bachtis et al., 2024).
-Since these directions can be determined before training by performing PCA on the data, one can pre-train an RBM to encode them accurately.
-This avoids crossing the initial phase transitions encountered during the standard training and the associated degradation of the gradient calculation.
-4 The low-rank RBM pretrained
-In Decelle & Furtlehner (2021a), it was shown that it is possible to train exactly an RBM containing a reduced number of modes in the weight matrix $\bm{W}$ by exploiting a mapping between the RBM and a Restricted Coulomb Machine and solving a convex optimization problem, see the SI A.2.
-In other words, it is possible to train an RBM whose coupling matrix has this simplified form
-$\bm{W}=\sum_{\alpha=1}^{d}w_{\alpha}\bar{\bm{u}}_{\alpha}\bm{u}_{\alpha}^{\top},\qquad\text{with}\qquad(\bm{u}_{\alpha},\bar{\bm{u}}_{\alpha})\in{\mathbb{R}}^{N_{v}}\times{\mathbb{R}}^{N_{h}},$ (3)
-where $\{w_{\alpha}\}$ are the singular values of the coupling matrix and the right singular vectors $\{\bm{u}_{\alpha}\}_{\alpha=1}^{d}$ correspond exactly to the first $d$ principal directions of the data set.
-Under this assumption, $p(\bm{v})$ a function of only $d$ order parameters, given by the magnetizations along each of the $\boldsymbol{u}_{\alpha}$ components, i.e.
-$m_{\alpha}(\bm{v})=\bm{u}_{\alpha}\cdot\bm{v}/\sqrt{{N_{\mathrm{v}}}}$.
-In particular,
-$H(\bm{v})=-\sum_{\alpha=0}^{d}\theta_{\alpha}m_{\alpha}-\sum_{a}\log\cosh\left(\sqrt{N_{\mathrm{v}}}\bar{u}_{a}\sum_{\alpha=1}^{d}w_{\alpha}m_{\alpha}+\eta_{a}\right)=\mathcal{H}(\bm{m}(\bm{v})),$ (4)
-where $\bm{m}=\left(m_{1},\ldots,m_{\alpha}\right)$ and $\theta_{\alpha}$ is the projection of the visible bias on the extended $\bm{u}$ matrix.
-In order to initialize the visible bias as suggested in Montavon & Müller (2012), a direction $\bm{u}_{0}$ is used for the magnetization $m_{0}$ that is only present in the bias term.
-This is an improvement of this work with respect to Decelle & Furtlehner (2021a), as it allows us to gain almost for free one additional direction.
-The optimal parameters can be determined by solving a convex optimization problem, detailed in SI A.2.
-Once the model is trained, we obtain a probability distribution $p(\bm{m})$ in a much lower-dimensional space than the original $p(\bm{v})$, which can be easily sampled globally using the inverse transform sampling method.
-However, since this approach requires discretizing the $\bm{m}$-space for both training and generation, it becomes impractical for intrinsic space dimensions $d>4$ due to memory constraints.
-These low-rank RBMs are trained to replicate the statistics of the dataset projected onto its first $d$ principal components.
-Despite their simplicity, the low-rank models can still generate an approximate version of the dataset, as shown in Supplemental Figure 15 for the five datasets previously presented.
-In the initial stage of the standard RBM learning process, the model encodes the strongest PCA components of the data through multiple critical transitions (Decelle et al., 2017; 2018; Bachtis et al., 2024).
-Pre-training with the low-rank construction allows us to bypass these transitions and avoid out-of-equilibrium effects caused by critical slowing down associated to these transitions, which becomes prohibitive in the case of very clustered datasets.
-Once the main directions are incorporated, training can efficiently continue with standard algorithms like PCD, as the mixing times of pre-trained machines tend to be much shorter that at the transitions.
-Figure 2: We compare the equilibrium samples generated by RBMs trained on the MNIST01, Mickey, HGD, Ising2D and CelebA datasets using two different training schemes: PCD, and PCD initialized on low-rank RBMs.
-To assess the fitting of the modes, we show a scatter plot of the projections of the samples in the first two principal directions of each dataset, together with the histograms along each direction.
-Black dots and histograms refer to the original data distribution, while the red color is used to represent the generated data.
-In D we show the evolution of the log-likelihood in time for the two training methods computed with the online AIS trajectory.
-In continuous lines we show the train and in dashed, the test.
-Results are obtained by averaging with respect to 10 trainings initialized at the same low-rank RBM, and the standard deviation between all is shown in a shade around the line, which is hardly seen due to the small variation from run to run.
-Overcoming these transitions has dramatic implications for the quality of the models we can train and how accurately they reproduce the statistics of the data, especially for short training times.
-In Fig.
-2, we show for 5 datasets the equilibrium samples drawn from 2 RBMs trained with identical number of samples, minibatch size, $k=100$ Gibbs steps, and learning rate $\gamma=0.01$, but different training strategies.
-In particular, we consider an RBMs trained from scratch with the standard PCD (Tieleman, 2008), and a final machine trained with PCD and pre-trained with a low-rank RBM.
-We also considered machines trained with the recently proposed Jarzynski reweighing method (Carbone et al., 2023) (see SI E for our specific implementation in the RBM), but we do not include the analysis here because in most of the datasets the training method eventually becomes very unstable reaching very poor results.
-In all datasets, the quality of the generated samples and the test log-likelihood are significantly better when pre-training is used.
-We see that the pre-training+PCD training perfectly balances the different modes (here we show the projections on the first 2 PCA directions, but the agreement extends further to several other principal directions).
-We can also compare the log-likelihood of all 5 datasets and find that the pre-trained RBM achieves higher values.
-In all cases, the log-likelihood is measured using the annealed importance sampling trajectory (Tr-AIS) method, and the samples are drawn using the trajectory PT (PTT) algorithm.
-Both algorithms will be described in the next section 5.
-Already from the scatter plots, we see that the pre-training is key in obtaining models where all modes are properly balanced in clustered datasets, but it’s also beneficial in the maximum test log-likelihood we can achieve.
-We observe that the pretraining is not particularly useful when the dataset is not much clustered, as it is the case of CelebA and the entire MNIST dataset as shown in Fig.
-16 of the SI I.
-This result can be easily understood as the encoding of the first modes of the data do not involve splitting the distribution into two isolated modes, something that happens in the case of clustered data.
-5 Trajectory annealing
-The evolution of parameters during training can be seen as a smoother annealing process compared to the standard temperature annealing used in algorithms like annealing importance sampling (AIS) for log-likelihood estimation or the PT algorithm for sampling.
-This is because it avoids the first-order transitions that typically occur with temperature in clustered datasets.
-However, these methods can still be applied to trajectory annealing by using the parameters from different stages of training, rather than scaling the model parameters by a common factor $\beta_{i}$, as is done in temperature-based protocols.
+This algorithm is typically referred to as PCD.
+In this scheme, the process of training can be mimicked as a slow cooling process, only that instead of varying a single parameter, e.g.
+the temperature, a whole set of parameters $\bm{\Theta}_{t}=(\bm{w}_{t},\bm{\theta}_{t},\bm{\eta}_{t})$ is updated at every step to $\bm{\Theta}_{t+1}=\bm{\Theta}_{t}+\gamma\bm{\nabla}\mathcal{L}_{t}$ with $\bm{\nabla}\mathcal{L}_{t}$ being the gradient in equation 2 estimated using the configurations in $\bm{P}_{t}$, and $\gamma$ being the learning rate.
+Typical MCMC mixing times in RBMs are initially short but increase as training progresses (decelle2021equilibrium).
+Studies reveal sharp mixing time peaks when the model encodes new modes (bachtis2024cascade), a phenomenon known as critical slowing down hohenberg1977theory, associated with critical phase transitions.
+In RBMs, these transitions are related with the alignment of the singular vectors of $\bm{W}$ with the dataset’s principal components (Decelle_2017; decelle2018thermodynamics; decelle2021restricted; bachtis2024cascade).
+However, these dominant modes can be identified via PCA before training and efficiently encoded through convex optimization, circumventing standard maximum likelihood training.
+This approach prevents gradient degradation due to diverging mixing times at the first second-order transitions.
+The advantages of using the low-rank construction from decelle2021exact as a pretraining method will be discussed in Section 5.
+4 Trajectory annealing paradigm
+One major challenge with structured datasets is quantifying the model’s quality, since sampling the equilibrium measure of a well-trained model is often too time-consuming.
+This affects the reliability of generated samples and indirect measures such as log-likelihood (LL) estimation through Annealing Importance Sampling (AIS) (krause_algorithms_2020), making them inaccurate and meaningless.
+The evolution of parameters during training can be seen as a smoother annealing process compared to the standard temperature annealing used in algorithms like annealing importance sampling (AIS) for LL estimation or in sampling algorithms like simulated annealing or PT.
+This is because it avoids the first-order transitions that typically occur with temperature in clustered datasets, see SI H for more detailed explanations of this phenomenon in clustered datasets.
+However, these methods can still be applied to trajectory annealing using the parameters of different stages of training, rather than scaling the model parameters by a common factor $\beta_{i}$, as is done in temperature-based protocols.
 4.1 Trajectory Annealing importance sampling
-The standard method for computing the log-likelihood (LL) in the literature is Annealed Importance Sampling (AIS) (Neal, 2001).
-For a detailed review of its application in RBMs, see Krause et al.
-(2020).
-In this approach, the partition function of a model with parameters $\bm{\Theta}=\{\bm{w},\bm{\theta},\bm{\eta}\}$ is estimated by gradually increasing the inverse temperature $\beta$ from 0 to 1 over $N_{\beta}$ steps.
-At $\beta=0$ (where $p_{0}=p_{\mathrm{ref}}$), the partition function $Z_{0}$ is tractable, enabling the estimation of $Z_{i}$ at neighboring $\beta_{i}$ values through path sampling methods.
-The reference configuration $p_{\mathrm{ref}}$ is commonly taken uniform, but improvements have been reported in RBM when fixing it to the independent site distribution that describes the empirical center of the dataset.
-For a detailed description of the algorithm, see SI D.
+The standard method for log-LL estimation in the literature is AIS (neal2001annealed); see (krause_algorithms_2020) for a review on its application to RBMs.
+A detailed algorithm description is provided in SI D.
+AIS estimates the partition function of a model with parameters $\bm{\Theta}=\{\bm{W},\bm{\theta},\bm{\eta}\}$ by gradually increasing the inverse temperature $\beta$ from 0 to 1 over $N_{\beta}$ steps.
+At $\beta=0$, where $p_{0}=p_{\mathrm{ref}}$, the partition function $Z_{0}$ is tractable, allowing estimation of $Z_{i}$ at neighboring $\beta_{i}$ values via path sampling.
+The reference distribution $p_{\mathrm{ref}}$ is typically uniform, but improved performance has been reported when using the independent site distribution centered on the dataset, see SI D for details.
 AIS is quite costly because it requires a large number of annealing steps $N_{\beta}$ for reliable estimations, making it impractical to compute the log-likelihood online during training.
@@
 The results for the HGD dataset are presented in Fig.
-3, while similar qualitative results for the other datasets can be found in the SI, see Figs.
-11, 12, 13 and 14.
+2, while similar qualitative results for the other datasets can be found in the SI, see Figs.
+15, 16, 17 and 18.
 In Fig.
-3, we display the estimation errors in the LL using three different methods: standard temperature Annealed Importance Sampling (AIS) (panel A), AIS with a reference distribution (panel B), and Tr-AIS (panel C).
+2, we display the estimation errors in the LL using three different methods: standard temperature AIS (panel A), AIS with non-flat reference distribution (panel B), and Tr-AIS (panel C).
 For the temperature AIS estimates, we show results for different values of $N_{\beta}$ in distinct colors.
 In panel C, we compare the online estimation in black with the offline Tr-AIS estimation, using the same models analyzed in panels A and B.
-Additionally, we present the error in LL estimation using a few machines and the trajectory parallel tempering (PTT), which will be detailed in the next section.
+Additionally, we present the error in LL estimation using just a few machines saved during the training and the trajectory parallel tempering (PTT), which will be detailed in the next section.
 Overall, the online Tr-AIS estimation consistently outperforms the other methods, and only the PTT estimation yields similar results in offline analyses.
-Figure 3: Comparison of the error in (LL) estimation (with respect to the exact value) using different methods: AIS (left), AIS with a reference distribution (middle), and Tr-AIS (ours, right).
-For Tr-AIS, we present three approaches to computing it: online during training, offline using the models saved during training, and with PTT.
-The RBM was pretrained and trained with a total of 20 hidden nodes (allowing for an exact computation of the LL) for 10,000 gradient steps.
-PTT uses a subset of saved models such that the acceptance rate between two consecutive models is 0.25.
-5.2 Standard Gibbs sampling vs. Parallel Trajectory Tempering
-One major challenge with structured datasets is quantifying the model’s quality, since sampling the equilibrium measure of a well-trained model is often too time-consuming.
-This affects the reliability of generated samples and indirect measures as log-likelihood’s estimation through Annealing Importance Sampling (AIS) (Krause et al., 2020), making them inaccurate and meaningless.
+Figure 2: Comparison of LL estimation error (relative to the exact value) across different methods: AIS (A), AIS with a reference distribution fixed to the independent site distribution matching the dataset’s empirical center (B), and Tr-AIS (ours, C).
+For Tr-AIS, we evaluate three approaches: online during training, offline using saved models, and with PTT.
+The RBM, pretrained and trained with 20 hidden nodes (allowing exact $L$ computation), for $10^{4}$ gradient steps on the HGD dataset.
+PTT selects a subset of saved models ensuring a 0.25 acceptance rate between consecutive models.
+Lines show the mean $L$ over 10 independent runs, with shaded areas representing one standard deviation.
+4.2 Exploiting the training trajectory for sampling
+The trajectory annealing paradigm can also be exploited to significantly accelerate the sampling of equilibrium configurations, as we show in Fig.
+3.
+Before presenting our parallel trajectory annealing (PTT) algorithm, we first discuss the limitations of standard methods like Alternating Gibbs Sampling (AGS), and introduce the standard parallel tempering (PT) algorithm.
+The AGS procedure involves iteratively alternating between two steps: conditioning on all hidden variables given fixed visible variables, and then conditioning on all visible variables given fixed hidden variables.
 Figure 3: Comparison of the performance of different MCMC sampling methods on RBMs trained with the pretraining+PCD procedure on the MNIST01 (row-1), the HGD (row-2) and Ising 2D datasets (row-3).
-In A and B columns, we show the trajectory of two independent Markov chains (red and orange) iterated for $10^{4}$ MCMC steps using either PTT or alternate Gibbs sampling (AGS), projected onto the first two principal components of the dataset.
+In A and B columns, we show the trajectory of two independent Markov chains (red and orange) iterated for $10^{4}$ MCMC steps using either PTT or AGS, projected onto the first two principal components of the dataset.
 The position of the chains is plotted every 10 MCMC steps.
 The black contour represents the density profile of the dataset.
-In column C, we show the average number of jumps between the two regions separated by the dashed grew line in the PCA plots using different MCMC methods: Alternative Gibbs Sampling (AGS), Parallel Tempering (PT) (Hukushima & Nemoto, 1996), without or with reference configuration (PTref) (Krause et al., 2020), the Stacked Tempering (Roussel et al., 2023), and the Trajectory Parallel Tempering (PTT) proposed in this work.
+In column C, we show the averaged number of jumps between the two regions separated by the dashed grey line in the PCA plots using different $MC$ methods: AGS, Parallel Tempering (PT) (hukushima1996exchange), without or with reference configuration (PTref) (krause_algorithms_2020), the Stacked Tempering (roussel2023accelerated), and the Trajectory Parallel Tempering (PTT) proposed in this work.
 The average is calculated over a population of 1000 chains and the shadow around the lines indicates the error of the mean.
-To illustrate this problem, let us consider the MNIST01, the HGD and the Ising 2D datasets.
-MNIST01 and Ising 2D are strongly bimodal in the PCA and the HGD trimodal, as shown in Fig.
+Problems with the standard AGS method– Let us consider the MNIST01, HGD, and Ising 2D datasets.
+MNIST01 and Ising 2D exhibit strong bimodality in their PCA representations, while HGD is trimodal, as shown in Fig.
 1.
-Let us consider that we want to sample the equilibrium measure of the RBMs trained using low-rank RBM pretraining.
-In order to draw new samples from these models, one would typically run MCMC processes from random initialization and iterate them until convergence.
-The mixing time in such case is controlled by the jumping time between clusters.
-To accurately estimate the relative weight between modes, the MCMC processes must be ergodic, requiring many back-and-forth jumps.
-However, as shown in Figs.
-4–A1-3, Alternate Gibbs sampling (AGS) dynamics are extremely slow, rarely producing jumps even after $10^{4}$ MCMC steps.
+Suppose that we have trained an RBM on each of these datasets, for example using the low-rank RBM pretraining (see Section 5), and we aim to sample the equilibrium measure of these models.
+Typically, this involves running MCMC processes from random initializations and iterating them until convergence.
+The mixing time in such cases is determined by the jumping time between clusters.
+Accurate estimation of the relative weights between modes requires the MCMC processes to be ergodic, necessitating frequent back-and-forth jumps.
+However, as shown in the chains in Figs.
+3–B1-3, AGS dynamics are exceedingly slow, rarely producing jumps even after $10^{4}$ MCMC steps.
 The red curves in Figs.
-4–C1-3 show the mean number of jumps over 1000 independent chains as a function of MCMC steps, indicating that a proper equilibrium generation would require at least $10^{6}-10^{7}$ MCMC steps.
-In order to speed up sampling, we consider other traditional optimized MCMC methods such as the PT (with or without reference), the Stack Tempering (Roussel et al., 2023) and our Parallel Trajectory Tempering (PTT) that will be explained below.
-The evolution of the number of jumps between modes observed with each of these algorithms, as a function of the total number of AGS steps performed (i.e.
-the number of parallel models multiplied by the number of AGS steps performed at each model) is shown in Fig.
-4–C1-3.
-Before discussing our PTT algorithm and the results, let us briefly introduce the PT algorithm.
-5.2.1 Parallel Tempering approaches
-Parallel tempering (PT) is commonly used to simulate multiple systems in parallel at different inverse temperatures $\beta$ (Hukushima & Nemoto, 1996; Marinari & Parisi, 1992), with occasional temperature swaps between models based on the Metropolis rule.
+3–C1-3 illustrate the mean number of jumps over 1000 independent chains as a function of MCMC steps, indicating that achieving proper equilibrium generation would require at least $10^{6}$-$10^{7}$ MCMC steps.
+To accelerate sampling, we therefore explore optimized Monte Carlo methods inspired by Parallel Tempering.
+Trajectory Parallel Tempering– Parallel tempering (PT) is used to simulate multiple systems in parallel at different inverse temperatures $\beta$ (hukushima1996exchange; marinari1992simulated), with occasional temperature swaps between models based on the Metropolis rule.
 Since this move satisfies detailed balance, the dynamics will converge to the Boltzmann distribution $p_{\beta}(\bm{x})\propto\exp(-\beta\mathcal{H}(\bm{x}))$ at each $\beta$.
-However, a more general approach involves simulating multiple systems, each with its own set of parameters.
-We can define a system as $p_{\bm{\Theta}^{t}}(\bm{x})$, where $\bm{\Theta}^{t}$ denotes the parameters of the model indexed by $t$.
-This includes the classic PT case, where temperature scaling is applied, $\bm{\Theta}^{t}=\beta^{(t)}\bm{\Theta}$, with $\bm{\Theta}$ being the parameters of the model to be sampled.
-In this work, we introduce a new algorithm that exchanges parameters of models saved at different stages of learning, using the learning trajectory to define the parameter set $\bm{\Theta}^{t}$, where $t$ refers to a specific point in the training process.
+However, a more general approach involves simulating multiple systems with distincts set of parameters.
+For a set of parameters $\bm{\Theta}^{t}$ we associate the system $p_{\bm{\Theta}^{t}}(\bm{x})$.
+This includes the classic PT case, with $\bm{\Theta}_{\mathrm{PT}}^{t}=\beta{(t)}\bm{\Theta}$ where the temperature ladder is typically defined as $\beta(t)=t/N_{T}$, where $t=0,\ldots,N_{T}-1$, and $N_{T}$ represents the number of systems simulated in parallel.
+Alternatively, recent work (roussel2023accelerated) introduced a variant of this algorithm called Stacked Tempering, which uses a series of nested RBMs, where the hidden configurations of one machine are exchanged with the visible configurations of the next in the stack, analogous to the mechanism in deep belief networks.
+We introduce a novel variant where model parameters from different training phases are exchanged.
+Using the learning trajectory, we define the parameter set $\bm{\Theta}^{t}$, with $t$ indexing models at specific training steps.
 We call this method Parallel Trajectory Tempering (PTT).
-The rule for swapping configurations $\bm{x}=(\bm{v},\bm{h})$ between neighboring machines indexed by $t$ and $t-1$ is given by
-$p_{\mathrm{acc}}(\bm{x}^{t}\leftrightarrow\bm{x}^{t-1})=\min\left(1,\exp\left(\Delta\mathcal{H}^{t}(\bm{x}^{t})-\Delta\mathcal{H}^{t}(\bm{x}^{t-1})\right)\right),$ (5)
+Configuration exchanges $\bm{x}\!=\!(\bm{v},\bm{h})$ between neighboring models at $t$ and t1t\!-\!1 follow the rule:
+$p_{\mathrm{acc}}(\bm{x}^{t}\!\leftrightarrow\!\bm{x}^{t-1})\!=\!\min\left[1,\exp\left(\Delta\mathcal{H}^{t}(\bm{x}^{t})\!-\!\Delta\mathcal{H}^{t}(\bm{x}^{t-1})\right)\right]\>\>\text{with}\>\>\Delta\mathcal{H}_{t}(\bm{x})\!\equiv\!\mathcal{H}_{t}(\bm{x})-\mathcal{H}_{t-1}(\bm{x}),$ (3)
 where again the index $t$ indicates that the Hamiltonian is evaluated using the parameters $\bm{\Theta}^{t}$.
 This move satisfies detailed balance with our target equilibrium distribution $p(\bm{x})=\exp(-\mathcal{H}(\bm{x}))/Z$, ensuring that the moves lead to the same equilibrium measure.
-The reason for using the trajectory is that PT algorithms work very well when a system undergoes a continuous phase transition.
-As discussed in the appendix G, adjusting the inverse temperature can lead to discontinuous transitions for which PT is not known to perform very well, while following the trajectory, the system undergoes several continuous transitions as discussed in Decelle et al.
-(2018); Bachtis et al.
-(2024).
-Our PTT approach relies on the ease of sampling configurations from an initial RBM.
-In classical PCD training, the early “non-specialized” models thermalize quickly because they behave like Gaussian distributions.
-For RCM+PCD, independent configurations can be efficiently sampled from the low-rank RBM using a static Monte Carlo algorithm, as described in Section A.1 of the SI.
-The trajectory flow then significantly speeds up convergence to equilibrium, especially in the presence of a multimodal structure.
-The time interval between successive models is chosen to keep the acceptance probability for exchanges between neighboring models (Eq.
-5) around 0.25 (we observed that the total time (number of AGS multiplied by number of machines) needed to thermalize a PTT chain did not depend on the acceptance fixed in the interval [0.2,0.5].
-In practice, effective models can be saved during training by monitoring the acceptance rate between the last saved model and the current one.
+The reason for using the trajectory is that PT-like algorithms work very well when a system undergoes a continuous phase transition.
+As discussed in the SI G, adjusting the inverse temperature can lead to discontinuous transitions for which PT is not known to perform very well, while following the trajectory, the system undergoes several continuous transitions as discussed in (decelle2018thermodynamics; bachtis2024cascade).
+The pseudocodes for the standard PT and our new PTT algorithms are provided in SI B.1.
+Figs.
+3–A1-3 show the evolution of two independent Markov chains on the most trained model using the PTT sampling, showing a significant increase in the number of jumps compared to Figs.
+3–B1-3.
+We also compare the number of mode jumps observed with PTT with those of other algorithms, as a function of the total number of AGS steps performed.
+For the PT algorithm, we test different values of $N_{T}$ and include the variant with a reference configuration (krause_algorithms_2020).
+For Stacked Tempering, we replicate the method of (roussel2023accelerated).
+The total AGS steps are taken as the product of the number of parallel models and the AGS steps per model, as shown in Fig.
+3–C1-3.
+PTT consistently outperforms all other algorithms in all datasets.
+Our PTT approach leverages the ease of sampling configurations from an initial RBM.
+During classical PCD training, early “non-specialized” models thermalize quickly, behaving like Gaussian distributions.
+The same is true for RBMs pre-trained from the low-rank RBM of decelle2021exact, as independent configurations for the first machine can be efficiently sampled by a static Monte Carlo algorithm (see SI A.1).
+The trajectory flow greatly accelerates equilibrium convergence, particularly in multimodal structures.
+The time interval between successive models is set to maintain an acceptance probability for exchanges (Eq.
+3) around 0.25.
+Notably, the total time required to thermalize a PTT chain (AGS steps × number of machines) is unaffected by acceptance values within [0.2, 0.5] (see SI B.2 for details).
+In practice, effective models (i.e.
+the parameters of the model) can be saved during training by monitoring the acceptance rate between the last saved model and the current one.
 Pre-trained models require far fewer saved models, as many are naturally positioned near major phase transitions.
 The number of models used for each sampling process, along with a detailed description of the algorithm, is provided in the SI.
+PTT offers the advantage of requiring no extra training (unlike Stacked Tempering) or unnecessary temperature simulations and can be fully applied to other energy-based models.
+Instead, the equilibrium samples obtained from different models during sampling can be used to compute the log-likelihood, similar to the standard PT approach (krause_algorithms_2020).
 In Fig.
-4–C1-3, we present the number of jumps between clusters as a function of the total number of AGS steps, where in PT-like schemes this refers to 1 (AGS step + one swap proposal) multiplied by the total number of models run in parallel.
-In all cases, PTT sampling outperforms the other methods, showing particularly remarkable results for HGD.
-Additionally, our method has the advantage of not requiring extra training like Stacked Tempering or simulating unnecessary temperatures.
-Moreover, the equilibrium samples obtained from different models during sampling can be used to compute the log-likelihood, similar to the standard PT approach Krause et al.
-(2020).
+2–right we show that the quality of this estimation is comparable with the online trajectory AIS.
+Another desirable advantage of the PT algorithms is that one can easily check the thermalisation of the MCMC process by investigating the ergodicity of the random walk in models (banos2010nature), as shown in SI B.2.
+In addition, we evaluated other indicators of sample quality and potential overfitting in SI  I.
+Our method consistently provides high-quality samples across all datasets and, thanks to precise likelihood computation, can effectively detect overfitting.
+On the failure of thermal algorithms– Thermal sampling algorithms (such as Simulated Annealing or PT) have difficulties in efficiently sampling models that describe structured data sets.
+This is mainly because in these types of models, remote modes can simply disappear or appear discontinuously when the temperature is slightly changed.
+This phenomenon is known in physics as a discontinuous or first-order phase transition.
+For a detailed introduction to this effect and an illustration of its occurrence when changing the temperature in an RBM toy model, see SI H.
+The presence of this or these temperature transitions can affect the performance of the PT algorithm, as a change between two neighboring temperatures is rarely accepted (and therefore much denser temperature ladders are required), but in multimodal models the problem can be even more harmful and difficult to detect: PT can even forget configurations in certain minority clusters while maintaining high swap acceptances.
+We illustrate this problem in Fig.
+4-A, where we compare the equilibrium distribution $p(\bm{m})$ of a well-trained machine trained with the HGD dataset when sampled at two different temperatures $\beta=0.9$ and $\beta=1$.
+This marginal distribution can be efficiently calculated using the tethered MC algorithm bereux2023learning.
+We can see that from $\beta=1$ to $\beta=0.9$ the East Asian cluster (the upper right cluster) disappears from the measure.
+Since the jump to this cluster is strongly suppressed with AGS, as shown in Fig.
+3-C2, simulating the system at higher temperatures does not help to find this cluster more easily, but prevents it from being found with a standard 1-temperature swap + 1AGS scheme (because it will be never nucleated at low temperatures).
+Indeed, in  4-B we show the distribution of samples generated at $\beta=1$ by a long PT sampling run (with all swap acceptances above 0.8) completely forgetting the Asian cluster.
+One might wonder if this is indeed the right balance between clusters learned by the model, so we checked it by performing a long ($10^{6}$ AGS steps) run initialized with these PT-generated samples, finding that the missing clusters gradually resettle.
+Instead, the PTT samples correctly identify all the clusters and the sampled distribution statistics remain unchanged when a normal AGS run is re-run from that initialisation.
+5 The low-rank RBM pretraining
+Figure 4: On the left: Panel A compares the marginal distribution $p(m)$ along the first two principal components for the pretrained+PCD RBM on the HGD dataset at inverse temperatures $\beta\!=\!1$ and $\beta\!=\!0.9$.
+Panel B presents sampling results at $\beta\!=\!1$ using PT, PTT, and $10^{6}$ AGS steps, initialized from configurations generated by both algorithms.
+Right panels: Panels C and D compare equilibrium samples from RBMs trained with PCD alone vs.
+PCD initialized on low-rank RBMs for MNIST01, Mickey, HGD, Ising2D, and CelebA.
+Scatter plots and histograms of projections onto the first two principal components show data (black) vs.
+generated samples (red).
+Panel E tracks log-likelihood evolution (train: solid, test: dashed) using online Tr-AIS.
+Results, averaged over 10 low-rank RBM initializations, show minimal variance, as indicated by narrow shaded regions.
+In (decelle2021exact), it was shown that an RBM with only a few modes in its $\bm{W}$ can be trained precisely through a mapping to a Restricted Coulomb Machine and solving a convex optimization problem (see SI A.2).
+This approach enables training RBMs with simplified $\bm{W}$:
+$\bm{W}=\sum_{\alpha=1}^{d}w_{\alpha}\bar{\bm{u}}_{\alpha}\bm{u}_{\alpha}^{\top},\qquad\text{with}\qquad(\bm{u}_{\alpha},\bar{\bm{u}}_{\alpha})\in{\mathbb{R}}^{{N_{\mathrm{v}}}}\times{\mathbb{R}}^{{N_{\mathrm{h}}}},$ (4)
+where $\{w_{\alpha}\}$ are the singular values of the coupling matrix and the right singular vectors $\{\bm{u}_{\alpha}\}_{\alpha=1}^{d}$ correspond exactly to the first $d$ principal directions of the data set.
+Under this assumption, $p(\bm{v})$ is a function of only $d$ order parameters, given by the magnetizations along each of the $\boldsymbol{u}_{\alpha}$ components, i.e.
+$m_{\alpha}(\bm{v})=\bm{u}_{\alpha}\cdot\bm{v}/\sqrt{{N_{\mathrm{v}}}}$.
+In particular,
+$H(\bm{v})=-\sum_{\alpha=0}^{d}\theta_{\alpha}m_{\alpha}-\sum_{a}\log\cosh\left(\sqrt{N_{\mathrm{v}}}\sum_{\alpha=1}^{d}\bar{u}_{\alpha a}w_{\alpha}m_{\alpha}+\eta_{a}\right)=\mathcal{H}(\bm{m}(\bm{v})),$ (5)
+where $\bm{m}=\left(m_{1},\ldots,m_{\alpha}\right)$ and $\theta_{\alpha}$ is the projection of the visible bias onto the extended $\bm{u}$ matrix.
+To initialize the visible bias as proposed in (montavon2012deep), we add an additional direction $\bm{u}_{0}$ associated with the magnetization $m_{0}$ generated by the part of the visible bias that is not contained in the intrinsic space of $\bm{W}$.
+This is an improvement over (decelle2021exact) because it effectively adds an additional direction at minimal cost.
+We also found that the addition of the bias term is crucial to obtain reliable low-rank RBMs for image data.
+The optimal parameters are obtained by solving a convex optimization problem (SI A.2).
+This yields a probability distribution $p(\bm{m})$ in a lower-dimensional space than $p(\bm{v})$, which can be efficiently sampled using inverse transform sampling.
+However, discretizing the $\bm{m}$-space limits practicality to intrinsic dimensions $d\leq 4$ due to memory constraints.
+These low-rank RBMs replicate dataset statistics projected onto the first $d$ principal components and, despite their simplicity, can approximate the dataset, as demonstrated in Supplemental Figure 19 for the 5 datasets.
+At the start of RBM training, the model encodes the strongest PCA components through multiple critical transitions (Decelle_2017; decelle2018thermodynamics; bachtis2024cascade).
+Pretraining with a low-rank construction bypasses these transitions, avoiding out-of-equilibrium effects from critical slowing down, which becomes prohibitive in highly clustered datasets.
+Once the main directions are incorporated, mixing times stabilize at much lower values, as barriers between clusters shrink when modes align properly bereux2023learning, making it a more reliable starting point for PCD training.
+However, since mixing times still increase over training, pretraining primarily mitigates poor initialization effects, while issues from insufficient MCMC steps may reappear later.
+Bypassing these transitions significantly improves model quality and data statistics reproduction, especially for short training times.
 In Fig.
-3–right we show that the quality of this estimation is comparable with the online trajectory AIS.
-Another desirable advantage of the PT algorithms is that one can easily check the thermalisation of the MCMC process by investigating the ergodicity of the random walk in models Banos et al.
-(2010).
-In addition, we evaluated other indicators of sample quality and potential overfitting.
-As detailed analysis is in Sect.
-H of the SI.
-Our method consistently provides high-quality samples across all datasets and, thanks to precise likelihood computation, can effectively detect overfitting.
+4, we compare equilibrium samples from two RBMs trained on five datasets with identical settings (see SI C) but different strategies: one trained from scratch using standard PCD (tieleman2008training), the other using PCD after low-rank RBM pretraining.
+We also tested Jarzynski reweighting (carbone2024fast) (see SI E), but results were unstable on most datasets and are omitted.
+Across all datasets, pretraining improves sample quality and test log-likelihood.
+The pre-trained+PCD RBM balances modes effectively, as confirmed by PCA projections, and consistently achieves higher log-likelihood values using online Tr-AIS.
+Samples are drawn with the PTT algorithm, both described in Section 4.1.
+Scatter plots highlight the importance of pretraining for balanced models in clustered datasets, significantly improving test log-likelihood.
+However, for less clustered datasets like CelebA and full MNIST, pretraining offers little benefit, as splitting distributions into isolated modes is unnecessary (see Fig.
+20 in SI J).
 6 Conclusions
-In this work, we tackle the challenges of training high-quality equilibrium RBMs on highly structured datasets.
-We focused on improving the sampling process and accurately evaluating test log-likelihood throughout training.
-Our findings reveal that for datasets that are tightly clustered in their principal component analysis representation, introducing a pre-training phase can enhance training quality.
-This phase involves encoding the principal directions of the data into the model using a convex optimization approach.
-However, this pre-training method, that is specific to RBMs, appears to have limited benefits for less structured datasets or for long trainings, as chains get out of equilibrium, and should be combined in the future with optimized MCMC methods to achieve its maximum capacity.
-Furthermore, we developed a new approach that exploits the progressive learning of features during RBM training to enhance evaluation and sampling.
-This approach enables online and precise computation of log-likelihood and leads to an efficient sampling strategy akin to the standard parallel tempering algorithm.
-Our results demonstrate that both our proposed methods outperform traditional algorithms without the need for simulating models at unnecessary temperatures or training additional models to facilitate sampling.
-Furthermore, both the trajectory AIS method for log-likelihood estimation and the trajectory parallel tempering algorithm are highly versatile, making them applicable to more complex energy-based models.
-Additionally, the low-rank RBM model can serve as an effective pre-initialization step for deeper architectures, and the trajectory PT algorithm is suitable for direct use in any EBM, regardless of complexity.
-7 Code availability
-The code and datasets are available at https://github.com/anonymousrbm/iclr-2025
+In this work, we address the challenges of training, evaluating, and sampling high-quality equilibrium Restricted Boltzmann Machines (RBMs) on highly structured datasets.
+Our primary focus is on improving the sampling process and accurately estimating the log-likelihood during training.
+We introduce a novel approach that leverages the progressive feature learning inherent in RBM training to enhance both evaluation and sampling.
+This method facilitates precise, online computation of log-likelihood and enables an efficient sampling strategy, overcoming the limitations of the standard parallel tempering algorithm.
+Additionally, our results indicate that for datasets with tightly clustered structures in their principal component analysis (PCA) representation, incorporating a pre-training phase can improve training outcomes.
+This pre-training involves encoding the principal directions of the data into the model through a convex optimization technique.
+However, this RBM-specific pre-training method shows limited advantages for less structured datasets or extended training durations, where chains deviate from equilibrium.
+In the latter case, the presented training approach should be combined in the future with optimized MCMC methods to achieve its maximum capacity.
+Code availability
+The code and datasets are available at https://github.com/DsysDML/fastrbm
+Acknowledgments
+Authors acknowledge financial support by the Comunidad de Madrid and the Complutense University of Madrid through the Atracción de Talento program (Refs.
+2019-T1/TIC-13298 & Refs.
+2023-5A/TIC-28934), the project PID2021-125506NA-I00 financed by the “Ministerio de Economía y Competitividad, Agencia Estatal de Investigación” (MICIU/AEI/10.13039/501100011033), the Fondo Europeo de Desarrollo Regional (FEDER, UE) and the French ANR grant Scalp (ANR-24-CE23-1320).
 Supplemental Information
@@
 A simple way to encode these $d$-modes is to parameterize the $w$-matrix as:
-$\bm{w}=\sum_{\alpha=1}^{d}w_{\alpha}\bar{\bm{u}}_{\alpha}\bm{u}_{\alpha}^{\top},\qquad\text{with}\qquad(\bm{u}_{\alpha},\bar{\bm{u}}_{\alpha})\in{\mathbb{R}}^{N_{v}}\times{\mathbb{R}}^{N_{h}},$ (6)
+$\bm{W}=\sum_{\alpha=1}^{d}w_{\alpha}\bar{\bm{u}}_{\alpha}\bm{u}_{\alpha}^{\top},\qquad\text{with}\qquad(\bm{u}_{\alpha},\bar{\bm{u}}_{\alpha})\in{\mathbb{R}}^{{N_{\mathrm{v}}}}\times{\mathbb{R}}^{{N_{\mathrm{h}}}},$ (6)
 where $\bm{u}$ and $\hat{\bm{u}}$ are respectively the right-hand and left-hand singular vectors of $\bm{W}$, the former being directly given by the PCA, while $w_{\alpha}$ are the singular values of $\bm{W}$.
@@
 Newton method which converge really fast since we are in small dimension.
-Note that in practice we will obviously use finite estimates of $\phi$, assuming $N_{v}$ is large enough.
-As a result we get $\bm{\mu}^{*}(\bm{m})$ satisfying implicit equations given by the constraints given at given $N_{v}$:
+Note that in practice we will obviously use finite estimates of $\phi$, assuming $N_{\mathrm{v}}$} is large enough.
+As a result we get $\bm{\mu}^{*}(\bm{m})$ satisfying implicit equations given by the constraints given at given $N_{\mathrm{v}}$}:
 $m_{\alpha}=\frac{1}{\sqrt{{N_{\mathrm{v}}}}}\sum_{i=1}^{{N_{\mathrm{v}}}}u_{i}^{\alpha}\tanh\left({\sqrt{{N_{\mathrm{v}}}}\sum_{\beta=1}^{d}u_{i}^{\beta}\mu^{*}_{\beta}}\right).$ (16)
@@
 In Ref.
-Decelle & Furtlehner (2021a), it has been shown that such RBM can be trained by mean of the Restricted Coulomb Machine, where the gradient is actually convex in the parameter’s space.
+(decelle2021exact), it has been shown that such RBM can be trained by mean of the Restricted Coulomb Machine, where the gradient is actually convex in the parameter’s space.
 It is then possible to do a mapping from the RCM to the RBM to recover the RBM’s parameters.
@@
 A.2 The Restricted Coulomb Machine
-As introduced in Decelle & Furtlehner (2021a), it is possible to exactly train a surrogate model for the RBM, called the Restricted Coulomb Machine (RCM), on a low dimensional dataset without explicitly sampling the machine allowing to learn even heavily clustered datasets.
+As introduced in (decelle2021exact), it is possible to precisely train a surrogate model for the RBM, called the Restricted Coulomb Machine (RCM), on a low dimensional dataset without explicitly sampling the machine allowing to learn even heavily clustered datasets.
 The RCM is an approximation of the marginal distribution of the RBM with $\{-1,1\}$ binary variables:
@@
 We then project both the parameters and variables of the RBM on the first $d$ principal components of the dataset:
-$m_{\alpha}\coloneqq\frac{1}{\sqrt{N_{v}}}\sum_{i=1}^{N_{v}}v_{i}u_{i\alpha},\quad w_{\alpha a}\coloneqq\sum_{i=1}^{N_{v}}w_{ia}u_{i\alpha},\quad\theta_{\alpha}\coloneqq\frac{1}{\sqrt{N_{v}}}\sum_{i=1}^{N_{v}}\theta_{i}u_{i\alpha}$ (20)
-with $\alpha\in\{0,\ldots,d\}$ and $\bm{u}$ the projection matrix of the PCA.
+$m_{\alpha}\coloneqq\frac{1}{\sqrt{{N_{\mathrm{v}}}}}\sum_{i=1}^{{N_{\mathrm{v}}}}v_{i}u_{i\alpha},\quad w_{\alpha a}\coloneqq\sum_{i=1}^{{N_{\mathrm{v}}}}W_{ia}u_{i\alpha},\quad\theta_{\alpha}\coloneqq\frac{1}{\sqrt{{N_{\mathrm{v}}}}}\sum_{i=1}^{{N_{\mathrm{v}}}}\theta_{i}u_{i\alpha}$ (20)
+with $\alpha\in\{1,\ldots,d\}$ and $\bm{u}$ the projection matrix of the PCA.
 The projected distribution of the model is then given by
-$p_{\text{RBM}}(\bm{m})=\frac{\exp\left(N_{v}\left[\mathcal{S}(\bm{m})+\sum_{\alpha=1}^{d}\theta_{\alpha}m_{\alpha}+\frac{1}{N_{v}}\sum_{a=1}^{N_{h}}\log\cosh\left(\sqrt{N_{v}}\sum_{\alpha=1}^{d}m_{\alpha}w_{\alpha a}+\eta_{a}\right)\right]\right)}{Z}$ (21)
+$p_{\text{RBM}}(\bm{m})=\frac{\exp\left({N_{\mathrm{v}}}\left[\mathcal{S}(\bm{m})+\sum_{\alpha=1}^{d}\theta_{\alpha}m_{\alpha}+\frac{1}{{N_{\mathrm{v}}}}\sum_{a=1}^{{N_{\mathrm{h}}}}\log\cosh\left(\sqrt{{N_{\mathrm{v}}}}\sum_{\alpha=1}^{d}m_{\alpha}w_{\alpha a}+\eta_{a}\right)\right]\right)}{Z}$ (21)
 where we ignore the fluctuations related to the transverse directions and $\mathcal{S}[\bm{m}]$ accounts for the non-uniform prior on $\bm{m}$ due to the projection of the uniform prior on $\bm{s}$ for the way to compute it.
@@
 which is valid for $x$ large enough. The probability of the RCM is thus given by:
-$p_{\text{RCM}}(\bm{m})=\frac{\exp\left(N_{v}\left[\mathcal{S}(\bm{m})+\sum_{\alpha=1}^{d}\theta_{\alpha}m_{\alpha}+\sum_{a=1}^{N_{h}}q_{a}\left|\sum_{\alpha=1}^{d}n_{\alpha}m_{\alpha}+z_{a}\right|\right]\right)}{Z}$ (23)
+$p_{\text{RCM}}(\bm{m})=\frac{\exp\left({N_{\mathrm{v}}}\left[\mathcal{S}(\bm{m})+\sum_{\alpha=1}^{d}\theta_{\alpha}m_{\alpha}+\sum_{a=1}^{{N_{\mathrm{h}}}}q_{a}\left|\sum_{\alpha=1}^{d}n_{\alpha}m_{\alpha}+z_{a}\right|\right]\right)}{Z}$ (23)
 where
-$q_{a}=\sqrt{N_{v}\sum_{\alpha=1}^{d}w_{\alpha a}^{2}},\quad n_{a}=\frac{w_{\alpha a}}{\sqrt{\sum_{\alpha=1}^{d}w_{\alpha a}^{2}}},\quad z_{a}=\frac{\eta_{a}}{\sqrt{N_{v}\sum_{\alpha=1}^{d}w_{\alpha a}^{2}}}.$ (24)
+$q_{a}=\sqrt{{N_{\mathrm{v}}}\sum_{\alpha=1}^{d}w_{\alpha a}^{2}},\quad n_{a}=\frac{w_{\alpha a}}{\sqrt{\sum_{\alpha=1}^{d}w_{\alpha a}^{2}}},\quad z_{a}=\frac{\eta_{a}}{\sqrt{{N_{\mathrm{v}}}\sum_{\alpha=1}^{d}w_{\alpha a}^{2}}}.$ (24)
 This can be easily inverted as
-$w_{\alpha a}=\frac{1}{\sqrt{N_{v}}}q_{a}n_{a}\qquad\text{and}\qquad\eta_{a}=q_{a}z_{a},$
+$w_{\alpha a}=\frac{1}{\sqrt{{N_{\mathrm{v}}}}}q_{a}n_{a}\qquad\text{and}\qquad\eta_{a}=q_{a}z_{a},$
 in order to obtain the RBM from the RCM.
@@
 To relax the problem, since we’re in low dimension, we can define a family of hyperplanes $(\bm{n},\bm{z})$ covering the space and let the model only learn the weights of each to the hyperplane.
-We can then discard the ones with a weight low enough for the approximation equation 22 to be bad.
+We can then discard the ones with a low weights to keep the approximation (equation 22) good enough.
 The gradients are given by
@@
 A.3 Learning the bias
-To initialize the model with its bias set to the dataset’s mean, we define the 00-th direction of the longitudinal space as the direction obtained by projecting the configuration space onto the normalized mean vector of the dataset.
+To initialize the model with its bias set to the dataset’s mean, we define the 0-th direction of the longitudinal space as the direction obtained by projecting the configuration space onto the normalized mean vector of the dataset.
 The RBM will only contribute to this direction through the visible bias:
-$H(\bm{m}(v))=-\sum_{\alpha=0}^{N_{v}}\theta_{\alpha}m_{\alpha}-\frac{1}{N_{v}}\sum_{a}\log\cosh\left(\sqrt{N_{v}}\sum_{\alpha=1}^{d}m_{\alpha}w_{\alpha a}+\eta_{a}\right).$ (27)
+$\mathcal{H}(\bm{m}(v))=-\sum_{\alpha=0}^{{N_{\mathrm{v}}}}\theta_{\alpha}m_{\alpha}-\frac{1}{{N_{\mathrm{v}}}}\sum_{a}\log\cosh\left(\sqrt{{N_{\mathrm{v}}}}\sum_{\alpha=1}^{d}m_{\alpha}w_{\alpha a}+\eta_{a}\right).$ (27)
 This implies that this direction is independent of the others within the model.
@@
 where $\Delta\mathcal{H}_{t}(\bm{x})=\mathcal{H}_{t}(\bm{x})-\mathcal{H}_{t-1}(\bm{x})$.
-We continue alternating between the update step and the swap step until a total of $N_{\mathrm{mcmc}}$ steps is reached. The sampling procedure is illustrated in the following pseudo-code:
-Input: Set of models $\{\mathcal{H}_{t}\}$, $t=0,\dots,t_{\mathrm{f}}$, Number of Gibbs steps $k$, Number of MCMC steps $N_{\mathrm{mcmc}}$
-Output: Configurations $\bm{x}_{t}$ for $t=1,\ldots,t_{\mathrm{f}}$
-Initialize: Sample $N$ chains from the RCM model $\bm{x}_{0}^{(0)}\sim\mathrm{RCMSampling}(\mathcal{H}_{0})$
-for $t=1$ to $t_{\mathrm{f}}$ do
-     $\bm{x}_{t}^{(0)}\sim\mathrm{GibbsSampling}(\mathcal{H}_{t},\bm{x}_{t-1}^{(0)},\tilde{k})$
+We continue alternating between the update step and the swap step until a total of $N_{\mathrm{mcmc}}$ steps is reached.
+A comparison of performances between PTT and standard Gibbs sampling is reported in Fig.3.
+B.1 Pseudo-code of PTT vs PT
+We provide here the pseudo code for both Parallel Tempering (Alg.
+1) and Parallel Trajectory Tempering (Alg.
+2) sampling schemes.
+The difference between both algorithms lies in the way the parallel models are selected.
+In the case of Parallel Tempering, an annealing on the temperature is made, whereas for Parallel Trajectory Tempering several models are saved during training.
+Algorithm 1 Parallel Tempering
+$\{\beta_{j}\}_{j=\{1,\ldots,\texttt{N}_{\beta}\}},\bm{\Theta}=\{\bm{w},\bm{\theta},\bm{\eta}\},\texttt{N}_{\texttt{iter}},\texttt{N}_{\texttt{increment}},\texttt{N}_{\texttt{sample}}$
+for $j=1,\ldots,N_{\beta}$ do
+  $\texttt{chains}[j]=\texttt{RandomBinaryChains}(\texttt{N}_{\texttt{sample}})$
 end for
-for $n=1$ to $\lfloor N_{\mathrm{mcmc}}/k\rfloor$ do
-     for $t=1$ to $t_{\mathrm{f}}$ do
-         $\bm{x}_{t}^{(n)}\sim\mathrm{GibbsSampling}(\mathcal{H}_{t},\bm{x}_{t}^{(n-1)},k)$
+for $i=1,\ldots,\texttt{N}_{\texttt{iter}}$ do
+  for $j=1,\ldots,\texttt{N}_{\beta}$ do
+   $\texttt{chains}[j]\leftarrow\texttt{AlternateGibbsSampling}(\texttt{N}_{\texttt{increment}},\beta_{j}\bm{\Theta},\texttt{chains}[j])$
   end for
-     Resample $\bm{x}_{0}^{(n)}\sim\mathrm{RCMSampling}(\mathcal{H}_{0})$
-     for $t=1$ to $t_{\mathrm{f}}$ do
-         Compute acceptance probability
-$p_{\mathrm{acc}}(\bm{x}_{t}^{(n)}\leftrightarrow\bm{x}_{t-1}^{(n)})=\min\left(1,\exp\left(\Delta\mathcal{H}_{t}(\bm{x}_{t}^{(n)})-\Delta\mathcal{H}_{t}(\bm{x}_{t-1}^{(n)})\right)\right)$
-         Swap $\bm{x}_{t}^{(n)}$ and $\bm{x}_{t-1}^{(n)}$ with probability $p_{\mathrm{acc}}(\bm{x}_{t}^{(n)}\leftrightarrow\bm{x}_{t-1}^{(n)})$
+  for $j=1,\ldots,\texttt{N}_{\beta}-1$ do
+   $\texttt{chains}[j],\texttt{chains}[j+1]\leftarrow\texttt{SwapConfigurations}(\texttt{chains}[j],\beta_{j}\bm{\Theta},\texttt{chains}[j+1],\beta_{j+1}\bm{\Theta})$
   end for
 end for
-A comparison of performances between PTT and standard Gibbs sampling is reported in Fig.4.
-B.1 Controlling equilibration times with PTT
+Algorithm 2 Parallel Trajectory Tempering
+$\{\bm{\Theta}_{j}=\{\bm{w}_{j},\bm{\theta}_{j},\bm{\eta}_{j}\}\}_{j=\{1,\ldots,\texttt{N}_{\texttt{model}}\}},\texttt{N}_{\texttt{iter}},\texttt{N}_{\texttt{increment}},\texttt{N}_{\texttt{sample}}$
+for $j=1,\ldots,\texttt{N}_{\texttt{model}}$ do
+  $\texttt{chains}[j]=\texttt{RandomBinaryChains}(\texttt{N}_{\texttt{sample}})$
+end for
+for $i=1,\ldots,\texttt{N}_{\texttt{iter}}$ do
+  for $j=1,\ldots,\texttt{N}_{\texttt{model}}$ do
+   $\texttt{chains}[j]\leftarrow\texttt{AlternateGibbsSampling}(\texttt{N}_{\texttt{increment}},\bm{\Theta}_{j},\texttt{chains}[j])$
+  end for
+  for $j=1,\ldots,\texttt{N}_{\texttt{model}}-1$ do
+   $\texttt{chains}[j],\texttt{chains}[j+1]\leftarrow\texttt{SwapConfigurations}(\texttt{chains}[j],\bm{\Theta}_{j},\texttt{chains}[j+1],\bm{\Theta}_{j+1})$
+  end for
+end for
+B.2 Controlling equilibration times with PTT
 One of the advantages of PT-like approaches is that they allow easy control over thermalization times by studying the random walk in temperatures (or models for the PTT) visited by each run over time, as proposed in Ref.
-Banos et al.
-(2010).
+(banos2010nature).
 From now on, we will refer only to the number of models used in the PTT sampling, denoted as $N_{\mathrm{m}}$.
 In an ergodic sampling process, each run should spend roughly the same amount of time at each of the $N_{\mathrm{m}}$ different model indices, as all models have the same probability.
-The mean temperature/model index is known to be $\left\langle n\right\rangle=(N_{\mathrm{m}}-1)/2$ (assuming the model indices run from 00 to $N_{\mathrm{m}}-1$).
+The mean temperature/model index is known to be $\left\langle n\right\rangle=(N_{\mathrm{m}}-1)/2$ (assuming the model indices run from 0 to $N_{\mathrm{m}}-1$).
 This simplifies the computation of the time-autocorrelation function of the model indices $n_{i}(t)$ visited by each sample $i$ at time $t$:
@@
 $C(t)\sim Ae^{-t/\tau_{\mathrm{exp}}}\quad\text{for large }t,$
-as well as the integrated autocorrelation time, $\tau_{\mathrm{int}}$, using the self-consistent equation Sokal (1997); Amit & Martin-Mayor (2005):
+as well as the integrated autocorrelation time, $\tau_{\mathrm{int}}$, using the self-consistent equation (sokal1997monte; amit2005field):
 $\tau_{\mathrm{int}}=\frac{1}{2}+\sum_{t=0}^{6\tau_{\mathrm{int}}}C(t).$
-These two times provide different insights: $\tau_{\mathrm{exp}}$ gives insight into the time needed to thermalize the simulations (setting the length of the runs to at least $20\tau_{\mathrm{exp}}$ is generally a safe (and very conservative) choice to ensure thermalization Sokal (1997)), while $\tau_{\mathrm{int}}$ gives the time needed to generate independent samples.
+These two times provide different insights: $\tau_{\mathrm{exp}}$ gives insight into the time needed to thermalize the simulations (setting the length of the runs to at least $20\tau_{\mathrm{exp}}$ is generally a safe (and very conservative) choice to ensure thermalization (sokal1997monte)), while $\tau_{\mathrm{int}}$ gives the time needed to generate independent samples.
 To keep track of the fluctuations, we compute a different $C^{m}(t)$ for each $m=1,\ldots,N_{\mathrm{m}}$ system index simulated in parallel in the PTT, where each $C^{m}(t)$ is averaged over 1000 independent realizations (with all runs initialized to the same model index).
@@
 Appendix D Computation of the likelihood using AIS
-To compute the likelihood of the RBMs we used the Annealed Importance Sampling method introduced by Neal (2001) and recently studied in the case of the RBM Krause et al. (2020).
+To compute the likelihood of the RBMs we used the Annealed Importance Sampling method introduced by (neal2001annealed) and recently studied in the case of the RBM (krause_algorithms_2020).
 D.1 Offline AIS
-We follow the method described in Krause et al.
-(2020).
+We follow the method described in (krause_algorithms_2020).
 The AIS estimate is based on creating a sequence of machines, at one end a machine from which we can compute the partition function and at the other end the machine from which we want to compute the partition function.
@@
 For each machine, we denote its set of parameters $\bm{\Theta}_{i}=\{\bm{w}_{i},\bm{\theta}_{i},\bm{\eta}_{i}\}$.
-Following Krause et al.
-(2020), defining a transition rule $T_{i}(\bm{x}_{i-1}\to\bm{x}_{i})$ in order to pass from a configuration $\bm{x}_{i-1}$ of the system $i-1$ to $\bm{x}_{i}$ of the system $i$, that we have the following identity
+Following (krause_algorithms_2020), defining a transition rule $T_{i}(\bm{x}_{i-1}\to\bm{x}_{i})$ in order to pass from a configuration $\bm{x}_{i-1}$ of the system i1i-1 to $\bm{x}_{i}$ of the system $i$, that we have the following identity
 $\frac{Z_{N_{\rm rbm}}}{Z_{0}}=\left\langle\exp\left(-\sum_{i=1}^{N_{\rm rbm}}\mathcal{H}_{i+1}(\bm{x}_{i})-\mathcal{H}_{i}(\bm{x}_{i})\right)\right\rangle_{\rm ForwardPath}.$ (29)
@@
 The RBMs are chosen following the prescription that the exchange rate between configuration (see Eq.
-5) is about $\sim 0.3$).
+3) is about $\sim 0.3$).
 Then the formula Eq.
@@
 Appendix E Training of the RBM using the Jarzynski equation
-In this section, we describe the procedure introduced in Carbone et al.
-(2024) for training energy-based models by leveraging the Jarzynki equation, and we adapt it to the specific case of the RBM.
-At variance with Carbone et al.
-(2024), we introduce an additional assumption on the transition probability of the parameters that will lead us to the the reweighting implemented in population annealing methods Weigel et al.
-(2021).
+In this section, we describe the procedure introduced in (carbone2024efficient) for training energy-based models by leveraging the Jarzynki equation, and we adapt it to the specific case of the RBM.
+At variance with (carbone2024efficient), we introduce an additional assumption on the transition probability of the parameters that will lead us to the the reweighting implemented in population annealing methods (PhysRevE.103.053301).
 In one of its formulations, the Jarzynski equation states that we can relate the ensemble average of an observable $\mathcal{O}$ with the average obtained through many repetitions of an out-of-equilibrium dynamical process.
@@
 In other words, the vast majority of the chains we are simulating are statistically irrelevant, and we expect to get large fluctuations in the estimate of the gradient because of the small effective number of chains contributing to the statistical average.
-A good observable for monitoring this effect is the Effective Sample Size (ESS), defined as Carbone et al.
-(2024)
+A good observable for monitoring this effect is the Effective Sample Size (ESS), defined as (carbone2024efficient)
 $\mathrm{ESS}=\frac{\left(R^{-1}\sum_{r=1}^{R}e^{-W^{(r)}}\right)^{2}}{R^{-1}\sum_{r=1}^{R}e^{-2W^{(r)}}}\in[0,1],$ (38)
@@
 In order to be fair, the comparison is done at a constant number of gradient updates of $10000$.
-In this regime, the machines that perform the best have a learning rate of $\gamma\sim 0.1,0.005$ and about $N_{h}=500,1000$ hidden nodes.
+In this regime, the machines that perform the best have a learning rate of $\gamma\sim 0.1,0.005$ and about ${N_{\mathrm{h}}}=500,1000$ hidden nodes.
 We can expect that smaller learning rates reach the same value of the log-likelihood but for a much longer training.
@@
 where $g(E)$ the number of states with a given energy $E$, and thus $S(E)=\log g(E)$ is the entropy.
-Finally, the free energy is defined as $F(E)=E-S(E)$ and $f(E)=F(E)/N$ is the free energy per variable.
+Finally, the free energy is defined as F(E)=ES(E)F(E)=E-S(E) and $f(E)=F(E)/N$ is the free energy per variable.
 All $Z$, $E$ and $F$ depend on the model parameters.
@@
 In the case of the RBM, the first learning transition can be mapped to a very simple model, the so-called Curie-Weiss model, as recently described in Ref.
-Bachtis et al.
-(2024).
+(bachtis2024cascade).
 The Curie-Weiss model is fully solvable, and for this reason the meaning of a continuous transition is perhaps easier to understand there.
@@
 Case $H=0$:
-we expect that for $\beta$ small each spin behaves as an isolated Rademacher random variables (each spin being ±$\pm 1$ with probability one-half), while for large value of $\beta$, the distribution of the system is dominated by configurations where the variables have (almost) all the same signs.
+we expect that for $\beta$ small each spin behaves as an isolated Rademacher random variables (each spin being $\pm 1$ with probability one-half), while for large value of $\beta$, the distribution of the system is dominated by configurations where the variables have (almost) all the same signs.
 The key mathematical aspect is to study the system in the limit where $N\to\infty$.
@@
 A particular interesting properties is that at $\beta_{c}$, the function $\Omega(m)$ develops a non-analyticity in its second-order derivative, that can be linked with long-range fluctuations.
-For a recent review Kochmański et al.
-(2013), and for more rigorous results Sinai (2014).
+For a recent review (kochmanski2013curie), and for more rigorous results (sinai2014theory).
 Case $H\neq 0$:
@@
 In the infinite size limit, the system will always dominated by the state where $\Omega$ is at its lowest value.
-However, if one prepare a system close to a metastable state, the system will remain trap there until it disappears.
+However, if one prepares a system close to a metastable state, the system will remain trapped there until it disappears.
 Figure 8: Left: The case $H=0$ for various temperature.
@@
 How is this related to the training of the RBM ?
-First, it has been shown in previous works Decelle et al.
-(2018); Bachtis et al.
-(2024) that the RBM’s training undergoes a set of continuous transitions, at least at the beginning of the learning.
+First, it has been shown in previous works (decelle2018thermodynamics; bachtis2024cascade) that the RBM’s training undergoes a set of continuous transitions, at least at the beginning of the learning.
 Hence, changing the parameters of the RBM following the learning trajectories, the transitions are continuous.
-Then in Decelle & Furtlehner (2021a) it is shown that the relaxed version of the RBM, namely the RCM undergoes a first-order transition when changing the temperature.
+Then in (decelle2021exact) it is shown that the relaxed version of the RBM, namely the RCM undergoes a first-order transition when changing the temperature.
 Schematically to understand that, it is enough to see how the parameters $\beta$ acts on the system.
@@
 Changing the temperature tends to produces discontinuous transition where $p_{\beta(t)\bm{\Theta}}(\bm{s})$ can happen to be very far from $p_{\beta(t+\delta t)\bm{\Theta}}(\bm{s})$.
-Appendix H Overfitting and privacy loss as quality indicators
+Appendix H First order transitions in RBMs
+In this section, we examine the appearance of first-order transitions when adjusting the temperature of a well-trained RBM on clustered datasets—a phenomenon that significantly hinders the performance of thermal sampling algorithms.
+We first illustrate this issue using an analytically tractable toy model, followed by an analysis of the first-order temperature transition observed in the RBM trained on the HGD dataset, as discussed in the main text.
+H.1 Theoretical analysis in a simple model
+We propose a simple model of dataset on which we will show that the learned RBM is having a first order transition in temperature.
+We first consider an artificial dataset using a Curie-Weiss model.
+We defined the following Hamiltonian on the variables $s_{i}=\pm 1$
+$\mathcal{H}=-\frac{1}{2{N_{\mathrm{v}}}}\left(\sum_{i}s_{i}\right)^{2},\quad p(\bm{s})=\frac{1}{Z}\exp(-\beta_{T}\mathcal{H}).$
+When the inverse temperature $\beta_{T}$ is below one, the distribution in the $N_{\mathrm{v}}$$\to\infty$ limit is unimodal, with the Boltzmann average of the variables $s$ being $\langle s\rangle=0$.
+For $\beta_{T}>1$, the system becomes bimodal, with $\langle s_{i}\rangle=m$, where $m$ satisfies the self-consistent equation $m=\tanh(\beta_{T}m)$.
+We propose to model this dataset using a Bernoulli-Gauss RBM with one hidden node.
+The visible nodes, $\sigma=\{0,1\}$, follow a Bernoulli distribution, while the hidden node, $\tau$, follows a Gaussian distribution with zero mean and variance $1/{N_{\mathrm{v}}}$, combined with a local visible bias.
+When the inverse temperature $\beta_{T}$ is below one, the distribution in the $N_{\mathrm{v}}$$\to\infty$ limit is unimodal and the average w.r.t.
+the Boltzmann distribution of the values of the variables $s$’s is $\langle s\rangle=0$.
+When $\beta_{T}$ is above one, the system becomes bimodal with $\langle|s_{i}|\rangle=m$ when $m$ is defined by the self-consistent equation $m=\tanh(\beta_{T}m)$.
+We now propose to learn this dataset using a Bernoulli-Gauss RBM with one hidden node, where the visible nodes are $\sigma=\{0,1\}$, and the hidden one $\tau$ is Gaussian distribution of zero mean and variance $1/{N_{\mathrm{v}}}$, and a local visible bias.
+The Hamiltonian is then
+$\mathcal{H}=-\sum_{i}\sigma_{i}W_{i}\tau-\sum_{i}\eta_{i}\sigma_{i},$
+where $W_{i}$ is the weight matrix and $\eta_{i}$ the biases.
+The leaning then consist in converting the binary variable of the Curie-Weiss model into $\{0,1\}$ and to learn the coresponding dataset.
+By a rapid inspection (in the infinite size limit), we can easily show that the optimal learned parameters for the RBM are
+$w_{i}=2\sqrt{\beta_{T}}\text{ and }\eta_{i}=-2\beta_{T}.$
+Now, thanks to the simplicity of the model, we can actually compute the free energy of the model where the Hamiltonian has been rescaled by an annealing temperature $\beta$, we obtain
+$-f(m_{T})$ $=\frac{m_{\tau}^{2}}{2}-N^{-1}\sum_{i}\log\left[1+\exp(\beta w_{i}m_{\tau}+\beta\eta_{i})\right]$
+$\text{ where }m_{\tau}=\frac{1}{N}\sum_{i}\beta w_{i}{\rm sigm}\left[1+\exp(\beta w_{i}m_{\tau}+\beta\eta_{i}\right]$
+In Fig.
+9, we plot $\beta f(m_{T})$ for ten different values of $\beta$ within the range $[0.8,1.05]$.
+It is evident that $\beta=1$ represents the coexistence point, indicating a first-order transition in temperature.
+Note that the probability $p(m)\approx e^{-N\beta f(m_{T})}$ for large $N$, implying that for $\beta>1$, the typical configuration corresponds to $m>1$, whereas for $\beta<1$, it corresponds to $m<1$.
+This clearly shows that even on simple case we observe such transition when doing annealing.
+Figure 9: Free energy $\beta f(m_{\tau})$ for $\beta\in[0.8,1.05]$ for the optimal RBM’s parameters learned on a CW model with $\beta_{T}=1.4$.
+At $\beta=1.0$ the two minima are perfectly adjusted.
+As we change the temperature of the RBM, they get destabilized showing the presence of a first order transition.
+H.2 First order phase transition in temperature on a trained RBM on a real dataset
+We now consider the RBM trained on the HGD dataset using the pretraining and 50 000 updates of PCD training.
+As we vary the temperature, we recover the probability $p(\bm{m})$ learned by the model using the TMC method (bereux2023learning), projected along the second principal component in Fig.
+11 and on both first and second principal components in Fig.
+10.
+While $\beta<1.0$, the probability distribution is dominated by the central mode, equilibrating between all modes $\beta=1.0$.
+As $\beta$ goes above 1.0, the distribution becomes dominated by the two external clusters.
+The effect is clearly visible when sampling the model using the PT algorithm (Fig.
+12).
+The sampling was performed with 200 temperatures with a minimum acceptance rate of 0.78 between two temperatures.
+We see a dataset cluster being ignored with only a few samples in it.
+However, when resampling sampling the model for $10^{6}$ AGS steps starting from the samples obtained through PT, we see a clear change in the distribution, equilibrating with the missing cluster.
+In contrast, performing the same experiment with PTT instead of PT, we observe no shift in the empirical distribution of the samples after $10^{6}$ AGS, as the upper cluster is already sampled.
+Figure 10: Projection of the negative log probability (top) and probability (bottom) learned by the model, constrained to visible magnetizations $m_{0}$ and $m_{1}$, for various values of $\beta$.
+The red dots represent the training data projected onto the first two principal components.
+Figure 11: Negative log probability (left) and probability (right) learned by the model constrained to have a given $m_{1}$ for several values of $\beta$.
+The equilibrium distribution of the model is displayed as the black dotted line ($\beta=1$).
+Figure 12: Comparison of the sampling results of PT (left) and PTT (right).
+The top row corresponds to sample generated with each methods compared with the dataset distribution.
+The bottom row compares samples obtained by performing $10^{6}$ AGS starting from the PT and PTT samples with their starting position.
+Appendix I Overfitting and privacy loss as quality indicators
 In this section, we examine the quality of the samples generated, regarding overfitting and privacy criteria which have been defined for genomic data in particular.
@@
 1–C, to evaluate the ability of various state-of-the-art generative models to generate realistic fake genomes while minimizing privacy concerns (i.e., reducing overfitting).
-Recent studies (Yelmen et al., 2021; 2023) have thoroughly investigated this for a variety of generative models.
+Recent studies (yelmen2021creating; yelmen2023deep) have thoroughly investigated this for a variety of generative models.
 Both studies concluded that the RBM was the most accurate method for generating high-quality and private synthetic genomes.
 The comparison between models relies primarily on the Nearest Neighbor Adversarial Accuracy ($AA_{\rm TS}$) and privacy loss indicators, introduced in Ref.
-Yale et al.
-(2020), which quantify the similarity and the level of ”privacy” of the data generated by a model w.r.t.
+(yale2020generation), which quantify the similarity and the level of ”privacy” of the data generated by a model w.r.t.
 the training set.
@@
 Fig.
-10 shows the comparison of $AA_{\rm TS}$ and privacy loss values obtained with our two models, demonstrating that the pre-trained RBM clearly outperforms the other model, and even achieves better results ($AA_{\rm TS}$ values much closer to 0.5) than those discussed in Yelmen et al.
-(2021; 2023).
+14 shows the comparison of $AA_{\rm TS}$ and privacy loss values obtained with our two models, demonstrating that the pre-trained RBM clearly outperforms the other model, and even achieves better results ($AA_{\rm TS}$ values much closer to 0.5) than those discussed in (yelmen2021creating).
 Figure 13: Evolution of the AATS (first row) and privacy loss (second row) indicator during the training of pretrained RBMs.
-The online estimation of the train and test LL is given at the third row.
-The overfitting is clearly detected in the HGD and MICKEY dataset and the divergence between the train and test LL coincides with a sharp rise in the privacy loss.
-Figure 10: Comparison of the evolution of overfitting metrics along the log-likelihood between a pretrained RBM (A) and a RBM trained from scratch using PCD (B).
+The online estimation of the train and test $L$ is given at the third row.
+The overfitting is clearly detected in the HGD and MICKEY dataset and the divergence between the train and test $L$ coincides with a sharp rise in the privacy loss.
+Figure 14: Comparison of the evolution of overfitting metrics along the log-likelihood between a RBM trained from scratch using PCD (A) and a pretrained RBM (B).
 (C) shows the comparison between the AASyn and AATruth between the two machines.
 The machines were selected in order to have a comparable privacy loss (D), right before starting to overfit.
-Figure 11: Comparison of log-likelihood estimation w.r.t. the exact ll on the MNIST-01 dataset.
-Figure 12: Comparison of log-likelihood estimation w.r.t. the exact ll on the MNIST dataset.
-Figure 13: Comparison of log-likelihood estimation w.r.t. the exact ll on the Ising dataset.
-Figure 14: Comparison of log-likelihood estimation w.r.t. the exact ll on the MICKEY dataset.
-Appendix I Extra figures
+Figure 15: Comparison of log-likelihood estimation w.r.t. the exact $l$ on the MNIST-01 dataset.
+Figure 16: Comparison of log-likelihood estimation w.r.t. the exact $l$ on the MNIST dataset.
+Figure 17: Comparison of log-likelihood estimation w.r.t. the exact $l$ on the Ising dataset.
+Figure 18: Comparison of log-likelihood estimation w.r.t. the exact $l$ on the MICKEY dataset.
+Appendix J Extra figures
 Figure 19: Samples generated with the low-rank RBM trained with 3 directions plus a bias.